### v1

# Measuring cosmic baryon density with FRB and GW data

Abstract
The dispersion measure (DM) analysis of extragalactic fast radio bursts (FRBs) has accounted for all cosmic baryons but remains limited by systematic uncertainties from various parameter degeneracies. We show that the prominent degeneracy between the baryon density ($\Omega_{\rm b}$) and the Hubble constant ($H_{0}$) can be disentangled through an independent $H_{0}$ value uniquely inferred from the absolute luminosity distance measurements of gravitational-wave (GW) standard sirens. By combining $104$ localized FRBs with $47$ GW events, we obtain a robust, late-Universe measurement of $\Omega_{\rm b}~{}$$=0.0488\pm 0.0064$ ($1\sigma$), in concordance with early-Universe observations of CMB + BBN. Notably, incorporating GW data helps not only avoid biases induced by the $H_{0}$ tension, but also mitigate those from the parameters of diffuse baryon fraction and DM distribution models. Although the current precision ($\sim 13\%$) is limited by sample size, the growing detections of both FRBs and GWs will make their synergy a powerful probe of low-redshift cosmology.
Cosmological parameters (339); Baryon density (139); Radio bursts (1339); Gravitational waves (678)
1 Introduction
The concordance cosmology predicts consistency among independent probes across all epochs (Ostriker & Steinhardt, 1995). However, while the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) provide precise estimates of the baryon content in the early Universe (Cooke et al., 2018), about $30\%$ baryons remain undetected in the local Universe (Shull et al., 2012). These missing baryons are thought to reside in diffuse ionized gas that is too faint to be directly observed, known as the “missing baryon problem” (Fukugita et al., 1998; Cen & Ostriker, 1999).
Fast radio bursts (FRBs) are bright, millisecond-duration radio transients of extragalactic origin (Lorimer et al., 2007; Thornton et al., 2013). Their observed dispersion measures (DMs) trace the integrated column density of free electrons along the line of sight (LoS), making them sensitive to baryons in the intergalactic medium (IGM) (McQuinn, 2013). Thus, FRBs with known redshifts derived from precise host galaxy localization can be used to trace the missing baryons (Deng & Zhang, 2014; Connor et al., 2025). Macquart et al. (2020) (hereafter Macquart20) first employed $5$ localized FRBs to estimate the present-day cosmic baryon density of $\Omega_{\rm b}=0.051^{+0.021}_{-0.020}$ at the $95\%$ confidence level, which agrees with the joint CMB and BBN result and likely resolves the local baryon problem. Yang et al. (2022) (hereafter Yang22) used an updated sample of $22$ localized FRBs and obtained $\Omega_{\rm b}=0.0490_{-0.0033}^{+0.0036}$ at the $1\sigma$ level, reinforcing the evidence for a complete cosmic baryon census. With the high event rate and increasing number of localized events, FRBs have emerged as a promising cosmological probe (Zhou et al., 2014; Gao et al., 2014); see Bhandari & Flynn (2021),Wu & Wang (2024), and Glowacki & Lee (2025) for recent reviews.
However, FRB cosmology based on DM is hindered by parameter degeneracies, which result in two main sources of systematic uncertainties (Kumar & Linder, 2019). The first is from extracting the IGM contribution ($\mathrm{DM}_{\rm IGM}$) from the total DM budget (see Eq. (1)). This is mainly due to the difficulty in modeling both the host contribution ($\mathrm{DM}_{\rm host}$) and the LoS inhomogeneity in the IGM. Macquart20 modeled their probability distributions and Zhang et al. (2020, 2021) further employed the IllustrisTNG simulations to derive priors on these empirical DM model parameters for different FRB populations, which effectively reduce biases induced by host galaxy diversity and IGM fluctuations. However, the second uncertainty from the analysis of the extracted $\mathrm{DM}_{\rm IGM}$ remains challenging to control. This is because its redshift relation (known as the Macquart relation) exhibits intrinsic degeneracies among $\Omega_{\rm b}$, the Hubble constant ($H_{0}$), and the baryon fraction in diffuse ionized gas ($f_{\rm d}$) (see Eq. (3)). Given the limited constraints on $f_{\rm d}$, it is commonly modeled as either a constant or a redshift-dependent function in the literature (Li et al., 2020; Qiang & Wei, 2020; Wang & Wei, 2023; Liu et al., 2025). Although $H_{0}$ has been precisely measured, estimates from early- (e.g., Planck 2018 CMB (Aghanim et al., 2020) (hereafter Planck18)) and late-Universe (e.g., SH0ES Cepheid-calibrated Type Ia supernovae (SNe) (Riess et al., 2022)) observations show a significant discrepancy now exceeding $5\sigma$, which has widely been known as the “$H_{0}$ tension” (Guo et al., 2019; Di Valentino et al., 2021; Abdalla et al., 2022; Hu & Wang, 2023). This inconsistency would introduce systematic uncertainties into the estimation of $\Omega_{\rm b}$ using FRB samples. It should be noted that Macquart20 and Yang22 derived their respective estimates of $\Omega_{\rm b}$ by assuming $H_{0}=70$ km s-1  Mpc-1and a flat prior over 6767676774747474 km s-1  Mpc-1, which provides a reasonable strategy for handling the second uncertainty given the $H_{0}$ tension (note that many studies also constrain $H_{0}$ by adopting early-Universe priors on $\Omega_{\rm b}h^{2}$; see, e.g., Hagstotz et al. (2022)). Alternatively, jointly analyzing FRBs and datasets that independently measure $H_{0}$ would also be welcome.
As an emerging cosmological probe, gravitational-wave (GW) standard sirens offer unique advantages for measuring $H_{0}$ (Schutz, 1986). This primarily arises from the fact that the amplitudes of GW waveforms are inversely proportional to the luminosity distances ($D_{\rm L}$) to the sources, allowing for a direct measurement of $D_{\rm L}$ through waveform analysis. Since $D_{\rm L}\propto 1/H_{0}$, GW standard sirens serve as a powerful and self-calibrating probe for determining $H_{0}$, provided that the redshifts of the GW sources are also known (Dalal et al., 2006; Cutler & Holz, 2009; Zhao et al., 2011; Cai & Yang, 2017; Abbott et al., 2017a, b, c; Chen et al., 2018; Cai et al., 2018; Wang et al., 2018; Zhang et al., 2019a, b; Zhang, 2019; Wang et al., 2020; Zhao et al., 2020b; Gray et al., 2020; Jin et al., 2020; Yu et al., 2020; Wang et al., 2022; Zhu et al., 2022a, b; Wu et al., 2023; Muttoni et al., 2023; Yun et al., 2023; Yu et al., 2024; Han et al., 2024; Feng et al., 2024; Song et al., 2024; Jin et al., 2024a, b; Li et al., 2024; Dong et al., 2024; Xiao et al., 2024; Zhu et al., 2024; Zheng et al., 2024; Han et al., 2025). During the first three observing runs of the LIGO, Virgo, and KAGRA (LVK) collaboration, approximately 100 GW events were detected and collected in the third Gravitational-Wave Transient Catalog (GWTC-3). Using 47 high–signal-to-noise ratio (SNR) GW standard sirens from GWTC-3, Abbott et al. (2023a) obtained a $\sim 10\%$ precision on $H_{0}$. Song et al. (2025) further combined these 47 GW standard sirens with an additional strong lensing system and achieved a model-independent constraint on $H_{0}$ with a precision of $\sim 8\%$. Therefore, the $H_{0}$ information uniquely provided by GW data enables a direct estimate of $\Omega_{\rm b}$ disentangled from $H_{0}$ in the Macquart relation, which offers new insight into resolving systematic effects in FRB cosmology.
Unlike previous studies that combined FRBs with traditional observations such as SNe (Jaroszynski, 2019; Gao et al., 2023) or CMB (Zhao et al., 2020a; Qiu et al., 2022), we highlight that combining with GWs inherently involves calibration-free distances and avoids reliance on early-time physics. More recently, Zhang & Zhang (2025) and Gao et al. (2025b) proposed a novel method based on FRB/persistent radio sources (PRS) associations, which awaits the detection of more such events. The FRB + GW synergy is promising, with numerous detections expected from both communities in next-generation surveys (Zhang et al., 2023).
In this Letter, we combine $104$ localized FRBs with $47$ GW events from the GWTC-3 catalog to jointly constrain $\Omega_{\rm b}$ and $H_{0}$, without relying on any cosmological priors. To control the aforementioned systematics, we allow all DM parameters [^1] In this context, we refer to the DM parameters as ($f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$), while the empirical DM model parameters refer specifically to ($F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$). The parameters are defined in Sec. 2.1. to vary freely and incorporate GW data to break parameter degeneracies. Note that we do not assume any FRB/GW associations (see Wei et al. (2018) and Qiang et al. (2025) for studies based on such associations).
2 Data and methodology
2.1 Localized FRB data
In this section, we first present the likelihood function for localized FRBs based on DM analysis following Macquart20, then the FRB sample used in this work.
The observed dispersion measure can be decomposed into several components:
$\rm{DM}_{\rm{obs}}=\rm{DM}_{\rm{MW,ISM}}+\rm{DM}_{\rm{MW,halo}}+\rm{DM}_{\rm{IGM}}+\rm{DM}_{\rm{host}},$ (1)
[…]
$P_{\rm{halo}}(\mathrm{DM}_{\rm{halo}})=\frac{1}{\sqrt{2\pi}\sigma_{\rm{halo}}}\exp\left[-\frac{(\mathrm{DM}_{\rm{halo}}-\mu_{\rm{halo}})^{2}}{2\sigma_{\rm{halo}}^{2}}\right].$ (2)
The average $\mathrm{DM}_{\mathrm{IGM}}$ in the standard $\Lambda$CDM (cosmological constant $\Lambda$ + cold dark matter) cosmology is given by the Macquart relation,
$\langle\mathrm{DM}_{\mathrm{IGM}}\rangle=$ $\frac{3cf_{\rm e}}{8\pi Gm_{\mathrm{p}}}\Omega_{\mathrm{b}}H_{0}f_{\mathrm{d}}\int_{0}^{z}\frac{(1+z)dz}{\sqrt{\Omega_{\rm m}(1+z)^{3}+(1-\Omega_{\rm m})}},$ (3)
where $\Omega_{\rm m}$ is the present matter density parameter, $c$ is the speed of light, $G$ is the gravitational constant, $m_{\rm{p}}$ is the proton mass, and $f_{\rm e}$ is the number of ionized electrons per baryon, which can be approximated as $7/8$ for $z<3$.
The IGM plasma is inhomogeneous, and the true value of $\rm DM_{\rm IGM}$ would vary significantly around the mean value $\langle{\rm{DM}}_{{{\rm{IGM}}}}\rangle$ due to the LoS variance caused by intersecting foreground galaxy halos. Thus, the probability distribution of $\rm DM_{\rm IGM}$ is
$P_{\rm{IGM}}(\Delta)=A\Delta^{-\beta}\exp\left[-\frac{\left(\Delta^{-\alpha}-C_{0}\right)^{2}}{2\alpha^{2}\sigma_{\rm{IGM}}^{2}}\right],~{}\Delta>0,$ (4)
where $\Delta\equiv{\rm{DM}}_{\rm{IGM}}/\left\langle{\rm{DM}}_{{{\rm{IGM}}}}\right\rangle$. $\alpha$ and $\beta$, which describe the inner gas density profile of halos, are best fitted as $\alpha=3$ and $\beta=3$ according to hydrodynamic simulations. The effective standard deviation $\sigma_{\mathrm{IGM}}=Fz^{-1/2}$ is parameterized by Macquart20, where $F$ is the parameter reflecting the galactic feedback. Here, $A$ is a normalization constant, and $C_{0}$ is chosen to ensure that the mean of this distribution is unity. Note that we include the factor $1/\langle{\rm DM_{\rm IGM}}\rangle$ suggested by (Zhang et al., 2025b) when normalizing Eq. (4).
The distribution of $\rm{DM}_{\rm{host}}$ is modeled as
$P_{\rm{host}}(\rm{DM}_{\rm{host}})$ $=\frac{1}{\sqrt{2\pi}\sigma_{\rm{host}}\mathrm{DM}_{\rm{host}}}$
×$\times\exp\left[-\frac{(\ln\mathrm{DM}_{\rm{host}}-\mu_{\rm{host}})^{2}}{2\sigma_{\rm{host}}^{2}}\right],$ (5)
where a correction $\mathrm{DM}_{\text{host }}\rightarrow\mathrm{DM}_{\text{host}}/(1+z)$ has been applied. The mean value and variance of the distribution are ${\rm e}^{\mu_{\rm host}}$ and ${\rm e}^{2\mu_{\rm host}+\sigma_{\rm{host}}^{2}}({\rm e}^{\sigma_{\rm{host}}^{2}}-1)$, respectively.
Thus, the total likelihood of the localized FRB data set $\{x_{\rm FRB}\}$ composed of $N_{\rm FRB}$ sources, is given by
$\mathcal{L}_{\mathrm{FRB}}(\{x_{\rm FRB}\})=\sum_{i=1}^{N_{\mathrm{FRB}}}P_{i}\left(\mathrm{DM}^{\prime}_{i}\right).$ (6)
For the $i$-th FRB, the possibility distribution of its extragalactic DM component ($\mathrm{DM}^{\prime}=\rm DM_{\rm IGM}+\rm DM_{\rm host}$) is computed by convolving Eqs. (2), (4), and (2.1):
$P_{i}\left(\mathrm{DM}^{\prime}_{i}\right)$ $=\int_{30\,\mathrm{pc\,cm}^{-3}}^{100\,\mathrm{pc\,cm}^{-3}}P_{\rm halo}\left(\mathrm{DM}_{\rm halo}\right)\,\mathrm{dDM}_{\rm halo}$
×$\times\int_{0}^{\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}}P_{\rm host}\left(\mathrm{DM}_{\rm host}\right)$
×$\times P_{\rm IGM}\left(\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}-\mathrm{DM}_{\rm host}\right)\,\mathrm{dDM}_{\rm host}.$ (7)
We also consider a scenario in which the empirical DM model parameters are fixed, though our main results treat them as free. Specifically, we adopt values calibrated from cosmological hydrodynamical simulations, notably the IllustrisTNG simulation, which provides redshift-dependent distributions for both $P_{\rm IGM}(\mathrm{DM}_{\rm IGM})$ (Zhang et al., 2021) and $P_{\rm host}(\mathrm{DM}_{\rm host})$ (Zhang et al., 2020). The parameters $A(z)$, $C_{0}(z)$, and $\sigma_{\rm IGM}(z)$ in Eq. (4), as well as $\mu_{\rm host}(z)$ and $\sigma_{\rm host}(z)$ in Eq. (2.1), are interpolated from best-fit simulation results across redshift. Note that $\mu_{\rm host}$ and $\sigma_{\rm host}$ are further fixed based on the host galaxy types (see the seventh column of Table 2).
Finally, we select a sample of $104$ localized FRBs from an initial set of $119$, based on the following criteria: (1) $\mathrm{DM}_{\mathrm{MW,ISM}}$ is required to be less than $40\%$ of $\mathrm{DM}_{\rm obs}$ (Connor et al., 2025); (2) $\mathrm{DM}_{\rm obs}-\rm DM_{MW,ISM}$ is taken to exceed 80 $\,\rm{pc}\,\rm{cm}^{-3}\,$, ensuring a positive $\mathrm{DM}_{\rm host}$; (3) FRB 20190520B and FRB 20220831A are excluded due to excess local DM contributions from their host environments. The final sample is listed in Appendix B (see Table 2).
2.2 GW standard siren data
We use 47 GW standard sirens with SNRs greater than 11 and an inverse false alarm rate exceeding 4 yr from GWTC-3, covering a redshift range of approximately $z<0.8$. The sample consists of 42 BBHs, three neutron star-black holes (NSBHs), and two binary neutron stars (BNSs).
We obtain the redshift of GW170817 from the electromagnetic (EM) counterpart (Abbott et al., 2017a, b, c). For GW events without EM counterparts, we obtain their redshifts through the dark siren method (Schutz, 1986; Mastrogiovanni et al., 2023; Gray et al., 2023) and incorporate the GLADE+ catalog (Dálya et al., 2022). In the following, we provide a brief introduction to the likelihood function of GW standard sirens. We denote the source parameters of each GW event by $\bm{\theta}$, and the population-level parameters by $\bm{\Lambda}$, including population model parameters and cosmological parameters. For GW events without EM counterparts, assuming a “scale-free” prior on the expected number of detected events $N_{\exp}$, i.e., $\pi(N_{\exp})\propto 1/N_{\exp}$, the likelihood function of the GW data set $\{x_{\rm GW}\}$ composed of $N_{\rm GW}$ GW events (Mastrogiovanni et al., 2021; Abbott et al., 2023b), is given by
$\mathcal{L}_{\rm GW}(\{x_{\rm GW}\}|\bm{\Lambda})\propto\prod_{i}^{N_{\rm GW}}\frac{\int{\rm d}\bm{\theta}\mathcal{L}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})\frac{{\rm d}N_{\rm CBC}(\bm{\Lambda})}{{\rm d}\bm{\theta}{\rm d}t}}{\int{\rm d}\bm{\theta}P_{\rm det}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})\frac{{\rm d}N_{\rm CBC}(\bm{\Lambda})}{{\rm d}\bm{\theta}{\rm d}t}},$ (8)
where $\mathcal{L}_{\rm GW}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})$ is the parameter estimation likelihood, quantifying the uncertainties in measuring the GW source parameters $\bm{\theta}$ through waveform analysis. ${\rm d}N_{\rm CBC}(\bm{\Lambda})/{\rm d}\bm{\theta}{\rm d}t$ is the CBC detection rate in the detector frame, which includes population models of GW sources, cosmological models, and redshift information from the GLADE+ galaxy catalog. In assessing the completeness of the GLADE+ catalog, we adopt a Schechter function to describe the $B$-band absolute magnitude distribution of galaxies (Kochanek et al., 2001). We further assume that the probability of a galaxy being the host of the GW source is proportional to its $B$-band luminosity. $P_{\rm det}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})$ is the detection probability of a GW event with parameters of $\bm{\theta}$ and $\bm{\Lambda}$. The denominator of Eq. (8) accounts for the GW selection effect.
For GW dark sirens, population models can help infer redshift for the sources, especially in scenarios where the galaxy catalog is significantly incomplete (Taylor et al., 2012; Ezquiaga & Holz, 2022; Mastrogiovanni et al., 2023; Gray et al., 2023). We adopt the same population model assumptions as Abbott et al. (2023a), but restrict our analysis to the Power Law + Peak BBH mass model, which is best supported by the data. Considering additional models would increase computational cost and lies beyond the scope of this work. For BNSs, we assume that neutron star masses are uniformly distributed in the range of 1–3 $M_{\odot}$. In the case of NSBH systems, the black hole component is drawn from the same distribution as the primary black holes in BBHs, while the neutron star component follows the mass distribution assumed for BNSs. The redshift distribution of GW sources is modeled using a phenomenological prescription that traces the cosmic star formation rate. Consistent with Abbott et al. (2023a), we disregard potential redshift evolution in the BBH mass distribution as well as variations in the GW detection rate arising from the spin distribution. These effects are currently subdominant compared to the statistical uncertainties present in the existing GW observations.
In cases where EM counterparts are detected, the information from EM observations should be incorporated alongside GW data. Since GW and EM detections are statistically independent, the overall likelihood is given by
$\mathcal{L}_{\rm EM+GW}(x_{{\rm GW},i}|\bm{\theta})\propto\mathcal{L}_{\rm EM}(x_{{\rm GW},i}|z,\Omega)\mathcal{L}_{\rm GW}(x_{{\rm GW},i}|\bm{\theta}),$ (9)
where $\mathcal{L}_{\rm EM}(x_{{\rm GW},i}|z,\Omega)$ quantifies the measurement uncertainties of the redshift $z$ and the sky localization $\Omega$ from EM observations. We assume that the probability of having EM observations is 1 once an EM counterpart is present (Gray et al., 2020; Mastrogiovanni et al., 2024). This assumption holds only within the current GW detection range, as it remains much smaller than that of EM observations.
2.3 Joint analysis of FRB and GW data
Figure 1: Normalized posterior distributions of $\Omega_{\rm b}~{}$from 104 FRBs (orange) and 104 FRBs + 47 GWs (indigo). The median and $1\sigma$ credible interval from FRB + GW is $\Omega_{\rm b}=0.0488\pm 0.0064$, marginalized over $f_{\rm d}$ and empirical DM model parameters. The errorbars in different colors and linestyles are shown under different assumptions: FRB + $H_{0}$ (Planck18 prior) and FRB + $H_{0}$ (SH0ES prior) adopt Gaussian priors on $H_{0}$ from Planck18 (Aghanim et al., 2020) and SH0ES (Riess et al., 2022), respectively; FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW adopt priors on empirical DM model parameters from the IllustrisTNG simulation (Zhang et al., 2020, 2021) and on $f_{\rm d}$ from the redshift-dependent form of Macquart20 (Macquart et al., 2020), respectively. The blue, gray, and red shaded bands show the results from CMB + BBN (Cooke et al., 2018), Macquart20, and Yang22 (Yang et al., 2022), respectively. Note that Yang22 adopts the TNG prior, a flat prior on $H_{0}$in $\mathcal{U}(67,73)$ km s-1 Mpc-1, and $f_{\rm d}=0.82$. All results are at the $1\sigma$ level.
We jointly constrain cosmological and FRB-related DM parameters, and adopt flat priors for the following seven parameters: $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ (Yang et al., 2022), $\Omega_{\rm b}\in\mathcal{U}(0.01,0.10)$, $H_{0}\in\mathcal{U}(20,140)~{}\rm km\ s^{-1}\ Mpc^{-1}$, $f_{\rm d}\in\mathcal{U}(0.84,0.96)$, [^2] This prior on $f_{\rm d}$ is informed by observations of the baryon fraction in stars ($f_{*}$) and cold gas ($f_{\rm cold}$), independent of FRBs. Assuming a conservative flat prior for $f_{*}$ extended by a factor of five, the diffuse baryon fraction is inferred via $f_{\rm d}=1-f_{*}-f_{\rm cold}$ (Connor et al., 2025). $F\in\mathcal{U}(0.011,1.0)$, $\mu_{\rm host}\in\mathcal{U}(3.0,5.3)$, and $\sigma_{\rm host}\in\mathcal{U}(0.2,2.0)$ (Macquart et al., 2020). The GW population model parameters include $\alpha$, $\beta$, $m_{\rm min}$, $m_{\rm max}$, $\Delta m$, $\mu_{\rm g}$, $\sigma_{\rm g}$, $\lambda_{\rm peak}$, $\gamma$, $\kappa$, and $z_{\rm p}$, which are fixed to the inferred values of Abbott et al. (2023a).
We utilize the bilby [^3] https://github.com/bilby-dev/bilby package to perform Markov Chain Monto Carlo (MCMC) analysis and choose the emcee [^4] https://emcee.readthedocs.io/en/stable sampler. We use the ICAROGW [^5] https://github.com/simone-mastrogiovanni/icarogw package to do the hierarchic Bayesian analysis of GW standard sirens, and the GetDist [^6] https://github.com/cmbant/getdist package to plot the marginalized posterior distributions. The final likelihood $\mathcal{L}$ used in our analysis is the combination of $\mathcal{L}_{\mathrm{FRB}}$ and $\mathcal{L}_{\rm GW}$ in Eqs. (6) and (8), respectively, which is written as
$\ln\mathcal{L}=\ln\left(\mathcal{L}_{\mathrm{FRB}}\right)+\ln\left(\mathcal{L}_{\mathrm{GW}}\right).$ (10)
Table 1: $1\sigma$ constraints on cosmological parameters $\Omega_{\rm b}$ and $H_{0}$.
Data $\Omega_{\rm b}$ $H_{0}$
FRB $0.0520^{+0.0120}_{-0.0280}$ $77.0^{+20.0}_{-40.0}$
FRB + GW $0.0488\pm 0.0064$ $71.6^{+4.4}_{-8.6}$
FRB (TNG) + GW $0.0491\pm 0.0054$ $71.2^{+4.4}_{-8.0}$
FRB (TNG + $f_{\rm d}(z)$) + GW $0.0510^{+0.0055}_{-0.0050}$ $71.8^{+4.5}_{-8.6}$
FRB + $H_{0}$ (Planck18 prior) $0.0513\pm 0.0046$ $67.4\pm 0.5$
FRB + $H_{0}$ (SH0ES prior) $0.0473\pm 0.0043$ $73.0\pm 1.0$
3 Results and discussions
Figure 2: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm b}$ and $\Omega_{\rm b}h_{70}$ versus $H_{0}$, $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$, using FRB (orange) and FRB + GW (purple) data. The correlation coefficient $r$ is labeled in each panel to quantify parameter degeneracies.
In this section, we report and discuss the constraints on $\Omega_{\rm b}~{}$in Table 1. The posteriors for $\Omega_{\rm b}~{}$and the corresponding credible intervals are shown in Figure 1. Using FRB data alone yields a relatively weak constraint of $\Omega_{\mathrm{b}}=0.052_{-0.028}^{+0.012}$ at the $1\sigma$ level. Incorporating GW standard sirens significantly narrows this to $\Omega_{\mathrm{b}}=0.0488\pm 0.0064$, by breaking the strong degeneracy between $\Omega_{\mathrm{b}}$ and $H_{0}$ via the independent $H_{0}$ measurement provided by GWs. Remarkably, this late-Universe estimate exhibits concordance with early-Universe values from BBN and CMB at the $\sim 13\%$ level, linking the gap between the early and late Universe. Note that a similarly precise constraint is simultaneously obtained for the Hubble constant, $H_{0}=71.6^{+4.4}_{-8.6}~{}\mathrm{kms^{-1}~{}Mpc^{-1}}$, primarily driven by the GW dataset alone, which gives $H_{0}=72.6^{+5.2}_{-8.9}~{}\mathrm{kms^{-1}~{}Mpc^{-1}}$. More data is required to arbitrate the $H_{0}$ tension. A detailed discussion on $H_{0}$ and DM parameters is provided in Appendix A.
We compare our $\Omega_{\rm b}~{}$results with those from previous works. Macquart20 and Yang22 constrained $\Omega_{\rm b}~{}$using different samples of localized FRBs under external assumptions of $H_{0}=70$ km s-1  Mpc-1and a flat prior on $H_{0}\in\mathcal{U}(67,73)$ km s-1  Mpc-1, as shown by the gray and red bands in Figure 1, respectively. The constraint from Yang22 compares favorably to ours, since the current GW standard siren sample remains too limited to outperform their adopted $H_{0}$ prior. Applying the external $H_{0}$ priors from Planck18 or SH0ES to our FRB sample alone yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively; the corresponding error bars are shown in Figure 1. This suggests that the $H_{0}$ tension would lead to two biased $\Omega_{\rm b}~{}$estimates. Although the induced “$\Omega_{\rm b}$ inconsistency” in our sample is within $1\sigma$ uncertainty ($\sim 0.95\sigma$), it may become non-negligible with larger samples in updated DM analyses. Nevertheless, the joint FRB + GW result assumes no prior on $H_{0}$ and instead provides a late-Universe measurement of baryons, independent of both CMB and local distance ladder. This FRB + GW synergy may establish a novel paradigm for the cosmic baryon census, analogous to CMB + BBN, and serve as a powerful low-redshift cosmological probe in the era of next-generation FRB and GW surveys (Zhang et al., 2023).
We also investigate potential improvements to our results. FRB cosmology is subject to uncontrolled systematics mainly from modeling $\mathrm{DM}_{\mathrm{IGM}}$ and $\mathrm{DM}_{\mathrm{host}}$ distributions, as well as the weakly constrained $f_{\rm d}$, which affects the mean value of $\mathrm{DM}_{\mathrm{IGM}}$. To reduce these systematics, cosmological simulations are commonly used to calibrate these DM parameters ($F$, $\mu_{\rm host}$, $\sigma_{\rm host}$, and $f_{\rm d}$) at different redshifts. Accordingly, we consider two cases with simulation-based priors: (i) fixing the empirical DM model parameters to the best-fit results from the IllustrisTNG simulation (hereafter TNG) (Zhang et al., 2020, 2021); and (ii) further fixing $f_{\rm d}$ to a redshift-dependent form following Macquart20, as implemented in the public FRB code [^7] https://github.com/FRBs/FRB. The constraints from FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW are $\Omega_{\mathrm{b}}=0.0491\pm 0.0054$ and $\Omega_{\mathrm{b}}=0.0507\pm 0.0052$, respectively, with the corresponding errorbars shown in Figure 1. Both results are consistent with the FRB + GW constraint within the $1\sigma$ uncertainty, suggesting that the simulation-based priors can improve current constraints without introducing significant bias. However, such agreement at the current level does not ensure the absence of systematic errors when applied to larger future FRB and GW samples, in which systematic errors may become dominant. Therefore, a detailed assessment of how the DM parameters affect the inferred baryon density is of great importance.
To assess the impact of systematic uncertainties on $\Omega_{\mathrm{b}}$ and verify the robustness of our constraint, we analyze the parameter correlations using the FRB + GW dataset. The lower row of Figure 2 shows the marginalized posteriors of $\Omega_{\rm b}~{}$and other relevant parameters, with correlation coefficients $r$ labeled in each panel to quantify parameter degeneracies.
On the one hand, the degeneracy between $\Omega_{\rm b}$ and $H_{0}$ is the strongest among all parameter pairs ($r=-0.73$), implying that the inferred value of $\Omega_{\rm b}$ is highly sensitive to how $H_{0}$ is determined. While FRB + $H_{0}$ analyses rely on external priors, which are controversial due to the $H_{0}$ tension, FRB + GW analyses rely on GW data to provide $H_{0}$, which is uniquely derived from self-calibrating absolute distance measurements. Therefore, the $\Omega_{\rm b}$ constraint from FRB + GW is robust against the $H_{0}$-related bias and free from assumptions affected by the $H_{0}$ tension.
On the other hand, the correlations between $\Omega_{\rm b}~{}$and DM parameters are moderate. These DM parameters are not properly constrained, both in this work (see Appendix A for details) and earlier studies (James et al., 2022), and thus act as nuisance parameters that bias the inferred $\Omega_{\rm b}$. In the joint FRB + GW analysis, the correlation coefficients between $\Omega_{\rm b}$ and $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$ are $-0.28$, $0.40$, $-0.40$, and $0.17$, respectively. For comparison, the upper row of Figure 2 shows the constraint on $\Omega_{\rm b}h_{70}$, where $h_{70}=H_{0}/(70~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}})$. This parameter absorbs the strong degeneracy, enabling it to be effectively constrained using FRB data alone (Macquart et al., 2020; Connor et al., 2025). In this case, the corresponding correlations are notably stronger: $-0.41$, $0.57$, $-0.62$, and $0.27$, respectively. This comparison suggests that directly measuring $\Omega_{\rm b}$ reduces degeneracies with the nuisance parameters and is more robust against the associated systematic uncertainties than $\Omega_{\rm b}h_{70}$.
We conclude that incorporating GW data into the analysis of FRB dispersion measures enables a direct and effective measurement of $\Omega_{\rm b}$ and, more importantly, avoids biases from the $H_{0}$ tension and mitigates those associated with nuisance model parameters. This demonstrates that the synergy between FRBs and GWs provides a robust estimate of the cosmic baryon density.
4 Conclusion
In this Letter, we combine two emerging cosmological probes, i.e., $104$ localized FRBs and $47$ GW standard sirens, to measure the cosmic baryon density in a robust manner. DM-based FRB cosmology struggles with various parameter degeneracies, particularly between $H_{0}$ and $\Omega_{\rm b}$. GWs can independently infer $H_{0}$ from self-calibrated luminosity distances obtained through waveform analysis as sirens, thus breaking the $\Omega_{\rm b}$-$H_{0}$ degeneracy. This joint analysis yields a constraint of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), which is remarkably consistent with early-Universe values from BBN and CMB at the $\sim 13\%$ level and provides a purely late-Universe perspective on resolving the missing baryon problem. We also obtain a simultaneous constraint on the Hubble constant, $H_{0}=71.6^{+4.4}_{-8.6}~{}\mathrm{km\,s^{-1}\,Mpc^{-1}}$, primarily driven by the GW data. Together, these results exhibit $\Omega_{\rm b}$-$H_{0}$ concordance between the early and late Universe.
Moreover, we explore two prior-based scenarios: (i) $H_{0}$-prior-based cases using FRB data alone, where adopting Gaussian priors from Planck18 and SH0ES yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively. These results indicate that the $H_{0}$ tension would bias baryon density estimates, although the induced “$\Omega_{\rm b}~{}$inconsistency” remains within $1\sigma$ uncertainty in our sample. (ii) Simulation-based prior cases using FRB + GW, where fixing empirical DM model parameters from the IllustrisTNG simulation and further fixing $f_{\rm d}$ to a redshift-dependent form yield $\Omega_{\rm b}=0.0491\pm 0.0054$ and $0.0507\pm 0.0052$, respectively. These two results agree with that of FRB + GW at the $1\sigma$ level, indicating that simulation-based priors can improve current constraints without introducing significant bias. However, in both (i) and (ii), systematic effects from external assumptions (related to $H_{0}$, $f_{\rm d}$, and empirical DM model parameters) may become non-negligible when applied to larger future FRB and GW samples.
Hence, we assess the associated systematic uncertainties through a detailed correlation analysis. On the one hand, the dominant uncertainty is the degeneracy between $\Omega_{\rm b}$ and $H_{0}$, implying the importance of carefully addressing the $H_{0}$-related bias. Unlike prior-based analyses that rely on assumed values of $H_{0}$, the joint FRB + GW analysis avoids such bias by uniquely inferring $H_{0}$ from the absolute distances provided by GWs, independent of both CMB and local distance ladder. On the other hand, the correlations between $\Omega_{\rm b}$ and DM parameters—treated as nuisance parameters—are moderate, and notably weaker than those of $\Omega_{\rm b}h_{70}$, making the direct inference of $\Omega_{\rm b}$ more robust against the uncertainties induced by the nuisance parameters. Therefore, combining FRB and GW data not only avoids the bias induced by the $H_{0}$ tension, but also mitigates model-dependent uncertainties.
Overall, the novel combination of FRBs and GWs enables a direct, unique, and robust measurement of both the baryon content and cosmological distances at low redshifts. With an increasing number of GW detections expected in the upcoming LVK run, and potentially thousands of localized FRBs expected from CHIME/Outriggers (Masui, 2023) and DSA-110 (Law et al., 2024), their synergy will soon emerge as a powerful low-redshift cosmological probe.
5 Acknowledgments
We thank Tian-Nuo Li, Meng-Lin Zhao, Yang Liu, and Qin Wu for fruitful discussions. This work was supported by the National SKA Program of China (grant Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (grant Nos. 12473001, 11975072, 11875102, and 11835009), the China Manned Space Program (grant No. CMS-CSST-2025-A02), and the National 111 Project (grant No. B16009).
Appendix A Constraints on $H_{0}$ and DM parameters
Figure 3: Same as Figure 1 but for posterior distribution of $H_{0}$. The median and $1\sigma$ credible interval from FRB + GW is $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1. The orange and black errorbars correspond to constraints from FRB + $\Omega_{\rm b}h^{2}$ (BBN prior (Cooke et al., 2018)) and GW data alone, respectively. The green, yellow, red, and gray shaded bands denote the results from Planck18 (Aghanim et al., 2020), SH0ES (Riess et al., 2022)), Wu22 (Wu et al., 2022), and James22 (James et al., 2022), respectively. Note that Wu22 adopts the TNG prior, and both Wu22 and James22 adopt $\Omega_{\rm b}h^{2}$ priors from BBN (Cooke et al., 2018) and CMB (Aghanim et al., 2020), respectively. All results are at $1\sigma$ uncertainty.
In this Appendix, we report and discuss the constraints on $H_{0}$ and DM parameters ($f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$).
The constraints on $H_{0}$ are listed in Table 1, and posteriors and corresponding credible intervals for $H_{0}$ are shown in Figure 3. A direct measurement of $H_{0}$ using FRBs alone is difficult due to strong parameter degeneracies. Specifically, our measurement using FRBs alone is $H_{0}=77.0^{+20.0}_{-40.0}$ km s-1  Mpc-1, and the inclusion of GW data narrows the constraint to $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1. This improvement is mainly driven by the GW data itself, which yields $H_{0}=72.6^{+5.2}_{-8.9}$ km s-1  Mpc-1. Previous studies predominantly inferred $H_{0}$ by fixing the baryon density parameter at the early-Universe value. For example, Wu et al. (2022) (hereafter Wu22) and James et al. (2022) (hereafter James22) used different FRB samples to derive $H_{0}=69.31^{+6.21}_{-6.63}~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$ and $H_{0}=73^{+12}_{-8}~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$, by assuming external priors on $\Omega_{\rm b}h^{2}\sim\mathcal{N}(0.02235,0.00049)$ from BBN and $\Omega_{\rm b}h^{2}=0.02242$ from Planck18, respectively, which are shown as red and gray bands in Figure 3. Recently, Wang et al. (2025) and Gao et al. (2025a) also used $92$ and $108$ localized FRBs to obtain $H_{0}=69.04_{-2.07}^{+2.30}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ and $H_{0}=69.40_{-1.97}^{+2.14}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$, respectively, both assuming BBN and TNG priors. Using the same BBN prior, we obtain $H_{0}=66.2^{+5.3}_{-7.3}$ km s-1  Mpc-1from FRB + $\Omega_{\rm b}h^{2}$, which slightly favors the Planck18 result with a precision comparable to that of FRB + GW. Unlike prior-based methods, the FRB + GW constraint provides a late-Universe result of $H_{0}$. More data is required to offer further insight into arbitrating the $H_{0}$ tension. Notably, time-delay cosmography using strongly lensed FRBs by massive galaxies can offer an independent measurement of $H_{0}$, since the time delays between lensed FRB images can be determined with exceptional precision (Li et al., 2018; Zhang et al., 2025a).
For DM parameters, the constraints are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$. The inclusion of GW data has a negligible impact on this constraint. For the host contribution, the results of $\mu_{\rm host}$ and $\sigma_{\rm host}$ correspond to a median rest-frame host DM of $108\pm 20~{}\mathrm{pc~{}cm^{-3}}$ (corresponding mean DM is $168^{+57}_{-40}~{}\mathrm{pc~{}cm^{-3}}$), which is consistent with both $\mathrm{DM}_{\rm host}=103_{-48}^{+68}~{}\mathrm{pc~{}cm^{-3}}$ (Yang et al., 2025) and $\mathrm{DM}_{\rm host}=130_{-23}^{+25}~{}\mathrm{pc~{}cm^{-3}}$ (Connor et al., 2025).
Appendix B FRB data
In this Appendix, we list the properties of the $104$ localized FRBs used in this work in Table 2.
[…]
Table 2: Properties of $104$ localized FRBs.
[…]
(pc cm-3) (pc cm-3) () ()
FRB 20121102A 0.19273 557 188.4 82.9946 $33.1479$ 1 Chatterjee et al. (2017) Tendulkar et al. (2017)
FRB 20180301A 0.3304 552 151.7 93.2268 $4.6711$ 1 Bhandari et al. (2022)
FRB 20180924B 0.3214 361.42 40.5 326.1053 $-40.9$ 3 Bannister et al. (2019)
FRB 20181112A 0.4755 589.27 41.7 327.3485 $-52.9709$ 3 Prochaska et al. (2019)
FRB 20181223C 0.03024 111.61 19.9 180.9207 $27.5476$ 3 Bhardwaj et al. (2024)
FRB 20190102C 0.2912 364.5 57.4 322.4157 $-79.4757$ 3 Bhandari et al. (2020)
FRB 20190110C 0.12244 221.6 37.1 249.3185 $41.4434$ 2 Ibik et al. (2024)
FRB 20190303A 0.064 223.2 29.8 207.9958 $48.1211$ 2 Michilli et al. (2023)
FRB 20190418A 0.07132 182.78 70.2 65.8123 $16.0738$ 3 Bhardwaj et al. (2024)
FRB 20190523A 0.66 760.8 37.2 207.065 $72.4697$ 3 Ravi et al. (2019)
FRB 20190608B 0.11778 338.7 37.3 334.0199 $-7.8983$ 3 Chittidi et al. (2021)
FRB 20190611B 0.3778 321.4 57.8 320.7456 $-79.3976$ 3 Heintz et al. (2020)
FRB 20190614D 0.6 959.2 87.8 65.0755 $73.7067$ 3 Law et al. (2020)
FRB 20190711A 0.522 593.1 56.5 329.4193 $-80.358$ 1 Heintz et al. (2020)
FRB 20190714A 0.2365 504.13 38.5 183.9797 $-13.021$ 3 Heintz et al. (2020)
FRB 20191001A 0.234 506.92 44.2 323.3513 $-54.7478$ 3 Heintz et al. (2020)
FRB 20191106C 0.10775 332.2 25 199.5801 $42.9997$ 2 Ibik et al. (2024)
FRB 20191228A 0.2432 297.5 32.9 344.4304 $-28.5941$ 3 Bhandari et al. (2022)
FRB 20200223B 0.06024 201.8 45.6 8.2695 $8.83132$ Ibik et al. (2024)
FRB 20200430A 0.1608 380.1 27.2 229.7064 $12.3763$ 3 Heintz et al. (2020)
FRB 20200723B 0.0085 244 33 190.1583 $-5.135$ 3 Shin et al. (2024)
FRB 20200906A 0.3688 577.8 35.8 53.4962 $-14.0832$ 3 Bhandari et al. (2022)
FRB 20201124A 0.098 413.52 139.9 77.0146 $6.06072$ Ravi et al. (2022)
FRB 20210117A 0.2145 729.1 34.4 339.9792 $-16.1515$ 3 Bhandari et al. (2023)
FRB 20210320C 0.2797 384.8 39.3 204.4608 $-16.1227$ 3 Shannon et al. (2025)
FRB 20210410D 0.1415 571.2 56.2 326.0863 $-79.3182$ 3 Caleb et al. (2023)
FRB 20210603A 0.1772 500.147 39.5 10.2741 $21.2263$ 3 Cassanelli et al. (2024)
FRB 20210807D 0.1293 251.3 121.2 299.2214 $-0.7624$ 3 Gordon et al. (2023)
FRB 20211127I 0.0469 234.83 42.5 199.8082 $-18.8378$ 3 Gordon et al. (2023)
FRB 20211203C 0.3439 636.2 63.7 204.5625 $-31.3801$ 3 Gordon et al. (2023)
FRB 20211212A 0.0715 206 27.1 157.3509 $1.676889$ 3 Gordon et al. (2023)
FRB 20220105A 0.2785 583 22 208.8039 $22.4665$ 3 Gordon et al. (2023)
FRB 20220204A 0.4012 612.584 50.7 274.2263 $69.7225$ 3 Sharma et al. (2024)
FRB 20220207C 0.04304 262.38 76.1 310.1995 $72.8823$ 3 Law et al. (2024)
FRB 20220208A 0.351 437 101.6 319.3483 $71.54$ 3 Sharma et al. (2024)
FRB 20220307B 0.2507 499.15 128.2 350.8745 $72.1924$ 3 Law et al. (2024)
FRB 20220310F 0.477958 462.24 46.3 134.7204 $73.4908$ 3 Law et al. (2024)
FRB 20220330D 0.3714 468.1 38.6 165.7256 $71.7535$ 3 Sharma et al. (2024)
FRB 20220418A 0.622 623.25 36.7 219.1056 $70.0959$ 3 Law et al. (2024)
FRB 20220501C 0.381 449.5 30.6 352.3792 $-32.4907$ 3 Shannon et al. (2025)
FRB 20220506D 0.30039 396.97 84.6 318.0448 $72.8273$ 3 Law et al. (2024)
FRB 20220509G 0.0894 269.53 55.6 282.67 $70.2438$ 3 Law et al. (2024)
FRB 20220521B 1.354 1342.9 138.8 351.036 $71.138$ 3 Shannon et al. (2025)
FRB 20220529A 0.1839 246 40 19.1042 $20.6325$ 1 Li et al. (2025)
FRB 20220610A 1.016 1458.15 31 351.0732 $-33.5137$ 3 Ryder et al. (2023)
FRB 20220717A 0.36295 637.34 118.3 293.3042 $-19.2877$ 3 Rajwade et al. (2024)
FRB 20220725A 0.1926 290.4 30.7 353.3152 $-35.9902$ 3 Shannon et al. (2025)
FRB 20220726A 0.361 686.55 89.5 73.94567 $69.9291$ 3 Sharma et al. (2024)
FRB 20220825A 0.241397 651.24 78.5 311.9815 $72.585$ 3 Law et al. (2024)
FRB 20220914A 0.1139 631.28 54.7 282.0568 $73.3369$ 3 Law et al. (2024)
FRB 20220918A 0.491 656.8 153.1 17.5921 $70.8113$ 3 Shannon et al. (2025)
FRB 20220920A 0.158239 314.99 39.9 240.2571 $70.9188$ 3 Law et al. (2024)
FRB 20221012A 0.284669 441.08 54.3 280.7987 $70.5242$ 3 Law et al. (2024)
FRB 20221027A 0.229 452.5 47.2 129.6104 $71.7315$ 3 Sharma et al. (2024)
FRB 20221029A 0.975 1391.75 43.8 141.9634 $72.4523$ 3 Sharma et al. (2024)
FRB 20221101B 0.2395 491.554 131.2 342.2162 $70.6812$ 3 Sharma et al. (2024)
FRB 20221106A 0.2044 343.8 34.8 56.7048 $-25.5698$ 3 Shannon et al. (2025)
FRB 20221113A 0.2505 411.027 91.7 71.411 $70.3074$ 3 Sharma et al. (2024)
FRB 20221116A 0.2764 643.448 132.3 21.2102 $72.6539$ 3 Sharma et al. (2024)
FRB 20221219A 0.553 706.708 44.4 257.6298 $71.6268$ 3 Sharma et al. (2024)
FRB 20230124A 0.0939 590.574 38.6 231.9163 $70.9681$ 3 Sharma et al. (2024)
FRB 20230203A 0.1464 420.1 67.299 151.66159 $5.69413$ Amiri et al. (2025)
FRB 20230216A 0.531 828 38.5 155.9717 $1.4678$ 3 Sharma et al. (2024)
FRB 20230222A 0.1223 706.1 33.31 106.96036 $11.22452$ 3 Amiri et al. (2025)
FRB 20230222B 0.11 187.8 56.9977 238.73912 $0.89873$ Amiri et al. (2025)
FRB 20230307A 0.271 608.9 37.6 177.7813 $71.6956$ 3 Sharma et al. (2024)
FRB 20230311A 0.1918 364.3 67.2352 91.10966 $55.94595$ 3 Amiri et al. (2025)
FRB 20230501A 0.301 532.5 125.6 340.0272 $70.9222$ 3 Sharma et al. (2024)
FRB 20230506C 0.3896 772 68 12.0998 $42.0061$ 2 Anna-Thomas et al. (2025)
FRB 20230526A 0.157 361.4 31.9 22.2326 $-52.7173$ 3 Shannon et al. (2025)
FRB 20230626A 0.327 451.2 39.2 240.7125 $71.7142$ 3 Sharma et al. (2024)
FRB 20230628A 0.1265 344.952 39 166.7867 $72.2818$ 3 Sharma et al. (2024)
FRB 20230703A 0.1184 291.3 38.1566 184.62445 $48.72993$ 3 Amiri et al. (2025)
FRB 20230708A 0.105 411.51 60.3 303.1155 $-55.3563$ 3 Shannon et al. (2025)
FRB 20230712A 0.4525 587.567 39.2 167.3585 $72.5578$ 3 Sharma et al. (2024)
FRB 20230730A 0.2115 312.5 114.6465 54.66456 $3.15933$ Amiri et al. (2025)
FRB 20230808F 0.3472 653.2 31 53.3041 $-51.9353$ 3 Hanmer et al. (2025)
FRB 20230814A 0.553 696.4 104.8 335.9748 $73.0259$ 3 Connor et al. (2025)
FRB 20230902A 0.3619 440.1 34.1 52.1398 $-47.3335$ 3 Shannon et al. (2025)
FRB 20230930A 0.0925 456 70 10.5 $41.4$ 3 Anna-Thomas et al. (2025)
FRB 20231011A 0.0783 186.3 58.5524 18.2411 $41.7491$ 3 Amiri et al. (2025)
FRB 20231017A 0.245 344.2 31.1515 346.75429 $6.652683$ Amiri et al. (2025)
FRB 20231025B 0.3238 368.7 65.5787 270.78807 $63.98908$ 3 Amiri et al. (2025)
FRB 20231120A 0.07 438.9 43.8 143.6169 $71.7574$ 3 Sharma et al. (2024)
FRB 20231123A 0.0729 302.1 37.1085 82.62325 $4.47554$ 3 Amiri et al. (2025)
FRB 20231123B 0.2625 396.7 40.2 242.5382 $70.7851$ 3 Sharma et al. (2024)
FRB 20231128A 0.1079 331.6 64.7368 199.5782 $42.99271$ 1 Amiri et al. (2025)
FRB 20231204A 0.0644 221 34.9372 207.99903 $48.116$ 2 Amiri et al. (2025)
FRB 20231206A 0.0659 457.7 37.5114 112.44284 $56.25627$ 3 Amiri et al. (2025)
FRB 20231220A 0.3355 491.2 49.9 123.9087 $73.6599$ 3 Connor et al. (2025)
FRB 20231226A 0.1569 329.9 38.1 155.3638 $6.1103$ 3 Shannon et al. (2025)
FRB 20231229A 0.019 198.5 65.1335 26.46783 $5.112923$ Amiri et al. (2025)
FRB 20231230A 0.0298 131.4 32.2305 72.79761 $2.39398$ 3 Amiri et al. (2025)
FRB 20240114A 0.13 527.65 49.7 321.9161 $4.3292$ 1 Tian et al. (2024)
FRB 20240119A 0.376 483.1 38 224.4672 $71.6118$ 3 Connor et al. (2025)
FRB 20240123A 0.968 1462 90.2 68.2625 $71.9453$ 3 Connor et al. (2025)
FRB 20240201A 0.042729 374.5 38.6 149.9056 $14.088$ 3 Shannon et al. (2025)
FRB 20240209A 0.1384 176.518 63.1 289.8504 $86.0609$ 2 Eftekhari et al. (2025)
FRB 20240210A 0.023686 283.73 28.7 8.7796 $-28.2708$ 3 Shannon et al. (2025)
FRB 20240213A 0.1185 357.4 40 166.1683 $74.0754$ 3 Connor et al. (2025)
FRB 20240215A 0.21 549.5 47.9 268.4413 $70.2324$ 3 Connor et al. (2025)
FRB 20240229A 0.287 491.15 38 169.9835 $70.6762$ 3 Connor et al. (2025)
FRB 20240304A 0.2423 652.6 73.9 136.3304 $-16.2282$ 3 Shannon et al. (2025)
FRB 20240310A 0.127 601.8 30.1 17.6219 $-44.4394$ 3 Shannon et al. (2025)
Note. — The results are derived from $104$ localized FRBs alone or in combination with $47$ GW events, under different assumptions. For all cases, the $1\sigma$ constraints on DM parameters are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$. $H_{0}$ is in units of km s-1  Mpc-1.

### v2

@@
-Measuring cosmic baryon density with FRB and GW data
+Cosmic baryon census with fast radio bursts and gravitational waves
 Abstract
-The dispersion measure (DM) analysis of extragalactic fast radio bursts (FRBs) has accounted for all cosmic baryons but remains limited by systematic uncertainties from various parameter degeneracies.
-We show that the prominent degeneracy between the baryon density ($\Omega_{\rm b}$) and the Hubble constant ($H_{0}$) can be disentangled through an independent $H_{0}$ value uniquely inferred from the absolute luminosity distance measurements of gravitational-wave (GW) standard sirens.
-By combining $104$ localized FRBs with $47$ GW events, we obtain a robust, late-Universe measurement of $\Omega_{\rm b}~{}$$=0.0488\pm 0.0064$ ($1\sigma$), in concordance with early-Universe observations of CMB + BBN.
-Notably, incorporating GW data helps not only avoid biases induced by the $H_{0}$ tension, but also mitigate those from the parameters of diffuse baryon fraction and DM distribution models.
+The cosmic baryon density fraction ($\Omega_{\rm b}$) is intrinsically correlated with the Hubble constant ($H_{0}$) through the critical density of the Universe.
+In the context of the decade-long $H_{0}$ tension, the significant discrepancy between early- and late-Universe measurements of $H_{0}$ implies that fixing its value or imposing an external prior could bias the baryon census.
+To address this concern, we construct a late-Universe probe framework that unifies fast radio bursts (FRBs) and gravitational-wave (GW) standard sirens, which can respectively resolve the “missing baryon” problem and the $H_{0}$ tension through their dispersion measures and absolute luminosity distances.
+By combining $104$ localized FRBs with $47$ GW events, we obtain an $H_{0}$-free measurement of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), in concordance with early-Universe observations of CMB + BBN.
 Although the current precision ($\sim 13\%$) is limited by sample size, the growing detections of both FRBs and GWs will make their synergy a powerful probe of low-redshift cosmology.
-Cosmological parameters (339); Baryon density (139); Radio bursts (1339); Gravitational waves (678)
-1 Introduction
-The concordance cosmology predicts consistency among independent probes across all epochs (Ostriker & Steinhardt, 1995).
-However, while the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) provide precise estimates of the baryon content in the early Universe (Cooke et al., 2018), about $30\%$ baryons remain undetected in the local Universe (Shull et al., 2012).
-These missing baryons are thought to reside in diffuse ionized gas that is too faint to be directly observed, known as the “missing baryon problem” (Fukugita et al., 1998; Cen & Ostriker, 1999).
-Fast radio bursts (FRBs) are bright, millisecond-duration radio transients of extragalactic origin (Lorimer et al., 2007; Thornton et al., 2013).
-Their observed dispersion measures (DMs) trace the integrated column density of free electrons along the line of sight (LoS), making them sensitive to baryons in the intergalactic medium (IGM) (McQuinn, 2013).
-Thus, FRBs with known redshifts derived from precise host galaxy localization can be used to trace the missing baryons (Deng & Zhang, 2014; Connor et al., 2025).
-Macquart et al.
-(2020) (hereafter Macquart20) first employed $5$ localized FRBs to estimate the present-day cosmic baryon density of $\Omega_{\rm b}=0.051^{+0.021}_{-0.020}$ at the $95\%$ confidence level, which agrees with the joint CMB and BBN result and likely resolves the local baryon problem.
-Yang et al.
-(2022) (hereafter Yang22) used an updated sample of $22$ localized FRBs and obtained $\Omega_{\rm b}=0.0490_{-0.0033}^{+0.0036}$ at the $1\sigma$ level, reinforcing the evidence for a complete cosmic baryon census.
-With the high event rate and increasing number of localized events, FRBs have emerged as a promising cosmological probe (Zhou et al., 2014; Gao et al., 2014); see Bhandari & Flynn (2021),Wu & Wang (2024), and Glowacki & Lee (2025) for recent reviews.
-However, FRB cosmology based on DM is hindered by parameter degeneracies, which result in two main sources of systematic uncertainties (Kumar & Linder, 2019).
+I Introduction
+The concordance cosmology predicts consistency among independent probes across all epochs Ostriker:1995su.
+However, while the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) provide precise estimates of the baryon content in the early Universe Cooke:2017cwo, about $30\%$ baryons remain undetected in the local Universe Shull:2011aa.
+These missing baryons are thought to reside in diffuse ionized gas that is too faint to be directly observed, known as the “missing baryon problem” Fukugita:1997bi; cen1999baryons.
+Fast radio bursts (FRBs) are bright, millisecond-duration radio transients of extragalactic origin lorimer2007; thornton2013pop, which can be broadly classified into repeaters and non-repeaters 2016Natur.531..202S; Sun2025a; Sun2025b.
+Their observed dispersion measures (DMs) trace the integrated column density of free electrons along the line of sight (LoS), making them sensitive to baryons in the intergalactic medium (IGM) mcquinn2013locating.
+Thus, FRBs with known redshifts derived from precise host galaxy localization can be used to trace the missing baryons Deng:2013aga; connor2025gas.
+Ref.
+Macquart:2020lln (hereafter Macquart20) first employed $5$ localized FRBs to estimate the present-day cosmic baryon density of $\Omega_{\rm b}=0.051^{+0.021}_{-0.020}$ at the $95\%$ confidence level, which agrees with the joint CMB and BBN result and likely resolves the local baryon problem.
+Ref.
+Yang:2022ftm (hereafter Yang22) used an updated sample of $22$ localized FRBs and obtained $\Omega_{\rm b}=0.0490_{-0.0033}^{+0.0036}$ at the $1\sigma$ level, reinforcing the evidence for a complete cosmic baryon census.
+With the high event rate and increasing number of localized events, FRBs have emerged as a promising cosmological probe Zhou:2014yta; Gao:2014iva; dai2023ska; Zhang:2023gye; see Refs.
+Bhandari:2021thi, Wu:2024iyu, and Glowacki:2024cgu for recent reviews.
+However, FRB cosmology based on DM is hindered by parameter degeneracies, which result in two main sources of systematic uncertainties Kumar:2019qhc.
 The first is from extracting the IGM contribution ($\mathrm{DM}_{\rm IGM}$) from the total DM budget (see Eq.
@@
 This is mainly due to the difficulty in modeling both the host contribution ($\mathrm{DM}_{\rm host}$) and the LoS inhomogeneity in the IGM.
-Macquart20 modeled their probability distributions and Zhang et al.
-(2020, 2021) further employed the IllustrisTNG simulations to derive priors on these empirical DM model parameters for different FRB populations, which effectively reduce biases induced by host galaxy diversity and IGM fluctuations.
+Macquart20 modeled their probability distributions and Refs.
+Zhang:2020mgq; Zhang:2020xoc further employed the IllustrisTNG simulations to derive priors on these empirical DM model parameters for different FRB populations, which effectively reduce biases induced by host galaxy diversity and IGM fluctuations.
 However, the second uncertainty from the analysis of the extracted $\mathrm{DM}_{\rm IGM}$ remains challenging to control.
@@
 (3)).
-Given the limited constraints on $f_{\rm d}$, it is commonly modeled as either a constant or a redshift-dependent function in the literature (Li et al., 2020; Qiang & Wei, 2020; Wang & Wei, 2023; Liu et al., 2025).
-Although $H_{0}$ has been precisely measured, estimates from early- (e.g., Planck 2018 CMB (Aghanim et al., 2020) (hereafter Planck18)) and late-Universe (e.g., SH0ES Cepheid-calibrated Type Ia supernovae (SNe) (Riess et al., 2022)) observations show a significant discrepancy now exceeding $5\sigma$, which has widely been known as the “$H_{0}$ tension” (Guo et al., 2019; Di Valentino et al., 2021; Abdalla et al., 2022; Hu & Wang, 2023).
-This inconsistency would introduce systematic uncertainties into the estimation of $\Omega_{\rm b}$ using FRB samples.
-It should be noted that Macquart20 and Yang22 derived their respective estimates of $\Omega_{\rm b}$ by assuming $H_{0}=70$ km s-1  Mpc-1and a flat prior over 6767676774747474 km s-1  Mpc-1, which provides a reasonable strategy for handling the second uncertainty given the $H_{0}$ tension (note that many studies also constrain $H_{0}$ by adopting early-Universe priors on $\Omega_{\rm b}h^{2}$; see, e.g., Hagstotz et al.
-(2022)).
-Alternatively, jointly analyzing FRBs and datasets that independently measure $H_{0}$ would also be welcome.
-As an emerging cosmological probe, gravitational-wave (GW) standard sirens offer unique advantages for measuring $H_{0}$ (Schutz, 1986).
+Given the limited constraints on $f_{\rm d}$, it is commonly modeled as either a constant or a redshift-dependent function in the literature Li:2020qei; Qiang:2020vta; Wang:2022ami; Liu:2025fdf.
+$\Omega_{\rm b}$ is intrinsically linked to $H_{0}$ through the critical density of the Universe.
+Although $H_{0}$ has been precisely measured, estimates from early- (e.g., Planck 2018 CMB 2020Planck (hereafter Planck18)) and late-Universe (e.g., SH0ES Cepheid-calibrated Type Ia supernovae (SNe) Riess:2021jrx) observations show a significant discrepancy now exceeding $5\sigma$, which has widely been known as the $H_{0}$ tension” Guo:2018ans; DiValentino:2021izs; Abdalla:2022yfr; Hu:2023jqc; CosmoVerseNetwork:2025alb.
+This inconsistency would introduce systematic uncertainties into the baryon census using FRB samples.
+It should be noted that Macquart20 and Yang22 derived their respective estimates of $\Omega_{\rm b}$ by assuming $H_{0}=70$ km s-1  Mpc-1and a flat prior over 67677474 km s-1  Mpc-1, which provides a reasonable strategy for handling the second uncertainty given the $H_{0}$ tension (note that many studies also constrain $H_{0}$ by adopting early-Universe priors on $\Omega_{\rm b}h^{2}$; see, e.g., Ref.
+Hagstotz:2021jzu).
+Nevertheless, any external assumption on $H_{0}$ could bias the baryon estimate, motivating the need for a direct inference of $\Omega_{\rm b}$ that is entirely untethered to the Hubble tension, e.g., jointly analyzing FRBs and datasets that independently measure $H_{0}$ in the late Universe.
+As an emerging cosmological probe, gravitational-wave (GW) standard sirens offer unique advantages for measuring $H_{0}$ Schutz:1986gp.
 This primarily arises from the fact that the amplitudes of GW waveforms are inversely proportional to the luminosity distances ($D_{\rm L}$) to the sources, allowing for a direct measurement of $D_{\rm L}$ through waveform analysis.
-Since $D_{\rm L}\propto 1/H_{0}$, GW standard sirens serve as a powerful and self-calibrating probe for determining $H_{0}$, provided that the redshifts of the GW sources are also known (Dalal et al., 2006; Cutler & Holz, 2009; Zhao et al., 2011; Cai & Yang, 2017; Abbott et al., 2017a, b, c; Chen et al., 2018; Cai et al., 2018; Wang et al., 2018; Zhang et al., 2019a, b; Zhang, 2019; Wang et al., 2020; Zhao et al., 2020b; Gray et al., 2020; Jin et al., 2020; Yu et al., 2020; Wang et al., 2022; Zhu et al., 2022a, b; Wu et al., 2023; Muttoni et al., 2023; Yun et al., 2023; Yu et al., 2024; Han et al., 2024; Feng et al., 2024; Song et al., 2024; Jin et al., 2024a, b; Li et al., 2024; Dong et al., 2024; Xiao et al., 2024; Zhu et al., 2024; Zheng et al., 2024; Han et al., 2025).
-During the first three observing runs of the LIGO, Virgo, and KAGRA (LVK) collaboration, approximately 100 GW events were detected and collected in the third Gravitational-Wave Transient Catalog (GWTC-3).
-Using 47 high–signal-to-noise ratio (SNR) GW standard sirens from GWTC-3, Abbott et al.
-(2023a) obtained a $\sim 10\%$ precision on $H_{0}$.
-Song et al.
-(2025) further combined these 47 GW standard sirens with an additional strong lensing system and achieved a model-independent constraint on $H_{0}$ with a precision of $\sim 8\%$.
+Since $D_{\rm L}\propto 1/H_{0}$, GW standard sirens serve as a powerful and self-calibrating probe for determining $H_{0}$, provided that the redshifts of the GW sources are also known Dalal:2006qt; Cutler:2009qv; Zhao:2010sz; Cai:2016sby; LIGOScientific:2017adf; LIGOScientific:2017vwq; LIGOScientific:2017zic; Chen:2017rfc; Cai:2017aea; Wang:2018lun; Zhang:2019loq; Zhang:2018byx; Zhang:2019ylr; Wang:2019tto; Zhao:2019gyk; Gray:2019ksv; Jin:2020hmc; Yu:2020vyy; Mastrogiovanni:2021wsd; KAGRA:2021duu; Wang:2021srv; Zhu:2021aat; Zhu:2021bpp; Wu:2022dgy; Muttoni:2023prw; Yun:2023ygz; Yu:2023ico; Han:2023exn; Feng:2024lzh; Song:2022siz; Jin:2023sfc; Jin:2023tou; Li:2023gtu; Dong:2024bvw; Xiao:2024nmi; Zhu:2024qpp; Zheng:2024mbo; Han:2025fii; Du:2025odq; Song:2025bio; Jin:2025dvf; Dong:2025ikq.
+During the first three observing runs of the LIGO, Virgo, and KAGRA (LVK) collaboration, approximately 100 GW events were detected and compiled into the GWTC-3 catalog.
+Using 47 GW standard sirens from GWTC-3, Ref.
+LIGOScientific:2021aug achieved a $\sim 10\%$ precision on $H_{0}$, and Ref.
+Song:2025ddm further combined these sirens with an additional strong-lensing system to obtain a model-independent $\sim 8\%$ constraint.
+More recently, with the release of the LVK O4a data set, over 300 GW events have been reported, and by analyzing 142 of them while allowing the population-model parameters of both binary black holes (BBHs) and binary neutron stars (BNSs) to vary freely, the LVK collaboration has obtained a $\sim 14\%$ constraint on $H_{0}$ LIGOScientific:2025jau.
 Therefore, the $H_{0}$ information uniquely provided by GW data enables a direct estimate of $\Omega_{\rm b}$ disentangled from $H_{0}$ in the Macquart relation, which offers new insight into resolving systematic effects in FRB cosmology.
-Unlike previous studies that combined FRBs with traditional observations such as SNe (Jaroszynski, 2019; Gao et al., 2023) or CMB (Zhao et al., 2020a; Qiu et al., 2022), we highlight that combining with GWs inherently involves calibration-free distances and avoids reliance on early-time physics.
-More recently, Zhang & Zhang (2025) and Gao et al.
-(2025b) proposed a novel method based on FRB/persistent radio sources (PRS) associations, which awaits the detection of more such events.
-The FRB + GW synergy is promising, with numerous detections expected from both communities in next-generation surveys (Zhang et al., 2023).
-In this Letter, we combine $104$ localized FRBs with $47$ GW events from the GWTC-3 catalog to jointly constrain $\Omega_{\rm b}$ and $H_{0}$, without relying on any cosmological priors.
-To control the aforementioned systematics, we allow all DM parameters [^1] In this context, we refer to the DM parameters as ($f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$), while the empirical DM model parameters refer specifically to ($F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$).
+Unlike previous studies that combined FRBs with traditional observations such as SNe Jaroszynski:2018vgh; gjz2023 or CMB Zhao:2020ole; Qiu:2021cww, we highlight that combining with GWs inherently involves calibration-free distances and avoids reliance on early-time physics, thereby eliminating the bias induced by the $H_{0}$ tension.
+In this work, we present the first $H_{0}$-free baryon census at low redshift, within a late-Universe probe framework that unifies $104$ localized FRBs and $47$ GW events from the GWTC-3 catalog to jointly constrain $\Omega_{\rm b}$ and $H_{0}$.
+To control model-dependent systematics, we impose no FRB population model and fully marginalize over all DM parameters [^1] In this context, we refer to the DM parameters as ($f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$), while the empirical DM model parameters refer specifically to ($F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$).
 The parameters are defined in Sec.
-2.1.
-to vary freely and incorporate GW data to break parameter degeneracies.
-Note that we do not assume any FRB/GW associations (see Wei et al.
-(2018) and Qiang et al.
-(2025) for studies based on such associations).
-2 Data and methodology
-2.1 Localized FRB data
+II.1., while on the GW side we adopt a population model for the main analysis and additionally release its parameters to assess robustness.
+Note that we do not assume any FRB/GW associations (see Refs.
+Wei:2018cgd and Qiang:2024erm for studies based on such associations).
+This work is organized as follows.
+In Sec.
+II, we describe the datasets and likelihood construction for both FRBs and GWs.
+Results and related discussions are presented in Sec.
+III, and we conclude in Sec.
+IV.
+II Data and methodology
+II.1 Localized FRB data
 In this section, we first present the likelihood function for localized FRBs based on DM analysis following Macquart20, then the FRB sample used in this work.
@@
 $P_{\rm{halo}}(\mathrm{DM}_{\rm{halo}})=\frac{1}{\sqrt{2\pi}\sigma_{\rm{halo}}}\exp\left[-\frac{(\mathrm{DM}_{\rm{halo}}-\mu_{\rm{halo}})^{2}}{2\sigma_{\rm{halo}}^{2}}\right].$ (2)
-The average $\mathrm{DM}_{\mathrm{IGM}}$ in the standard $\Lambda$CDM (cosmological constant $\Lambda$ + cold dark matter) cosmology is given by the Macquart relation,
+[…]
 $\langle\mathrm{DM}_{\mathrm{IGM}}\rangle=$ $\frac{3cf_{\rm e}}{8\pi Gm_{\mathrm{p}}}\Omega_{\mathrm{b}}H_{0}f_{\mathrm{d}}\int_{0}^{z}\frac{(1+z)dz}{\sqrt{\Omega_{\rm m}(1+z)^{3}+(1-\Omega_{\rm m})}},$ (3)
@@
 Here, $A$ is a normalization constant, and $C_{0}$ is chosen to ensure that the mean of this distribution is unity.
-Note that we include the factor $1/\langle{\rm DM_{\rm IGM}}\rangle$ suggested by (Zhang et al., 2025b) when normalizing Eq.
+Note that we include the factor $1/\langle{\rm DM_{\rm IGM}}\rangle$ suggested by Ref.
+Zhang:2025wif when normalizing Eq.
 (4).
@@
 $P_{\rm{host}}(\rm{DM}_{\rm{host}})$ $=\frac{1}{\sqrt{2\pi}\sigma_{\rm{host}}\mathrm{DM}_{\rm{host}}}$
-×$\times\exp\left[-\frac{(\ln\mathrm{DM}_{\rm{host}}-\mu_{\rm{host}})^{2}}{2\sigma_{\rm{host}}^{2}}\right],$ (5)
+$\times\exp\left[-\frac{(\ln\mathrm{DM}_{\rm{host}}-\mu_{\rm{host}})^{2}}{2\sigma_{\rm{host}}^{2}}\right],$ (5)
 where a correction $\mathrm{DM}_{\text{host }}\rightarrow\mathrm{DM}_{\text{host}}/(1+z)$ has been applied.
@@
 $\mathcal{L}_{\mathrm{FRB}}(\{x_{\rm FRB}\})=\sum_{i=1}^{N_{\mathrm{FRB}}}P_{i}\left(\mathrm{DM}^{\prime}_{i}\right).$ (6)
-For the $i$-th FRB, the possibility distribution of its extragalactic DM component ($\mathrm{DM}^{\prime}=\rm DM_{\rm IGM}+\rm DM_{\rm host}$) is computed by convolving Eqs. (2), (4), and (2.1):
+For the $i$-th FRB, the possibility distribution of its extragalactic DM component ($\mathrm{DM}^{\prime}=\rm DM_{\rm IGM}+\rm DM_{\rm host}$) is computed by convolving Eqs. (2), (4), and (II.1):
 $P_{i}\left(\mathrm{DM}^{\prime}_{i}\right)$ $=\int_{30\,\mathrm{pc\,cm}^{-3}}^{100\,\mathrm{pc\,cm}^{-3}}P_{\rm halo}\left(\mathrm{DM}_{\rm halo}\right)\,\mathrm{dDM}_{\rm halo}$
-×$\times\int_{0}^{\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}}P_{\rm host}\left(\mathrm{DM}_{\rm host}\right)$
-×$\times P_{\rm IGM}\left(\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}-\mathrm{DM}_{\rm host}\right)\,\mathrm{dDM}_{\rm host}.$ (7)
+$\times\int_{0}^{\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}}P_{\rm host}\left(\mathrm{DM}_{\rm host}\right)$
+$\times P_{\rm IGM}\left(\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}-\mathrm{DM}_{\rm host}\right)\,\mathrm{dDM}_{\rm host}.$ (7)
 We also consider a scenario in which the empirical DM model parameters are fixed, though our main results treat them as free.
-Specifically, we adopt values calibrated from cosmological hydrodynamical simulations, notably the IllustrisTNG simulation, which provides redshift-dependent distributions for both $P_{\rm IGM}(\mathrm{DM}_{\rm IGM})$ (Zhang et al., 2021) and $P_{\rm host}(\mathrm{DM}_{\rm host})$ (Zhang et al., 2020).
+Specifically, we adopt values calibrated from cosmological hydrodynamical simulations, notably the IllustrisTNG simulation, which provides redshift-dependent distributions for both $P_{\rm IGM}(\mathrm{DM}_{\rm IGM})$ Zhang:2020xoc and $P_{\rm host}(\mathrm{DM}_{\rm host})$ Zhang:2020mgq.
 The parameters $A(z)$, $C_{0}(z)$, and $\sigma_{\rm IGM}(z)$ in Eq.
 (4), as well as $\mu_{\rm host}(z)$ and $\sigma_{\rm host}(z)$ in Eq.
-(2.1), are interpolated from best-fit simulation results across redshift.
-Note that $\mu_{\rm host}$ and $\sigma_{\rm host}$ are further fixed based on the host galaxy types (see the seventh column of Table 2).
-Finally, we select a sample of $104$ localized FRBs from an initial set of $119$, based on the following criteria: (1) $\mathrm{DM}_{\mathrm{MW,ISM}}$ is required to be less than $40\%$ of $\mathrm{DM}_{\rm obs}$ (Connor et al., 2025); (2) $\mathrm{DM}_{\rm obs}-\rm DM_{MW,ISM}$ is taken to exceed 80 $\,\rm{pc}\,\rm{cm}^{-3}\,$, ensuring a positive $\mathrm{DM}_{\rm host}$; (3) FRB 20190520B and FRB 20220831A are excluded due to excess local DM contributions from their host environments.
-The final sample is listed in Appendix B (see Table 2).
-2.2 GW standard siren data
-We use 47 GW standard sirens with SNRs greater than 11 and an inverse false alarm rate exceeding 4 yr from GWTC-3, covering a redshift range of approximately $z<0.8$.
-The sample consists of 42 BBHs, three neutron star-black holes (NSBHs), and two binary neutron stars (BNSs).
-We obtain the redshift of GW170817 from the electromagnetic (EM) counterpart (Abbott et al., 2017a, b, c).
-For GW events without EM counterparts, we obtain their redshifts through the dark siren method (Schutz, 1986; Mastrogiovanni et al., 2023; Gray et al., 2023) and incorporate the GLADE+ catalog (Dálya et al., 2022).
-In the following, we provide a brief introduction to the likelihood function of GW standard sirens.
-We denote the source parameters of each GW event by $\bm{\theta}$, and the population-level parameters by $\bm{\Lambda}$, including population model parameters and cosmological parameters.
-For GW events without EM counterparts, assuming a “scale-free” prior on the expected number of detected events $N_{\exp}$, i.e., $\pi(N_{\exp})\propto 1/N_{\exp}$, the likelihood function of the GW data set $\{x_{\rm GW}\}$ composed of $N_{\rm GW}$ GW events (Mastrogiovanni et al., 2021; Abbott et al., 2023b), is given by
-$\mathcal{L}_{\rm GW}(\{x_{\rm GW}\}|\bm{\Lambda})\propto\prod_{i}^{N_{\rm GW}}\frac{\int{\rm d}\bm{\theta}\mathcal{L}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})\frac{{\rm d}N_{\rm CBC}(\bm{\Lambda})}{{\rm d}\bm{\theta}{\rm d}t}}{\int{\rm d}\bm{\theta}P_{\rm det}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})\frac{{\rm d}N_{\rm CBC}(\bm{\Lambda})}{{\rm d}\bm{\theta}{\rm d}t}},$ (8)
-where $\mathcal{L}_{\rm GW}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})$ is the parameter estimation likelihood, quantifying the uncertainties in measuring the GW source parameters $\bm{\theta}$ through waveform analysis.
-${\rm d}N_{\rm CBC}(\bm{\Lambda})/{\rm d}\bm{\theta}{\rm d}t$ is the CBC detection rate in the detector frame, which includes population models of GW sources, cosmological models, and redshift information from the GLADE+ galaxy catalog.
-In assessing the completeness of the GLADE+ catalog, we adopt a Schechter function to describe the $B$-band absolute magnitude distribution of galaxies (Kochanek et al., 2001).
-We further assume that the probability of a galaxy being the host of the GW source is proportional to its $B$-band luminosity.
-$P_{\rm det}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})$ is the detection probability of a GW event with parameters of $\bm{\theta}$ and $\bm{\Lambda}$.
-The denominator of Eq.
-(8) accounts for the GW selection effect.
-For GW dark sirens, population models can help infer redshift for the sources, especially in scenarios where the galaxy catalog is significantly incomplete (Taylor et al., 2012; Ezquiaga & Holz, 2022; Mastrogiovanni et al., 2023; Gray et al., 2023).
-We adopt the same population model assumptions as Abbott et al.
-(2023a), but restrict our analysis to the Power Law + Peak BBH mass model, which is best supported by the data.
+(II.1), are interpolated from best-fit simulation results across redshift.
+Note that $\mu_{\rm host}$ and $\sigma_{\rm host}$ are further fixed based on the host galaxy types (see the seventh column of Table LABEL:tab:localized_FRBs).
+Finally, we select a sample of $104$ localized FRBs from an initial set of $119$, based on the following criteria: (1) $\mathrm{DM}_{\mathrm{MW,ISM}}$ is required to be less than $40\%$ of $\mathrm{DM}_{\rm obs}$ connor2025gas; (2) $\mathrm{DM}_{\rm obs}-\rm DM_{MW,ISM}$ is taken to exceed 80 $\rm{pc}\,\rm{cm}^{-3}\,$, ensuring a positive $\mathrm{DM}_{\rm host}$; (3) FRB 20190520B and FRB 20220831A are excluded due to excess local DM contributions from their host environments.
+The final sample is listed in Appendix B (see Tables LABEL:tab:localized_FRBs).
+II.2 GW standard siren data
+We use 47 GW standard sirens with SNRs greater than 11 and an inverse false alarm rate exceeding 4 yr from GWTC-3222https://gwosc.org/GWTC-3, covering a redshift range of approximately $z<0.8$.
+The sample consists of 42 BBHs, three neutron star-black holes (NSBHs), and two BNSs.
+We do not consider GW events detected in the LVK O4a run, mainly because during this run, Virgo is not online, and all GW events are detected by two LIGOs, resulting in relatively poorer sky localizations and luminosity distance measurements LIGOScientific:2025jau.
+We obtain the redshift of GW170817 from the electromagnetic (EM) counterpart LIGOScientific:2017adf; LIGOScientific:2017vwq; LIGOScientific:2017zic.
+For GW events without EM counterparts, we infer their redshifts through the dark siren method Schutz:1986gp; Mastrogiovanni:2023emh; Gray:2023wgj and incorporate the GLADE+ catalog Dalya:2021ewn.
+In the dark siren method, assuming the GW sources all originate from galaxies, we can cross-match sky localizations of GW events with galaxies collected in galaxy catalogs, and identify a serious potential host galaxy.
+These potential host galaxies provide redshift priors for GW events, which help to infer cosmological parameters in a hierarchical Bayesian framework.
+Furthermore, GW population models can help break degeneracies of masses and redshifts in GW detections, thereby providing additional redshift information Taylor:2011fs; Ezquiaga:2022zkx; Mastrogiovanni:2023emh; Gray:2023wgj.
+In this paper, we use the public pipeline ICAROGW Mastrogiovanni:2023zbw to compute the GW likelihood function, denoted as $\mathcal{L}_{\rm GW}$.
+The explicit formulas can be found in Refs Mastrogiovanni:2023emh; Gray:2023wgj.
+We adopt the same population model assumptions as Ref.
+LIGOScientific:2021aug, but restrict our analysis to the Power Law + Peak BBH mass model, which is best supported by the GWTC-3 data.
 Considering additional models would increase computational cost and lies beyond the scope of this work.
@@
 The redshift distribution of GW sources is modeled using a phenomenological prescription that traces the cosmic star formation rate.
-Consistent with Abbott et al.
-(2023a), we disregard potential redshift evolution in the BBH mass distribution as well as variations in the GW detection rate arising from the spin distribution.
+Consistent with Ref.
+LIGOScientific:2021aug, we disregard potential redshift evolution in the BBH mass distribution as well as variations in the GW detection rate arising from the spin distribution.
 These effects are currently subdominant compared to the statistical uncertainties present in the existing GW observations.
-In cases where EM counterparts are detected, the information from EM observations should be incorporated alongside GW data.
-Since GW and EM detections are statistically independent, the overall likelihood is given by
-$\mathcal{L}_{\rm EM+GW}(x_{{\rm GW},i}|\bm{\theta})\propto\mathcal{L}_{\rm EM}(x_{{\rm GW},i}|z,\Omega)\mathcal{L}_{\rm GW}(x_{{\rm GW},i}|\bm{\theta}),$ (9)
-where $\mathcal{L}_{\rm EM}(x_{{\rm GW},i}|z,\Omega)$ quantifies the measurement uncertainties of the redshift $z$ and the sky localization $\Omega$ from EM observations.
-We assume that the probability of having EM observations is 1 once an EM counterpart is present (Gray et al., 2020; Mastrogiovanni et al., 2024).
-This assumption holds only within the current GW detection range, as it remains much smaller than that of EM observations.
-2.3 Joint analysis of FRB and GW data
+II.3 Joint analysis of FRB and GW data
 Figure 1: Normalized posterior distributions of $\Omega_{\rm b}$~from 104 FRBs (orange) and 104 FRBs + 47 GWs (indigo).
 The median and $1\sigma$ credible interval from FRB + GW is $\Omega_{\rm b}=0.0488\pm 0.0064$, marginalized over $f_{\rm d}$ and empirical DM model parameters.
-The errorbars in different colors and linestyles are shown under different assumptions: FRB + $H_{0}$ (Planck18 prior) and FRB + $H_{0}$ (SH0ES prior) adopt Gaussian priors on $H_{0}$ from Planck18 (Aghanim et al., 2020) and SH0ES (Riess et al., 2022), respectively; FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW adopt priors on empirical DM model parameters from the IllustrisTNG simulation (Zhang et al., 2020, 2021) and on $f_{\rm d}$ from the redshift-dependent form of Macquart20 (Macquart et al., 2020), respectively.
-The blue, gray, and red shaded bands show the results from CMB + BBN (Cooke et al., 2018), Macquart20, and Yang22 (Yang et al., 2022), respectively.
-Note that Yang22 adopts the TNG prior, a flat prior on $H_{0}$in $\mathcal{U}(67,73)$ km s-1 Mpc-1, and $f_{\rm d}=0.82$.
+The errorbars in different colors and linestyles are shown under different assumptions: FRB + $H_{0}$ (Planck18 prior) and FRB + $H_{0}$ (SH0ES prior) adopt Gaussian priors on $H_{0}$ from Planck18 2020Planck and SH0ES Riess:2021jrx, respectively; FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW adopt priors on empirical DM model parameters from the IllustrisTNG simulation Zhang:2020mgq; Zhang:2020xoc and on $f_{\rm d}$ from the redshift-dependent form of Macquart20 Macquart:2020lln, respectively.
+The blue, gray, and red shaded bands show the results from CMB + BBN Cooke:2017cwo, Macquart20, and Yang22 Yang:2022ftm, respectively.
+Note that Yang22 adopts the TNG prior, a flat prior on $H_{0}$ $\in$ $\mathcal{U}(67,73)$ km s-1 Mpc-1, and $f_{\rm d}=0.82$.
 All results are at the $1\sigma$ level.
-We jointly constrain cosmological and FRB-related DM parameters, and adopt flat priors for the following seven parameters: $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ (Yang et al., 2022), $\Omega_{\rm b}\in\mathcal{U}(0.01,0.10)$, $H_{0}\in\mathcal{U}(20,140)~{}\rm km\ s^{-1}\ Mpc^{-1}$, $f_{\rm d}\in\mathcal{U}(0.84,0.96)$, [^2] This prior on $f_{\rm d}$ is informed by observations of the baryon fraction in stars ($f_{*}$) and cold gas ($f_{\rm cold}$), independent of FRBs.
-Assuming a conservative flat prior for $f_{*}$ extended by a factor of five, the diffuse baryon fraction is inferred via $f_{\rm d}=1-f_{*}-f_{\rm cold}$ (Connor et al., 2025).
-$F\in\mathcal{U}(0.011,1.0)$, $\mu_{\rm host}\in\mathcal{U}(3.0,5.3)$, and $\sigma_{\rm host}\in\mathcal{U}(0.2,2.0)$ (Macquart et al., 2020).
-The GW population model parameters include $\alpha$, $\beta$, $m_{\rm min}$, $m_{\rm max}$, $\Delta m$, $\mu_{\rm g}$, $\sigma_{\rm g}$, $\lambda_{\rm peak}$, $\gamma$, $\kappa$, and $z_{\rm p}$, which are fixed to the inferred values of Abbott et al.
-(2023a).
-We utilize the bilby [^3] https://github.com/bilby-dev/bilby package to perform Markov Chain Monto Carlo (MCMC) analysis and choose the emcee [^4] https://emcee.readthedocs.io/en/stable sampler.
-We use the ICAROGW [^5] https://github.com/simone-mastrogiovanni/icarogw package to do the hierarchic Bayesian analysis of GW standard sirens, and the GetDist [^6] https://github.com/cmbant/getdist package to plot the marginalized posterior distributions.
-The final likelihood $\mathcal{L}$ used in our analysis is the combination of $\mathcal{L}_{\mathrm{FRB}}$ and $\mathcal{L}_{\rm GW}$ in Eqs.
-(6) and (8), respectively, which is written as
+We jointly constrain cosmological and FRB-related DM parameters, and adopt flat priors for the following seven parameters: $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ Yang:2022ftm, $\Omega_{\rm b}\in\mathcal{U}(0.01,0.10)$, $H_{0}\in\mathcal{U}(20,140)~\rm km\ s^{-1}\ Mpc^{-1}$, $f_{\rm d}\in\mathcal{U}(0.84,0.96)$, [^3] This prior on $f_{\rm d}$ is informed by observations of the baryon fraction in stars ($f_{*}$) and cold gas ($f_{\rm cold}$), independent of FRBs.
+Assuming a conservative flat prior for $f_{*}$ extended by a factor of five, the diffuse baryon fraction is inferred via $f_{\rm d}=1-f_{*}-f_{\rm cold}$ connor2025gas.
+$F\in\mathcal{U}(0.011,1.0)$, $\mu_{\rm host}\in\mathcal{U}(3.0,5.3)$, and $\sigma_{\rm host}\in\mathcal{U}(0.2,2.0)$ Macquart:2020lln.
+For the GW population model parameters, we consider two scenarios.
+In the first, following Ref.
+LIGOScientific:2021aug, we fix all population-model parameters to their fiducial values.
+In the second, we allow the BBH mass-distribution parameters and the redshift-distribution parameters of GW sources to vary freely, adopting the same prior ranges as the spectral siren analysis in Ref.
+LIGOScientific:2021aug, to assess the impact of population-model uncertainties on our results.
+This treatment is similar to Ref.
+Gray:2023wgj; however, unlike Ref.
+LIGOScientific:2025jau, we do not free the mass-distribution parameters of neutron-star sources.
+We utilize the bilby [^4] https://github.com/bilby-dev/bilby package to perform Markov Chain Monte Carlo (MCMC) analysis and choose the emcee [^5] https://emcee.readthedocs.io/en/stable sampler.
+The GetDist [^6] https://github.com/cmbant/getdist package to plot the marginalized posterior distributions.
+The final likelihood $\mathcal{L}$ used in our analysis is the combination of $\mathcal{L}_{\mathrm{FRB}}$ and $\mathcal{L}_{\rm GW}$, which is written as
 $\ln\mathcal{L}=\ln\left(\mathcal{L}_{\mathrm{FRB}}\right)+\ln\left(\mathcal{L}_{\mathrm{GW}}\right).$ (8)
 Table 1: $1\sigma$ constraints on cosmological parameters $\Omega_{\rm b}$ and $H_{0}$.
+The results are derived from $104$ localized FRBs alone or in combination with $47$ GW events, under different assumptions.
+For all cases, the $1\sigma$ constraints on DM parameters are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$.
+Unless otherwise specified, the FRB + GW case assumes a flat prior of $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ and fixed GW population parameters.
+Note that $H_{0}$ is in units of km s-1  Mpc-1.
     Data     $\Omega_{\rm b}$     $H_{0}$
@@
     FRB + GW     $0.0488\pm 0.0064$     $71.6^{+4.4}_{-8.6}$
+    FRB + GW ($\Omega_{\rm m}\in\mathcal{U}(0,1)$)     $0.0454^{+0.0066}_{-0.0082}$     $72.7^{+4.7}_{-9.0}$
     FRB (TNG) + GW     $0.0491\pm 0.0054$     $71.2^{+4.4}_{-8.0}$
@@
     FRB + $H_{0}$ (SH0ES prior)     $0.0473\pm 0.0043$     $73.0\pm 1.0$
-3 Results and discussions
-Figure 2: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm b}$ and $\Omega_{\rm b}h_{70}$ versus $H_{0}$, $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$, using FRB (orange) and FRB + GW (purple) data.
-The correlation coefficient $r$ is labeled in each panel to quantify parameter degeneracies.
-In this section, we report and discuss the constraints on $\Omega_{\rm b}~{}$in Table 1.
-The posteriors for $\Omega_{\rm b}~{}$and the corresponding credible intervals are shown in Figure 1.
+    FRB + GW (GW parameters free)     $0.0448^{+0.0094}_{-0.0076}$     $79.9^{+5.0}_{-17.0}$
+III Results and discussions
+In this section, we first report the constraints on the baryon density as our main results in Sec.
+III.1, and then extend the analysis to assess the impacts of FRB and GW population modeling assumptions in Secs.
+III.2 and III.3, respectively.
+III.1 Main results
+First, we report the main results for the baryon-density estimate in the first block of Table 1.
+The posteriors for $\Omega_{\rm b}$~and the corresponding credible intervals are shown in Fig.
+1.
 Using FRB data alone yields a relatively weak constraint of $\Omega_{\mathrm{b}}=0.052_{-0.028}^{+0.012}$ at the $1\sigma$ level.
@@
 We compare our $\Omega_{\rm b}$~results with those from previous works.
-Macquart20 and Yang22 constrained $\Omega_{\rm b}~{}$using different samples of localized FRBs under external assumptions of $H_{0}=70$ km s-1  Mpc-1and a flat prior on $H_{0}\in\mathcal{U}(67,73)$ km s-1  Mpc-1, as shown by the gray and red bands in Figure 1, respectively.
+Macquart20 and Yang22 constrained $\Omega_{\rm b}$~using different samples of localized FRBs under external assumptions of $H_{0}=70$ km s-1  Mpc-1and a flat prior on $H_{0}\in\mathcal{U}(67,73)$ km s-1  Mpc-1, as shown by the gray and red bands in Fig.
+1, respectively.
 The constraint from Yang22 compares favorably to ours, since the current GW standard siren sample remains too limited to outperform their adopted $H_{0}$ prior.
-Applying the external $H_{0}$ priors from Planck18 or SH0ES to our FRB sample alone yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively; the corresponding error bars are shown in Figure 1.
+Applying the external $H_{0}$ priors from Planck18 or SH0ES to our FRB sample alone yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively; the corresponding error bars are shown in Fig.
+1.
 This suggests that the $H_{0}$ tension would lead to two biased $\Omega_{\rm b}$~estimates.
-Although the induced “$\Omega_{\rm b}$ inconsistency” in our sample is within $1\sigma$ uncertainty ($\sim 0.95\sigma$), it may become non-negligible with larger samples in updated DM analyses.
+Although the induced $\Omega_{\rm b}$ inconsistency” in our sample is within $1\sigma$ uncertainty ($\sim 0.95\sigma$), it may become non-negligible with larger samples in updated DM analyses.
 Nevertheless, the joint FRB + GW result assumes no prior on $H_{0}$ and instead provides a late-Universe measurement of baryons, independent of both CMB and local distance ladder.
-This FRB + GW synergy may establish a novel paradigm for the cosmic baryon census, analogous to CMB + BBN, and serve as a powerful low-redshift cosmological probe in the era of next-generation FRB and GW surveys (Zhang et al., 2023).
-We also investigate potential improvements to our results.
+This FRB + GW synergy may establish a novel paradigm for the cosmic baryon census, analogous to CMB + BBN, and serve as a powerful low-redshift cosmological probe in the era of next-generation FRB and GW surveys Zhang:2023gye.
+We also assess the dependence on the prior for $\Omega_{\rm m}$.
+Relaxing the prior from $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ to a broad flat prior $\Omega_{\rm m}\in\mathcal{U}(0,1)$ yields $\Omega_{\rm b}=0.0454^{+0.0066}_{-0.0082}$ and $H_{0}=72.7^{+4.7}_{-9.0}$ km s-1  Mpc-1, as shown in the third row of Table 1.
+Compared with the fiducial case ($\Omega_{\rm b}=0.0488\pm 0.0064$), the slightly lower value of $\Omega_{\rm b}$ is driven by the biased constraint on $\Omega_{\rm m}$ under the broad prior, $\Omega_{\rm m}=0.225^{+0.060}_{-0.180}$, which propagates through the positive correlation between $\Omega_{\rm b}$~and $\Omega_{\rm m}$.
+Therefore, we caveat that our present analysis moderately depends on the $\Omega_{\rm m}$ prior, which can be improved with larger future FRB and GW samples to better constrain $\Omega_{\rm m}$.
+III.2 Impact of FRB model parameters
+Figure 2: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm b}$ and $\Omega_{\rm b}h_{70}$ versus $H_{0}$, $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$, using FRB (orange) and FRB + GW (indigo) data.
+The correlation coefficient $r$ is labeled in each panel to quantify parameter degeneracies.
+Note that $H_{0}$ is in units of km s-1  Mpc-1.
+Next, we assess the impact of FRB population modeling on the baryon census, as listed in the second block of Table 1.
+In our fiducial analysis in Sec.
+III.1, no specific FRB population model is imposed and all DM parameters are left free to avoid potential bias.
+To explore how such modeling affects the $\Omega_{\rm b}$ inference, we apply several commonly used FRB population models and evaluate their potential to improve the precision of $\Omega_{\rm b}$.
 FRB cosmology is subject to uncontrolled systematics mainly from modeling $\mathrm{DM}_{\mathrm{IGM}}$ and $\mathrm{DM}_{\mathrm{host}}$ distributions, as well as the weakly constrained $f_{\rm d}$, which affects the mean value of $\mathrm{DM}_{\mathrm{IGM}}$.
 To reduce these systematics, cosmological simulations are commonly used to calibrate these DM parameters ($F$, $\mu_{\rm host}$, $\sigma_{\rm host}$, and $f_{\rm d}$) at different redshifts.
-Accordingly, we consider two cases with simulation-based priors: (i) fixing the empirical DM model parameters to the best-fit results from the IllustrisTNG simulation (hereafter TNG) (Zhang et al., 2020, 2021); and (ii) further fixing $f_{\rm d}$ to a redshift-dependent form following Macquart20, as implemented in the public FRB code [^7] https://github.com/FRBs/FRB.
-The constraints from FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW are $\Omega_{\mathrm{b}}=0.0491\pm 0.0054$ and $\Omega_{\mathrm{b}}=0.0507\pm 0.0052$, respectively, with the corresponding errorbars shown in Figure 1.
+Accordingly, we consider two cases with simulation-based priors: (i) fixing the empirical DM model parameters to the best-fit results from the IllustrisTNG simulation (hereafter TNG) Zhang:2020mgq; Zhang:2020xoc; and (ii) further fixing $f_{\rm d}$ to a redshift-dependent form following Macquart20, as implemented in the public FRB code [^7] https://github.com/FRBs/FRB.
+The constraints from FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW are $\Omega_{\mathrm{b}}=0.0491\pm 0.0054$ and $\Omega_{\mathrm{b}}=0.0507\pm 0.0052$, respectively, with the corresponding errorbars shown in Fig.
+1.
 Both results are consistent with the FRB + GW constraint within the $1\sigma$ uncertainty, suggesting that the simulation-based priors can improve current constraints without introducing significant bias.
@@
 To assess the impact of systematic uncertainties on $\Omega_{\mathrm{b}}$ and verify the robustness of our constraint, we analyze the parameter correlations using the FRB + GW dataset.
-The lower row of Figure 2 shows the marginalized posteriors of $\Omega_{\rm b}~{}$and other relevant parameters, with correlation coefficients $r$ labeled in each panel to quantify parameter degeneracies.
-On the one hand, the degeneracy between $\Omega_{\rm b}$ and $H_{0}$ is the strongest among all parameter pairs ($r=-0.73$), implying that the inferred value of $\Omega_{\rm b}$ is highly sensitive to how $H_{0}$ is determined.
+The lower row of Fig.
+2 shows the marginalized posteriors of $\Omega_{\rm b}$~and other relevant parameters, with correlation coefficients $r$ labeled in each panel to quantify parameter degeneracies.
+On the one hand, the degeneracy between $\Omega_{\rm b}$ and $H_{0}$ is the strongest among all parameter pairs (r=0.73r=-0.73), implying that the inferred value of $\Omega_{\rm b}$ is highly sensitive to how $H_{0}$ is determined.
 While FRB + $H_{0}$ analyses rely on external priors, which are controversial due to the $H_{0}$ tension, FRB + GW analyses rely on GW data to provide $H_{0}$, which is uniquely derived from self-calibrating absolute distance measurements.
@@
 On the other hand, the correlations between $\Omega_{\rm b}$~and DM parameters are moderate.
-These DM parameters are not properly constrained, both in this work (see Appendix A for details) and earlier studies (James et al., 2022), and thus act as nuisance parameters that bias the inferred $\Omega_{\rm b}$.
-In the joint FRB + GW analysis, the correlation coefficients between $\Omega_{\rm b}$ and $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$ are $-0.28$, $0.40$, $-0.40$, and $0.17$, respectively.
-For comparison, the upper row of Figure 2 shows the constraint on $\Omega_{\rm b}h_{70}$, where $h_{70}=H_{0}/(70~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}})$.
-This parameter absorbs the strong degeneracy, enabling it to be effectively constrained using FRB data alone (Macquart et al., 2020; Connor et al., 2025).
-In this case, the corresponding correlations are notably stronger: $-0.41$, $0.57$, $-0.62$, and $0.27$, respectively.
+These DM parameters are not properly constrained, both in this work (see Appendix A for details) and earlier studies James:2022dcx, and thus act as nuisance parameters that bias the inferred $\Omega_{\rm b}$.
+In the joint FRB + GW analysis, the correlation coefficients between $\Omega_{\rm b}$ and $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$ are 0.28-0.28, $0.40$, 0.40-0.40, and $0.17$, respectively.
+For comparison, the upper row of Fig.
+2 shows the constraint on $\Omega_{\rm b}h_{70}$, where $h_{70}=H_{0}/(70~\mathrm{km~s^{-1}~Mpc^{-1}})$.
+This parameter absorbs the strong degeneracy, enabling it to be effectively constrained using FRB data alone Macquart:2020lln; connor2025gas.
+In this case, the corresponding correlations are notably stronger: 0.41-0.41, $0.57$, 0.62-0.62, and $0.27$, respectively.
 This comparison suggests that directly measuring $\Omega_{\rm b}$ reduces degeneracies with the nuisance parameters and is more robust against the associated systematic uncertainties than $\Omega_{\rm b}h_{70}$.
-We conclude that incorporating GW data into the analysis of FRB dispersion measures enables a direct and effective measurement of $\Omega_{\rm b}$ and, more importantly, avoids biases from the $H_{0}$ tension and mitigates those associated with nuisance model parameters.
+We conclude that the impact of FRB population modeling on the inferred $\Omega_{\rm b}$ is negligible within our analysis.
+The results also show that, within our FRB + GW framework, incorporating GW data into the analysis of FRB dispersion measures enables a direct and effective measurement of $\Omega_{\rm b}$ and, more importantly, avoids biases from the $H_{0}$ tension and mitigates those associated with nuisance model parameters.
 This demonstrates that the synergy between FRBs and GWs provides a robust estimate of the cosmic baryon density.
-4 Conclusion
-In this Letter, we combine two emerging cosmological probes, i.e., $104$ localized FRBs and $47$ GW standard sirens, to measure the cosmic baryon density in a robust manner.
-DM-based FRB cosmology struggles with various parameter degeneracies, particularly between $H_{0}$ and $\Omega_{\rm b}$.
-GWs can independently infer $H_{0}$ from self-calibrated luminosity distances obtained through waveform analysis as sirens, thus breaking the $\Omega_{\rm b}$-$H_{0}$ degeneracy.
+III.3 Impact of GW model parameters
+Figure 3: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm m}$, $H_{0}$, and the typical GW population parameter $\mu_{g}$ using FRB + GW data.
+Two cases are shown: GW parameters fixed within the PowerLaw + Peak model (indigo) and the parameters allowed to vary freely and marginalised over (gray).
+The dotted lines indicate the median values of each posterior, except for $\mu_{g}$ in the GW-parameter-fixed case, where the dotted line marks the prior value $\mu_{g}=32.27$ adopted following Ref.
+LIGOScientific:2021aug.
+Note that $H_{0}$ is in units of km s-1  Mpc-1.
+Last, we assess the impact of GW population modeling on our baryon census, as listed in the third block of Table 1.
+In our fiducial analysis in Sec.
+III.1, no FRB population model is imposed, whereas the PowerLaw + Peak GW population model is adopted for the GW population, as is commonly used in GW-based cosmological inference LIGOScientific:2021aug.
+To investigate the robustness of our $\Omega_{\rm b}$~estimate against this assumption, we release the GW population model parameters and marginalize over them.
+We find that joint FRB + GW analysis with GW model parameters free yields a higher value of $H_{0}=79.9^{+5.0}_{-17.0}$ km s-1  Mpc-1and a correspondingly lower value of $\Omega_{\rm b}=0.0448^{+0.0094}_{-0.0076}$.
+The $\Omega_{\rm b}$ result ($\sim 20\%$ precision) remains consistent with the CMB + BBN determination at the $1\sigma$ level.
+To illustrate how the GW population prior affects $\Omega_{\rm b}$, we focus on the parameter $\mu_{g}$, which denotes the mean of the Gaussian component in the primary mass distribution within the PowerLaw + Peak model.
+This parameter is moderately correlated with the inferred $H_{0}$ from GW observations LIGOScientific:2021aug, making it suitable for accounting for the impact of population modeling.
+As shown in Fig.
+3, $\mu_{g}$ exhibits a negative degeneracy with $H_{0}$ and a corresponding positive degeneracy with $\Omega_{\rm b}$.
+Consequently, compared with the prior value $\mu_{g}=32.27$ adopted in the GW–parameter-fixed (indigo) case, the lower inferred value of $\mu_{g}=31.5$ in the GW–parameter-free (gray) case leads to a higher $H_{0}$ and thus a lower $\Omega_{\rm b}$.
+Thus, within our FRB + GW framework, the inferred $\Omega_{\rm b}$ is more sensitive to GW-related parameters than to FRB-related ones (see Sec.
+III.2).
+It is because $H_{0}$, the essential parameter to infer $\Omega_{\rm b}$ in our framework, is more strongly degenerate with the GW population parameters than with the FRB-related parameters.
+We conclude that, to obtain an unbiased and robust estimate of $\Omega_{\rm b}$, caution is required when adopting GW-population assumptions, due to their strong impact on the inferred $H_{0}$.
+Nevertheless, our test without imposing any GW population model reinforces that our fiducial result obtained with the PowerLaw + Peak model is reliable and not significantly biased.
+IV Conclusion
+In this paper, we combine two emerging cosmological probes, i.e., $104$ localized FRBs and $47$ GW standard sirens, within a unified framework to measure the cosmic baryon density in a $H_{0}$-free approach.
+Most previous baryon censuses assumed a fixed value of $H_{0}$ or adopted an external prior, which could bias the inference in the presence of the Hubble tension.
+If FRB DMs are instead used to trace the baryons, additional parameter degeneracies arise in the DM modeling, particularly between $H_{0}$ and $\Omega_{\rm b}$.
+GWs independently infer $H_{0}$ from self-calibrated luminosity distances obtained through waveform analysis as sirens, thus breaking the $\Omega_{\rm b}$-$H_{0}$ degeneracy.
 This joint analysis yields a constraint of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), which is remarkably consistent with early-Universe values from BBN and CMB at the $\sim 13\%$ level and provides a purely late-Universe perspective on resolving the missing baryon problem.
@@
 Moreover, we explore two prior-based scenarios: (i) $H_{0}$-prior-based cases using FRB data alone, where adopting Gaussian priors from Planck18 and SH0ES yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively.
-These results indicate that the $H_{0}$ tension would bias baryon density estimates, although the induced “$\Omega_{\rm b}~{}$inconsistency” remains within $1\sigma$ uncertainty in our sample.
+These results indicate that the $H_{0}$ tension would bias baryon density estimates, although the induced $\Omega_{\rm b}$~inconsistency” remains within $1\sigma$ uncertainty in our sample.
 (ii) Simulation-based prior cases using FRB + GW, where fixing empirical DM model parameters from the IllustrisTNG simulation and further fixing $f_{\rm d}$ to a redshift-dependent form yield $\Omega_{\rm b}=0.0491\pm 0.0054$ and $0.0507\pm 0.0052$, respectively.
@@
 Therefore, combining FRB and GW data not only avoids the bias induced by the $H_{0}$ tension, but also mitigates model-dependent uncertainties.
+We caveat that our baryon census still depends on the adopted $\Omega_{\rm m}$ prior and the assumed GW population model.
+Nevertheless, the $\Omega_{\rm b}$ values inferred when removing these assumptions show only modest shifts and remain consistent with CMB + BBN values at the $1\sigma$ level.
+Using larger samples of localized FRBs and GW standard sirens, future analyses would jointly constrain $\Omega_{\rm m}$, GW population parameters, and $\Omega_{\rm b}$ more robustly, thereby reducing this dependence.
 Overall, the novel combination of FRBs and GWs enables a direct, unique, and robust measurement of both the baryon content and cosmological distances at low redshifts.
-With an increasing number of GW detections expected in the upcoming LVK run, and potentially thousands of localized FRBs expected from CHIME/Outriggers (Masui, 2023) and DSA-110 (Law et al., 2024), their synergy will soon emerge as a powerful low-redshift cosmological probe.
-5 Acknowledgments
+With an increasing number of GW detections expected in the upcoming LVK run, and potentially thousands of localized FRBs expected from CHIME/Outriggers masui2023pinpointing and DSA-110 Law:2023ibd, their synergy will soon emerge as a powerful low-redshift cosmological probe.
+V Acknowledgments
 We thank Tian-Nuo Li, Meng-Lin Zhao, Yang Liu, and Qin Wu for fruitful discussions.
-This work was supported by the National SKA Program of China (grant Nos.
-2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (grant Nos.
-12473001, 11975072, 11875102, and 11835009), the China Manned Space Program (grant No.
-CMS-CSST-2025-A02), and the National 111 Project (grant No.
-B16009).
+This work was supported by the National Natural Science Foundation of China (Grants Nos.
+12533001, 12575049, 12473001, and 12473091), the National SKA Program of China (Grants Nos.
+2022SKA0110200 and 2022SKA0110203), the China Manned Space Program (Grant No.
+CMS-CSST-2025-A02), the National 111 Project (Grant No.
+B16009), and the Fundamental Research Funds for the Central Universities (Grant No.
+N2405008).
 Appendix A Constraints on $H_{0}$ and DM parameters
-Figure 3: Same as Figure 1 but for posterior distribution of $H_{0}$.
+Figure 4: Same as Fig.
+1 but for posterior distribution of $H_{0}$.
 The median and $1\sigma$ credible interval from FRB + GW is $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1.
-The orange and black errorbars correspond to constraints from FRB + $\Omega_{\rm b}h^{2}$ (BBN prior (Cooke et al., 2018)) and GW data alone, respectively.
-The green, yellow, red, and gray shaded bands denote the results from Planck18 (Aghanim et al., 2020), SH0ES (Riess et al., 2022)), Wu22 (Wu et al., 2022), and James22 (James et al., 2022), respectively.
-Note that Wu22 adopts the TNG prior, and both Wu22 and James22 adopt $\Omega_{\rm b}h^{2}$ priors from BBN (Cooke et al., 2018) and CMB (Aghanim et al., 2020), respectively.
+The orange and black errorbars correspond to constraints from FRB + $\Omega_{\rm b}h^{2}$ (BBN prior Cooke:2017cwo) and GW data alone, respectively.
+The green, yellow, red, and gray shaded bands denote the results from Planck18 2020Planck, SH0ES Riess:2021jrx), Wu22 Wu:2021jyk, and James22 James:2022dcx, respectively.
+Note that Wu22 adopts the TNG prior, and both Wu22 and James22 adopt $\Omega_{\rm b}h^{2}$ priors from BBN Cooke:2017cwo and CMB 2020Planck, respectively.
 All results are at $1\sigma$ uncertainty.
 In this Appendix, we report and discuss the constraints on $H_{0}$ and DM parameters ($f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$).
-The constraints on $H_{0}$ are listed in Table 1, and posteriors and corresponding credible intervals for $H_{0}$ are shown in Figure 3.
+The constraints on $H_{0}$ are listed in Table 1, and posteriors and corresponding credible intervals for $H_{0}$ are shown in Fig.
+4.
 A direct measurement of $H_{0}$ using FRBs alone is difficult due to strong parameter degeneracies.
@@
 Previous studies predominantly inferred $H_{0}$ by fixing the baryon density parameter at the early-Universe value.
-For example, Wu et al.
-(2022) (hereafter Wu22) and James et al.
-(2022) (hereafter James22) used different FRB samples to derive $H_{0}=69.31^{+6.21}_{-6.63}~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$ and $H_{0}=73^{+12}_{-8}~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$, by assuming external priors on $\Omega_{\rm b}h^{2}\sim\mathcal{N}(0.02235,0.00049)$ from BBN and $\Omega_{\rm b}h^{2}=0.02242$ from Planck18, respectively, which are shown as red and gray bands in Figure 3.
-Recently, Wang et al.
-(2025) and Gao et al.
-(2025a) also used $92$ and $108$ localized FRBs to obtain $H_{0}=69.04_{-2.07}^{+2.30}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ and $H_{0}=69.40_{-1.97}^{+2.14}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$, respectively, both assuming BBN and TNG priors.
+For example, Refs.
+Wu:2021jyk (hereafter Wu22) and James:2022dcx (hereafter James22) used different FRB samples to derive $H_{0}=69.31^{+6.21}_{-6.63}~\mathrm{km~s^{-1}~Mpc^{-1}}$ and $H_{0}=73^{+12}_{-8}~\mathrm{km~s^{-1}~Mpc^{-1}}$, by assuming external priors on $\Omega_{\rm b}h^{2}\sim\mathcal{N}(0.02235,0.00049)$ from BBN and $\Omega_{\rm b}h^{2}=0.02242$ from Planck18, respectively, which are shown as red and gray bands in Fig.
+4.
+Recently, Refs.
+Wang:2025ugc and Gao:2025fcr also used $92$ and $108$ localized FRBs to obtain $H_{0}=69.04_{-2.07}^{+2.30}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$ and $H_{0}=69.40_{-1.97}^{+2.14}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$, respectively, both assuming BBN and TNG priors.
 Using the same BBN prior, we obtain $H_{0}=66.2^{+5.3}_{-7.3}$ km s-1  Mpc-1from FRB + $\Omega_{\rm b}h^{2}$, which slightly favors the Planck18 result with a precision comparable to that of FRB + GW.
@@
 More data is required to offer further insight into arbitrating the $H_{0}$ tension.
-Notably, time-delay cosmography using strongly lensed FRBs by massive galaxies can offer an independent measurement of $H_{0}$, since the time delays between lensed FRB images can be determined with exceptional precision (Li et al., 2018; Zhang et al., 2025a).
+Notably, time-delay cosmography using strongly lensed FRBs by massive galaxies can offer an independent measurement of $H_{0}$, since the time delays between lensed FRB images can be determined with exceptional precision Li:2017mek; Zhang:2024rra.
 For DM parameters, the constraints are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$.
 The inclusion of GW data has a negligible impact on this constraint.
-For the host contribution, the results of $\mu_{\rm host}$ and $\sigma_{\rm host}$ correspond to a median rest-frame host DM of $108\pm 20~{}\mathrm{pc~{}cm^{-3}}$ (corresponding mean DM is $168^{+57}_{-40}~{}\mathrm{pc~{}cm^{-3}}$), which is consistent with both $\mathrm{DM}_{\rm host}=103_{-48}^{+68}~{}\mathrm{pc~{}cm^{-3}}$ (Yang et al., 2025) and $\mathrm{DM}_{\rm host}=130_{-23}^{+25}~{}\mathrm{pc~{}cm^{-3}}$ (Connor et al., 2025).
+For the host contribution, the results of $\mu_{\rm host}$ and $\sigma_{\rm host}$ correspond to a median rest-frame host DM of $108\pm 20~\mathrm{pc~cm^{-3}}$ (corresponding mean DM is $168^{+57}_{-40}~\mathrm{pc~cm^{-3}}$), which is consistent with both $\mathrm{DM}_{\rm host}=103_{-48}^{+68}~\mathrm{pc~cm^{-3}}$ Yang:2024vqq and $\mathrm{DM}_{\rm host}=130_{-23}^{+25}~\mathrm{pc~cm^{-3}}$ connor2025gas.
 Appendix B FRB data
-In this Appendix, we list the properties of the $104$ localized FRBs used in this work in Table 2.
-[…]
+In this Appendix, we list the properties of the $104$ localized FRBs used in this work in Table LABEL:tab:localized_FRBs.
 Table 2: Properties of $104$ localized FRBs.
-[…]
+The listed parameters include redshift, right ascension (R.A.), declination (Decl.), observed dispersion measure (DM), the Milky Way ISM contribution derived from the NE2001 model, and types of their host galaxies.
+The host types are classified into three categories: type 1 refers to repeating FRBs in dwarf galaxies, type 2 to repeating FRBs in spiral galaxies, and type 3 to one-off FRBs.
+Name Redshift $\mathrm{DM}_{\mathrm{obs}}$ $\mathrm{DM}_{\mathrm{MW,ISM}}$ (NE2001) R.A. (J2000) Decl. (J2000) Host Type Reference
 (pc cm-3) (pc cm-3) () ()
-FRB 20121102A 0.19273 557 188.4 82.9946 $33.1479$ 1 Chatterjee et al. (2017) Tendulkar et al. (2017)
-FRB 20180301A 0.3304 552 151.7 93.2268 $4.6711$ 1 Bhandari et al. (2022)
-FRB 20180924B 0.3214 361.42 40.5 326.1053 $-40.9$ 3 Bannister et al. (2019)
-FRB 20181112A 0.4755 589.27 41.7 327.3485 $-52.9709$ 3 Prochaska et al. (2019)
-FRB 20181223C 0.03024 111.61 19.9 180.9207 $27.5476$ 3 Bhardwaj et al. (2024)
-FRB 20190102C 0.2912 364.5 57.4 322.4157 $-79.4757$ 3 Bhandari et al. (2020)
-FRB 20190110C 0.12244 221.6 37.1 249.3185 $41.4434$ 2 Ibik et al. (2024)
-FRB 20190303A 0.064 223.2 29.8 207.9958 $48.1211$ 2 Michilli et al. (2023)
-FRB 20190418A 0.07132 182.78 70.2 65.8123 $16.0738$ 3 Bhardwaj et al. (2024)
-FRB 20190523A 0.66 760.8 37.2 207.065 $72.4697$ 3 Ravi et al. (2019)
-FRB 20190608B 0.11778 338.7 37.3 334.0199 $-7.8983$ 3 Chittidi et al. (2021)
-FRB 20190611B 0.3778 321.4 57.8 320.7456 $-79.3976$ 3 Heintz et al. (2020)
-FRB 20190614D 0.6 959.2 87.8 65.0755 $73.7067$ 3 Law et al. (2020)
-FRB 20190711A 0.522 593.1 56.5 329.4193 $-80.358$ 1 Heintz et al. (2020)
-FRB 20190714A 0.2365 504.13 38.5 183.9797 $-13.021$ 3 Heintz et al. (2020)
-FRB 20191001A 0.234 506.92 44.2 323.3513 $-54.7478$ 3 Heintz et al. (2020)
-FRB 20191106C 0.10775 332.2 25 199.5801 $42.9997$ 2 Ibik et al. (2024)
-FRB 20191228A 0.2432 297.5 32.9 344.4304 $-28.5941$ 3 Bhandari et al. (2022)
-FRB 20200223B 0.06024 201.8 45.6 8.2695 $8.83132$ Ibik et al. (2024)
-FRB 20200430A 0.1608 380.1 27.2 229.7064 $12.3763$ 3 Heintz et al. (2020)
-FRB 20200723B 0.0085 244 33 190.1583 $-5.135$ 3 Shin et al. (2024)
-FRB 20200906A 0.3688 577.8 35.8 53.4962 $-14.0832$ 3 Bhandari et al. (2022)
-FRB 20201124A 0.098 413.52 139.9 77.0146 $6.06072$ Ravi et al. (2022)
-FRB 20210117A 0.2145 729.1 34.4 339.9792 $-16.1515$ 3 Bhandari et al. (2023)
-FRB 20210320C 0.2797 384.8 39.3 204.4608 $-16.1227$ 3 Shannon et al. (2025)
-FRB 20210410D 0.1415 571.2 56.2 326.0863 $-79.3182$ 3 Caleb et al. (2023)
-FRB 20210603A 0.1772 500.147 39.5 10.2741 $21.2263$ 3 Cassanelli et al. (2024)
-FRB 20210807D 0.1293 251.3 121.2 299.2214 $-0.7624$ 3 Gordon et al. (2023)
-FRB 20211127I 0.0469 234.83 42.5 199.8082 $-18.8378$ 3 Gordon et al. (2023)
-FRB 20211203C 0.3439 636.2 63.7 204.5625 $-31.3801$ 3 Gordon et al. (2023)
-FRB 20211212A 0.0715 206 27.1 157.3509 $1.676889$ 3 Gordon et al. (2023)
-FRB 20220105A 0.2785 583 22 208.8039 $22.4665$ 3 Gordon et al. (2023)
-FRB 20220204A 0.4012 612.584 50.7 274.2263 $69.7225$ 3 Sharma et al. (2024)
-FRB 20220207C 0.04304 262.38 76.1 310.1995 $72.8823$ 3 Law et al. (2024)
-FRB 20220208A 0.351 437 101.6 319.3483 $71.54$ 3 Sharma et al. (2024)
-FRB 20220307B 0.2507 499.15 128.2 350.8745 $72.1924$ 3 Law et al. (2024)
-FRB 20220310F 0.477958 462.24 46.3 134.7204 $73.4908$ 3 Law et al. (2024)
-FRB 20220330D 0.3714 468.1 38.6 165.7256 $71.7535$ 3 Sharma et al. (2024)
-FRB 20220418A 0.622 623.25 36.7 219.1056 $70.0959$ 3 Law et al. (2024)
-FRB 20220501C 0.381 449.5 30.6 352.3792 $-32.4907$ 3 Shannon et al. (2025)
-FRB 20220506D 0.30039 396.97 84.6 318.0448 $72.8273$ 3 Law et al. (2024)
-FRB 20220509G 0.0894 269.53 55.6 282.67 $70.2438$ 3 Law et al. (2024)
-FRB 20220521B 1.354 1342.9 138.8 351.036 $71.138$ 3 Shannon et al. (2025)
-FRB 20220529A 0.1839 246 40 19.1042 $20.6325$ 1 Li et al. (2025)
-FRB 20220610A 1.016 1458.15 31 351.0732 $-33.5137$ 3 Ryder et al. (2023)
-FRB 20220717A 0.36295 637.34 118.3 293.3042 $-19.2877$ 3 Rajwade et al. (2024)
-FRB 20220725A 0.1926 290.4 30.7 353.3152 $-35.9902$ 3 Shannon et al. (2025)
-FRB 20220726A 0.361 686.55 89.5 73.94567 $69.9291$ 3 Sharma et al. (2024)
-FRB 20220825A 0.241397 651.24 78.5 311.9815 $72.585$ 3 Law et al. (2024)
-FRB 20220914A 0.1139 631.28 54.7 282.0568 $73.3369$ 3 Law et al. (2024)
-FRB 20220918A 0.491 656.8 153.1 17.5921 $70.8113$ 3 Shannon et al. (2025)
-FRB 20220920A 0.158239 314.99 39.9 240.2571 $70.9188$ 3 Law et al. (2024)
-FRB 20221012A 0.284669 441.08 54.3 280.7987 $70.5242$ 3 Law et al. (2024)
-FRB 20221027A 0.229 452.5 47.2 129.6104 $71.7315$ 3 Sharma et al. (2024)
-FRB 20221029A 0.975 1391.75 43.8 141.9634 $72.4523$ 3 Sharma et al. (2024)
-FRB 20221101B 0.2395 491.554 131.2 342.2162 $70.6812$ 3 Sharma et al. (2024)
-FRB 20221106A 0.2044 343.8 34.8 56.7048 $-25.5698$ 3 Shannon et al. (2025)
-FRB 20221113A 0.2505 411.027 91.7 71.411 $70.3074$ 3 Sharma et al. (2024)
-FRB 20221116A 0.2764 643.448 132.3 21.2102 $72.6539$ 3 Sharma et al. (2024)
-FRB 20221219A 0.553 706.708 44.4 257.6298 $71.6268$ 3 Sharma et al. (2024)
-FRB 20230124A 0.0939 590.574 38.6 231.9163 $70.9681$ 3 Sharma et al. (2024)
-FRB 20230203A 0.1464 420.1 67.299 151.66159 $5.69413$ Amiri et al. (2025)
-FRB 20230216A 0.531 828 38.5 155.9717 $1.4678$ 3 Sharma et al. (2024)
-FRB 20230222A 0.1223 706.1 33.31 106.96036 $11.22452$ 3 Amiri et al. (2025)
-FRB 20230222B 0.11 187.8 56.9977 238.73912 $0.89873$ Amiri et al. (2025)
-FRB 20230307A 0.271 608.9 37.6 177.7813 $71.6956$ 3 Sharma et al. (2024)
-FRB 20230311A 0.1918 364.3 67.2352 91.10966 $55.94595$ 3 Amiri et al. (2025)
-FRB 20230501A 0.301 532.5 125.6 340.0272 $70.9222$ 3 Sharma et al. (2024)
-FRB 20230506C 0.3896 772 68 12.0998 $42.0061$ 2 Anna-Thomas et al. (2025)
-FRB 20230526A 0.157 361.4 31.9 22.2326 $-52.7173$ 3 Shannon et al. (2025)
-FRB 20230626A 0.327 451.2 39.2 240.7125 $71.7142$ 3 Sharma et al. (2024)
-FRB 20230628A 0.1265 344.952 39 166.7867 $72.2818$ 3 Sharma et al. (2024)
-FRB 20230703A 0.1184 291.3 38.1566 184.62445 $48.72993$ 3 Amiri et al. (2025)
-FRB 20230708A 0.105 411.51 60.3 303.1155 $-55.3563$ 3 Shannon et al. (2025)
-FRB 20230712A 0.4525 587.567 39.2 167.3585 $72.5578$ 3 Sharma et al. (2024)
-FRB 20230730A 0.2115 312.5 114.6465 54.66456 $3.15933$ Amiri et al. (2025)
-FRB 20230808F 0.3472 653.2 31 53.3041 $-51.9353$ 3 Hanmer et al. (2025)
-FRB 20230814A 0.553 696.4 104.8 335.9748 $73.0259$ 3 Connor et al. (2025)
-FRB 20230902A 0.3619 440.1 34.1 52.1398 $-47.3335$ 3 Shannon et al. (2025)
-FRB 20230930A 0.0925 456 70 10.5 $41.4$ 3 Anna-Thomas et al. (2025)
-FRB 20231011A 0.0783 186.3 58.5524 18.2411 $41.7491$ 3 Amiri et al. (2025)
-FRB 20231017A 0.245 344.2 31.1515 346.75429 $6.652683$ Amiri et al. (2025)
-FRB 20231025B 0.3238 368.7 65.5787 270.78807 $63.98908$ 3 Amiri et al. (2025)
-FRB 20231120A 0.07 438.9 43.8 143.6169 $71.7574$ 3 Sharma et al. (2024)
-FRB 20231123A 0.0729 302.1 37.1085 82.62325 $4.47554$ 3 Amiri et al. (2025)
-FRB 20231123B 0.2625 396.7 40.2 242.5382 $70.7851$ 3 Sharma et al. (2024)
-FRB 20231128A 0.1079 331.6 64.7368 199.5782 $42.99271$ 1 Amiri et al. (2025)
-FRB 20231204A 0.0644 221 34.9372 207.99903 $48.116$ 2 Amiri et al. (2025)
-FRB 20231206A 0.0659 457.7 37.5114 112.44284 $56.25627$ 3 Amiri et al. (2025)
-FRB 20231220A 0.3355 491.2 49.9 123.9087 $73.6599$ 3 Connor et al. (2025)
-FRB 20231226A 0.1569 329.9 38.1 155.3638 $6.1103$ 3 Shannon et al. (2025)
-FRB 20231229A 0.019 198.5 65.1335 26.46783 $5.112923$ Amiri et al. (2025)
-FRB 20231230A 0.0298 131.4 32.2305 72.79761 $2.39398$ 3 Amiri et al. (2025)
-FRB 20240114A 0.13 527.65 49.7 321.9161 $4.3292$ 1 Tian et al. (2024)
-FRB 20240119A 0.376 483.1 38 224.4672 $71.6118$ 3 Connor et al. (2025)
-FRB 20240123A 0.968 1462 90.2 68.2625 $71.9453$ 3 Connor et al. (2025)
-FRB 20240201A 0.042729 374.5 38.6 149.9056 $14.088$ 3 Shannon et al. (2025)
-FRB 20240209A 0.1384 176.518 63.1 289.8504 $86.0609$ 2 Eftekhari et al. (2025)
-FRB 20240210A 0.023686 283.73 28.7 8.7796 $-28.2708$ 3 Shannon et al. (2025)
-FRB 20240213A 0.1185 357.4 40 166.1683 $74.0754$ 3 Connor et al. (2025)
-FRB 20240215A 0.21 549.5 47.9 268.4413 $70.2324$ 3 Connor et al. (2025)
-FRB 20240229A 0.287 491.15 38 169.9835 $70.6762$ 3 Connor et al. (2025)
-FRB 20240304A 0.2423 652.6 73.9 136.3304 $-16.2282$ 3 Shannon et al. (2025)
-FRB 20240310A 0.127 601.8 30.1 17.6219 $-44.4394$ 3 Shannon et al. (2025)
-Note.
-— The results are derived from $104$ localized FRBs alone or in combination with $47$ GW events, under different assumptions.
-For all cases, the $1\sigma$ constraints on DM parameters are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$.
-$H_{0}$ is in units of km s-1  Mpc-1.
+FRB 20121102A 0.19273 557 188.4 82.9946 $33.1479$ 1 Chatterjee:2017dqg; Tendulkar2016
+FRB 20180301A 0.3304 552 151.7 93.2268 $4.6711$ 1 Bhandari2022
+FRB 20180924B 0.3214 361.42 40.5 326.1053 40.9-40.9 3 Bannister2019
+FRB 20181112A 0.4755 589.27 41.7 327.3485 52.9709-52.9709 3 Prochaska2019
+FRB 20181223C 0.03024 111.61 19.9 180.9207 $27.5476$ 3 Bhardwaj2024
+FRB 20190102C 0.2912 364.5 57.4 322.4157 79.4757-79.4757 3 Bhandari2020
+FRB 20190110C 0.12244 221.6 37.1 249.3185 $41.4434$ 2 Ibik2024
+FRB 20190303A 0.064 223.2 29.8 207.9958 $48.1211$ 2 Michilli2023
+FRB 20190418A 0.07132 182.78 70.2 65.8123 $16.0738$ 3 Bhardwaj2024
+FRB 20190523A 0.66 760.8 37.2 207.065 $72.4697$ 3 Ravi2019
+FRB 20190608B 0.11778 338.7 37.3 334.0199 7.8983-7.8983 3 Chittidi2021
+FRB 20190611B 0.3778 321.4 57.8 320.7456 79.3976-79.3976 3 Heintz2020
+FRB 20190614D 0.6 959.2 87.8 65.0755 $73.7067$ 3 Law2020
+FRB 20190711A 0.522 593.1 56.5 329.4193 80.358-80.358 1 Heintz2020
+FRB 20190714A 0.2365 504.13 38.5 183.9797 13.021-13.021 3 Heintz2020
+FRB 20191001A 0.234 506.92 44.2 323.3513 54.7478-54.7478 3 Heintz2020
+FRB 20191106C 0.10775 332.2 25 199.5801 $42.9997$ 2 Ibik2024
+FRB 20191228A 0.2432 297.5 32.9 344.4304 28.5941-28.5941 3 Bhandari2022
+FRB 20200223B 0.06024 201.8 45.6 8.2695 $28.8313$ 2 Ibik2024
+FRB 20200430A 0.1608 380.1 27.2 229.7064 $12.3763$ 3 Heintz2020
+FRB 20200723B 0.0085 244 33 190.1583 5.135-5.135 3 2024arXiv241007307S
+FRB 20200906A 0.3688 577.8 35.8 53.4962 14.0832-14.0832 3 Bhandari2022
+FRB 20201124A 0.098 413.52 139.9 77.0146 $26.0607$ 2 Ravi2022
+FRB 20210117A 0.2145 729.1 34.4 339.9792 16.1515-16.1515 3 Bhandari2023
+FRB 20210320C 0.2797 384.8 39.3 204.4608 16.1227-16.1227 3 Shannon2024
+FRB 20210410D 0.1415 571.2 56.2 326.0863 79.3182-79.3182 3 Caleb2023
+FRB 20210603A 0.1772 500.147 39.5 10.2741 $21.2263$ 3 Cassanelli2023
+FRB 20210807D 0.1293 251.3 121.2 299.2214 0.7624-0.7624 3 Gordon2023
+FRB 20211127I 0.0469 234.83 42.5 199.8082 18.8378-18.8378 3 Gordon2023
+FRB 20211203C 0.3439 636.2 63.7 204.5625 31.3801-31.3801 3 Gordon2023
+FRB 20211212A 0.0715 206 27.1 157.3509 $1.676889$ 3 Gordon2023
+FRB 20220105A 0.2785 583 22 208.8039 $22.4665$ 3 Gordon2023
+FRB 20220204A 0.4012 612.584 50.7 274.2263 $69.7225$ 3 Sharma2024
+FRB 20220207C 0.04304 262.38 76.1 310.1995 $72.8823$ 3 Law2024
+FRB 20220208A 0.351 437 101.6 319.3483 $71.54$ 3 Sharma2024
+FRB 20220307B 0.2507 499.15 128.2 350.8745 $72.1924$ 3 Law2024
+FRB 20220310F 0.477958 462.24 46.3 134.7204 $73.4908$ 3 Law2024
+FRB 20220330D 0.3714 468.1 38.6 165.7256 $71.7535$ 3 Sharma2024
+FRB 20220418A 0.622 623.25 36.7 219.1056 $70.0959$ 3 Law2024
+FRB 20220501C 0.381 449.5 30.6 352.3792 32.4907-32.4907 3 Shannon2024
+FRB 20220506D 0.30039 396.97 84.6 318.0448 $72.8273$ 3 Law2024
+FRB 20220509G 0.0894 269.53 55.6 282.67 $70.2438$ 3 Law2024
+FRB 20220521B 1.354 1342.9 138.8 351.036 $71.138$ 3 Shannon2024
+FRB 20220529A 0.1839 246 40 19.1042 $20.6325$ 1 Li:2025ckl
+FRB 20220610A 1.016 1458.15 31 351.0732 33.5137-33.5137 3 Ryder2023
+FRB 20220717A 0.36295 637.34 118.3 293.3042 19.2877-19.2877 3 Rajwade2024
+FRB 20220725A 0.1926 290.4 30.7 353.3152 35.9902-35.9902 3 Shannon2024
+FRB 20220726A 0.361 686.55 89.5 73.94567 $69.9291$ 3 Sharma2024
+FRB 20220825A 0.241397 651.24 78.5 311.9815 $72.585$ 3 Law2024
+FRB 20220914A 0.1139 631.28 54.7 282.0568 $73.3369$ 3 Law2024
+FRB 20220918A 0.491 656.8 153.1 17.5921 $70.8113$ 3 Shannon2024
+FRB 20220920A 0.158239 314.99 39.9 240.2571 $70.9188$ 3 Law2024
+FRB 20221012A 0.284669 441.08 54.3 280.7987 $70.5242$ 3 Law2024
+FRB 20221027A 0.229 452.5 47.2 129.6104 $71.7315$ 3 Sharma2024
+FRB 20221029A 0.975 1391.75 43.8 141.9634 $72.4523$ 3 Sharma2024
+FRB 20221101B 0.2395 491.554 131.2 342.2162 $70.6812$ 3 Sharma2024
+FRB 20221106A 0.2044 343.8 34.8 56.7048 25.5698-25.5698 3 Shannon2024
+FRB 20221113A 0.2505 411.027 91.7 71.411 $70.3074$ 3 Sharma2024
+FRB 20221116A 0.2764 643.448 132.3 21.2102 $72.6539$ 3 Sharma2024
+FRB 20221219A 0.553 706.708 44.4 257.6298 $71.6268$ 3 Sharma2024
+FRB 20230124A 0.0939 590.574 38.6 231.9163 $70.9681$ 3 Sharma2024
+FRB 20230203A 0.1464 420.1 67.299 151.66159 $35.6941$ 3 Amiri:2025sbi
+FRB 20230216A 0.531 828 38.5 155.9717 $1.4678$ 3 Sharma2024
+FRB 20230222A 0.1223 706.1 33.31 106.96036 $11.22452$ 3 Amiri:2025sbi
+FRB 20230222B 0.11 187.8 56.9977 238.73912 $30.8987$ 3 Amiri:2025sbi
+FRB 20230307A 0.271 608.9 37.6 177.7813 $71.6956$ 3 Sharma2024
+FRB 20230311A 0.1918 364.3 67.2352 91.10966 $55.94595$ 3 Amiri:2025sbi
+FRB 20230501A 0.301 532.5 125.6 340.0272 $70.9222$ 3 Sharma2024
+FRB 20230506C 0.3896 772 68 12.0998 $42.0061$ 2 AnnaThomas2025
+FRB 20230526A 0.157 361.4 31.9 22.2326 52.7173-52.7173 3 Shannon2024
+FRB 20230626A 0.327 451.2 39.2 240.7125 $71.7142$ 3 Sharma2024
+FRB 20230628A 0.1265 344.952 39 166.7867 $72.2818$ 3 Sharma2024
+FRB 20230703A 0.1184 291.3 38.1566 184.62445 $48.72993$ 3 Amiri:2025sbi
+FRB 20230708A 0.105 411.51 60.3 303.1155 55.3563-55.3563 3 Shannon2024
+FRB 20230712A 0.4525 587.567 39.2 167.3585 $72.5578$ 3 Sharma2024
+FRB 20230730A 0.2115 312.5 114.6465 54.66456 $33.1593$ 3 Amiri:2025sbi
+FRB 20230808F 0.3472 653.2 31 53.3041 51.9353-51.9353 3 2025MNRAS.538.1800H
+FRB 20230814A 0.553 696.4 104.8 335.9748 $73.0259$ 3 connor2025gas
+FRB 20230902A 0.3619 440.1 34.1 52.1398 47.3335-47.3335 3 Shannon2024
+FRB 20230930A 0.0925 456 70 10.5 $41.4$ 3 AnnaThomas2025
+FRB 20231011A 0.0783 186.3 58.5524 18.2411 $41.7491$ 3 Amiri:2025sbi
+FRB 20231017A 0.245 344.2 31.1515 346.75429 $36.65268$ 3 Amiri:2025sbi
+FRB 20231025B 0.3238 368.7 65.5787 270.78807 $63.98908$ 3 Amiri:2025sbi
+FRB 20231120A 0.07 438.9 43.8 143.6169 $71.7574$ 3 Sharma2024
+FRB 20231123A 0.0729 302.1 37.1085 82.62325 $4.47554$ 3 Amiri:2025sbi
+FRB 20231123B 0.2625 396.7 40.2 242.5382 $70.7851$ 3 Sharma2024
+FRB 20231128A 0.1079 331.6 64.7368 199.5782 $42.99271$ 1 Amiri:2025sbi
+FRB 20231204A 0.0644 221 34.9372 207.99903 $48.116$ 2 Amiri:2025sbi
+FRB 20231206A 0.0659 457.7 37.5114 112.44284 $56.25627$ 3 Amiri:2025sbi
+FRB 20231220A 0.3355 491.2 49.9 123.9087 $73.6599$ 3 connor2025gas
+FRB 20231226A 0.1569 329.9 38.1 155.3638 $6.1103$ 3 Shannon2024
+FRB 20231229A 0.019 198.5 65.1335 26.46783 $35.11292$ 3 Amiri:2025sbi
+FRB 20231230A 0.0298 131.4 32.2305 72.79761 $2.39398$ 3 Amiri:2025sbi
+FRB 20240114A 0.13 527.65 49.7 321.9161 $4.3292$ 1 Tian:2024ygd
+FRB 20240119A 0.376 483.1 38 224.4672 $71.6118$ 3 connor2025gas
+FRB 20240123A 0.968 1462 90.2 68.2625 $71.9453$ 3 connor2025gas
+FRB 20240201A 0.042729 374.5 38.6 149.9056 $14.088$ 3 Shannon2024
+FRB 20240209A 0.1384 176.518 63.1 289.8504 $86.0609$ 2 2025ApJ...979L..22E
+FRB 20240210A 0.023686 283.73 28.7 8.7796 28.2708-28.2708 3 Shannon2024
+FRB 20240213A 0.1185 357.4 40 166.1683 $74.0754$ 3 connor2025gas
+FRB 20240215A 0.21 549.5 47.9 268.4413 $70.2324$ 3 connor2025gas
+FRB 20240229A 0.287 491.15 38 169.9835 $70.6762$ 3 connor2025gas
+FRB 20240304A 0.2423 652.6 73.9 136.3304 16.2282-16.2282 3 Shannon2024
+FRB 20240310A 0.127 601.8 30.1 17.6219 44.4394-44.4394 3 Shannon2024
+Table 2: Continued

### v3

@@
 In the context of the decade-long $H_{0}$ tension, the significant discrepancy between early- and late-Universe measurements of $H_{0}$ implies that fixing its value or imposing an external prior could bias the baryon census.
-To address this concern, we construct a late-Universe probe framework that unifies fast radio bursts (FRBs) and gravitational-wave (GW) standard sirens, which can respectively resolve the “missing baryon” problem and the $H_{0}$ tension through their dispersion measures and absolute luminosity distances.
+To address this concern, we construct a late-Universe probe framework that unifies fast radio bursts (FRBs) and gravitational-wave (GW) standard sirens, which can respectively resolve the “missing baryon” problem and the $H_{0}$ tension through their dispersion measures (DMs) and absolute luminosity distances.
 By combining $104$ localized FRBs with $47$ GW events, we obtain an $H_{0}$-free measurement of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), in concordance with early-Universe observations of CMB + BBN.
+The result is tightly anchored by GW-inferred $H_{0}$ through the strong $\Omega_{\rm b}$-$H_{0}$ degeneracy.
 Although the current precision ($\sim 13\%$) is limited by sample size, the growing detections of both FRBs and GWs will make their synergy a powerful probe of low-redshift cosmology.
-I Introduction
-The concordance cosmology predicts consistency among independent probes across all epochs Ostriker:1995su.
-However, while the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) provide precise estimates of the baryon content in the early Universe Cooke:2017cwo, about $30\%$ baryons remain undetected in the local Universe Shull:2011aa.
-These missing baryons are thought to reside in diffuse ionized gas that is too faint to be directly observed, known as the “missing baryon problem” Fukugita:1997bi; cen1999baryons.
-Fast radio bursts (FRBs) are bright, millisecond-duration radio transients of extragalactic origin lorimer2007; thornton2013pop, which can be broadly classified into repeaters and non-repeaters 2016Natur.531..202S; Sun2025a; Sun2025b.
-Their observed dispersion measures (DMs) trace the integrated column density of free electrons along the line of sight (LoS), making them sensitive to baryons in the intergalactic medium (IGM) mcquinn2013locating.
-Thus, FRBs with known redshifts derived from precise host galaxy localization can be used to trace the missing baryons Deng:2013aga; connor2025gas.
-Ref.
-Macquart:2020lln (hereafter Macquart20) first employed $5$ localized FRBs to estimate the present-day cosmic baryon density of $\Omega_{\rm b}=0.051^{+0.021}_{-0.020}$ at the $95\%$ confidence level, which agrees with the joint CMB and BBN result and likely resolves the local baryon problem.
-Ref.
-Yang:2022ftm (hereafter Yang22) used an updated sample of $22$ localized FRBs and obtained $\Omega_{\rm b}=0.0490_{-0.0033}^{+0.0036}$ at the $1\sigma$ level, reinforcing the evidence for a complete cosmic baryon census.
-With the high event rate and increasing number of localized events, FRBs have emerged as a promising cosmological probe Zhou:2014yta; Gao:2014iva; dai2023ska; Zhang:2023gye; see Refs.
-Bhandari:2021thi, Wu:2024iyu, and Glowacki:2024cgu for recent reviews.
-However, FRB cosmology based on DM is hindered by parameter degeneracies, which result in two main sources of systematic uncertainties Kumar:2019qhc.
+Introduction.
+The concordance cosmology predicts consistency among independent probes across all epochs Ostriker and Steinhardt (1995).
+Yet, while the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) provide precise estimates of the baryon content in the early Universe Cooke et al.
+(2018), about $30\%$ baryons remain undetected in the local Universe Shull et al.
+(2012).
+These missing baryons are thought to reside in diffuse ionized gas that is too faint to be directly observed, known as the “missing baryon problem” Fukugita et al.
+(1998); Cen and Ostriker (1999).
+Fast radio bursts (FRBs) are bright, millisecond-duration radio transients of extragalactic origin Lorimer et al.
+(2007); Thornton et al.
+(2013), which can be broadly classified into repeaters and non-repeaters Spitler et al.
+(2016); Sun et al.
+(2025a, b).
+Their observed dispersion measures (DMs) trace the integrated column density of free electrons along the line of sight (LoS), making them sensitive to baryons in the intergalactic medium (IGM) McQuinn (2013).
+Localized FRBs with measured redshifts thus provide a direct probe of the missing baryons Deng and Zhang (2014); Connor et al.
+(2025).
+Using 5 and 22 localized FRBs, Refs.
+Macquart et al.
+(2020); Yang et al.
+(2022) (hereafter Macquart20 and Yang22) inferred $\Omega_{\rm b}$ values consistent with CMB + BBN, suggesting a nearly complete baryon census.
+With the high event rate and increasing number of localized events, FRBs have emerged as a promising cosmological probe to, e.g., probe dark energy Zhou et al.
+(2014); Gao et al.
+(2014); Walters et al.
+(2018); Dai (2023); Zhang et al.
+(2023), the Hubble constant ($H_{0}$) Li et al.
+(2018); Hagstotz et al.
+(2022); Wu et al.
+(2022); James et al.
+(2022); Liu et al.
+(2023); Zhao et al.
+(2022); Kalita et al.
+(2025); Fortunato et al.
+(2025); Zhang et al.
+(2025a); Wang et al.
+(2025a); Gao et al.
+(2025a); Zhang and Zhang (2025); Gao et al.
+(2025b); Zhuge et al.
+(2026); Xu et al.
+(2025); Jia et al.
+(2025); Wu et al.
+(2026), as well as baryonic feedback Sharma et al.
+(2026); Reischke and Hagstotz (2025), and the large-scale structure Hussaini et al.
+(2025); Wang et al.
+(2025b); see Refs.
+Bhandari and Flynn (2021); Petroff et al.
+(2022); Xiao et al.
+(2021); Zhang (2023); Wu and Wang (2024); Glowacki and Lee (2025) for recent reviews.
+However, FRB cosmology based on DM is seriously hindered by parameter degeneracies, which result in two main sources of systematic uncertainties Kumar and Linder (2019).
 The first is from extracting the IGM contribution ($\mathrm{DM}_{\rm IGM}$) from the total DM budget (see Eq.
@@
 Macquart20 modeled their probability distributions and Refs.
-Zhang:2020mgq; Zhang:2020xoc further employed the IllustrisTNG simulations to derive priors on these empirical DM model parameters for different FRB populations, which effectively reduce biases induced by host galaxy diversity and IGM fluctuations.
-However, the second uncertainty from the analysis of the extracted $\mathrm{DM}_{\rm IGM}$ remains challenging to control.
-This is because its redshift relation (known as the Macquart relation) exhibits intrinsic degeneracies among $\Omega_{\rm b}$, the Hubble constant ($H_{0}$), and the baryon fraction in diffuse ionized gas ($f_{\rm d}$) (see Eq.
+Zhang et al.
+(2020, 2021) further employed the IllustrisTNG (TNG) simulations to derive priors on these empirical model parameters for different FRB populations, which could help mitigate biases from host diversity and IGM fluctuations.
+However, the second, and more fundamental, uncertainty from the analysis of the extracted $\mathrm{DM}_{\rm IGM}$ remains challenging to control.
+This is because its redshift relation (known as the Macquart relation) exhibits intrinsic degeneracies among $\Omega_{\rm b}$, $H_{0}$, and the baryon fraction in diffuse ionized gas ($f_{\rm d}$) (see Eq.
 (3)).
-Given the limited constraints on $f_{\rm d}$, it is commonly modeled as either a constant or a redshift-dependent function in the literature Li:2020qei; Qiang:2020vta; Wang:2022ami; Liu:2025fdf.
-$\Omega_{\rm b}$ is intrinsically linked to $H_{0}$ through the critical density of the Universe.
-Although $H_{0}$ has been precisely measured, estimates from early- (e.g., Planck 2018 CMB 2020Planck (hereafter Planck18)) and late-Universe (e.g., SH0ES Cepheid-calibrated Type Ia supernovae (SNe) Riess:2021jrx) observations show a significant discrepancy now exceeding $5\sigma$, which has widely been known as the $H_{0}$ tension” Guo:2018ans; DiValentino:2021izs; Abdalla:2022yfr; Hu:2023jqc; CosmoVerseNetwork:2025alb.
+Given the limited constraints on $f_{\rm d}$, it is commonly modeled as either a constant or a redshift-dependent function in the literature Jaroszynski (2019); Li et al.
+(2020); Qiang and Wei (2020); Zhao et al.
+(2020a); Qiu et al.
+(2022); Wang and Wei (2023); Gao et al.
+(2023); Liu et al.
+(2025).
+On the other hand, $H_{0}$ is coupled to $\Omega_{\rm b}$ through the critical density of the Universe.
+Although $H_{0}$ has been precisely measured, estimates from early- (e.g., Planck 2018 CMB Aghanim et al.
+(2020) (hereafter Planck18)) and late-Universe (e.g., SH0ES Cepheid-calibrated Type Ia supernovae (SNe) Riess et al.
+(2022)) observations show a significant discrepancy now exceeding $5\sigma$, which has widely been known as the $H_{0}$ tension” Guo et al.
+(2019); Di Valentino et al.
+(2021); Abdalla et al.
+(2022); Hu and Wang (2023); Di Valentino et al.
+(2025).
 This inconsistency would introduce systematic uncertainties into the baryon census using FRB samples.
-It should be noted that Macquart20 and Yang22 derived their respective estimates of $\Omega_{\rm b}$ by assuming $H_{0}=70$ km s-1  Mpc-1and a flat prior over 67677474 km s-1  Mpc-1, which provides a reasonable strategy for handling the second uncertainty given the $H_{0}$ tension (note that many studies also constrain $H_{0}$ by adopting early-Universe priors on $\Omega_{\rm b}h^{2}$; see, e.g., Ref.
-Hagstotz:2021jzu).
-Nevertheless, any external assumption on $H_{0}$ could bias the baryon estimate, motivating the need for a direct inference of $\Omega_{\rm b}$ that is entirely untethered to the Hubble tension, e.g., jointly analyzing FRBs and datasets that independently measure $H_{0}$ in the late Universe.
-As an emerging cosmological probe, gravitational-wave (GW) standard sirens offer unique advantages for measuring $H_{0}$ Schutz:1986gp.
+It should be noted that Macquart20 and Yang22 derived their respective estimates of $\Omega_{\rm b}$ by assuming $H_{0}=70$ km s-1  Mpc-1and a flat prior over 67677474 km s-1  Mpc-1, which could unavoidably bias the baryon estimate in the context of the Hubble tension and thereby motivate a direct inference of $\Omega_{\rm b}$, e.g., jointly analyzing FRBs and datasets that independently measure $H_{0}$ in the late Universe.
+As an emerging cosmological probe, gravitational-wave (GW) standard sirens offer unique advantages for measuring $H_{0}$ Schutz (1986).
 This primarily arises from the fact that the amplitudes of GW waveforms are inversely proportional to the luminosity distances ($D_{\rm L}$) to the sources, allowing for a direct measurement of $D_{\rm L}$ through waveform analysis.
-Since $D_{\rm L}\propto 1/H_{0}$, GW standard sirens serve as a powerful and self-calibrating probe for determining $H_{0}$, provided that the redshifts of the GW sources are also known Dalal:2006qt; Cutler:2009qv; Zhao:2010sz; Cai:2016sby; LIGOScientific:2017adf; LIGOScientific:2017vwq; LIGOScientific:2017zic; Chen:2017rfc; Cai:2017aea; Wang:2018lun; Zhang:2019loq; Zhang:2018byx; Zhang:2019ylr; Wang:2019tto; Zhao:2019gyk; Gray:2019ksv; Jin:2020hmc; Yu:2020vyy; Mastrogiovanni:2021wsd; KAGRA:2021duu; Wang:2021srv; Zhu:2021aat; Zhu:2021bpp; Wu:2022dgy; Muttoni:2023prw; Yun:2023ygz; Yu:2023ico; Han:2023exn; Feng:2024lzh; Song:2022siz; Jin:2023sfc; Jin:2023tou; Li:2023gtu; Dong:2024bvw; Xiao:2024nmi; Zhu:2024qpp; Zheng:2024mbo; Han:2025fii; Du:2025odq; Song:2025bio; Jin:2025dvf; Dong:2025ikq.
+Since $D_{\rm L}\propto 1/H_{0}$, GW standard sirens serve as a powerful and self-calibrating probe for determining $H_{0}$, provided that the redshifts of the GW sources are also known Dalal et al.
+(2006); Cutler and Holz (2009); Zhao et al.
+(2011); Cai and Yang (2017); Abbott et al.
+(2017a, b, c); Chen et al.
+(2018); Cai et al.
+(2018); Wang et al.
+(2018); Zhang et al.
+(2019a, b); Zhang (2019); Wang et al.
+(2020); Zhao et al.
+(2020b); Gray et al.
+(2020); Jin et al.
+(2020); Yu et al.
+(2020); Mastrogiovanni et al.
+(2021); Abbott et al.
+(2023a); Wang et al.
+(2022); Zhu et al.
+(2022a, b); Wu et al.
+(2023); Muttoni et al.
+(2023); Yun et al.
+(2023); Yu et al.
+(2024); Han et al.
+(2024); Feng et al.
+(2024); Song et al.
+(2024); Jin et al.
+(2024a, b); Li et al.
+(2024); Dong et al.
+(2025a); Xiao et al.
+(2025); Zhu et al.
+(2024); Zheng et al.
+(2024); Han et al.
+(2026); Du et al.
+(2025); Song et al.
+(2026); Jin et al.
+(2026); Dong et al.
+(2025b).
 During the first three observing runs of the LIGO, Virgo, and KAGRA (LVK) collaboration, approximately 100 GW events were detected and compiled into the GWTC-3 catalog.
 Using 47 GW standard sirens from GWTC-3, Ref.
-LIGOScientific:2021aug achieved a $\sim 10\%$ precision on $H_{0}$, and Ref.
-Song:2025ddm further combined these sirens with an additional strong-lensing system to obtain a model-independent $\sim 8\%$ constraint.
-More recently, with the release of the LVK O4a data set, over 300 GW events have been reported, and by analyzing 142 of them while allowing the population-model parameters of both binary black holes (BBHs) and binary neutron stars (BNSs) to vary freely, the LVK collaboration has obtained a $\sim 14\%$ constraint on $H_{0}$ LIGOScientific:2025jau.
+Abbott et al.
+(2023b) achieved a $\sim 10\%$ precision on $H_{0}$, and Ref.
+Song et al.
+(2025) further combined these sirens with an additional strong-lensing system to obtain a model-independent $\sim 8\%$ constraint.
+More recently, with the release of the LVK O4a data set, over 300 GW events have been reported, and by analyzing 142 of them while varying the population-model parameters of both binary black holes (BBHs) and binary neutron stars (BNSs), the LVK collaboration has obtained a $\sim 14\%$ constraint on $H_{0}$ Abac et al.
+(2025).
 Therefore, the $H_{0}$ information uniquely provided by GW data enables a direct estimate of $\Omega_{\rm b}$ disentangled from $H_{0}$ in the Macquart relation, which offers new insight into resolving systematic effects in FRB cosmology.
-Unlike previous studies that combined FRBs with traditional observations such as SNe Jaroszynski:2018vgh; gjz2023 or CMB Zhao:2020ole; Qiu:2021cww, we highlight that combining with GWs inherently involves calibration-free distances and avoids reliance on early-time physics, thereby eliminating the bias induced by the $H_{0}$ tension.
-In this work, we present the first $H_{0}$-free baryon census at low redshift, within a late-Universe probe framework that unifies $104$ localized FRBs and $47$ GW events from the GWTC-3 catalog to jointly constrain $\Omega_{\rm b}$ and $H_{0}$.
-To control model-dependent systematics, we impose no FRB population model and fully marginalize over all DM parameters [^1] In this context, we refer to the DM parameters as ($f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$), while the empirical DM model parameters refer specifically to ($F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$).
-The parameters are defined in Sec.
-II.1., while on the GW side we adopt a population model for the main analysis and additionally release its parameters to assess robustness.
+In this Letter, we present the first $H_{0}$-free baryon census at low redshift, within a late-Universe probe framework that unifies $104$ localized FRBs and $47$ GW events from the GWTC-3 catalog to jointly constrain $\Omega_{\rm b}$ and $H_{0}$.
 Note that we do not assume any FRB/GW associations (see Refs.
-Wei:2018cgd and Qiang:2024erm for studies based on such associations).
-This work is organized as follows.
-In Sec.
-II, we describe the datasets and likelihood construction for both FRBs and GWs.
-Results and related discussions are presented in Sec.
-III, and we conclude in Sec.
-IV.
-II Data and methodology
-II.1 Localized FRB data
-In this section, we first present the likelihood function for localized FRBs based on DM analysis following Macquart20, then the FRB sample used in this work.
+Wei et al.
+(2018); Qiang et al.
+(2025); Jahns-Schindler and Spitler (2025) for studies based on such associations).
+Localized FRBs.
 The observed dispersion measure can be decomposed into several components:
@@
 Note that we include the factor $1/\langle{\rm DM_{\rm IGM}}\rangle$ suggested by Ref.
-Zhang:2025wif when normalizing Eq.
+Zhang et al.
+(2025b) when normalizing Eq.
 (4).
@@
 $\mathcal{L}_{\mathrm{FRB}}(\{x_{\rm FRB}\})=\sum_{i=1}^{N_{\mathrm{FRB}}}P_{i}\left(\mathrm{DM}^{\prime}_{i}\right).$ (6)
-For the $i$-th FRB, the possibility distribution of its extragalactic DM component ($\mathrm{DM}^{\prime}=\rm DM_{\rm IGM}+\rm DM_{\rm host}$) is computed by convolving Eqs. (2), (4), and (II.1):
+For the $i$-th FRB, the possibility distribution of its extragalactic DM component ($\mathrm{DM}^{\prime}=\rm DM_{\rm IGM}+\rm DM_{\rm host}$) is computed by convolving Eqs.
+(2), (4), and (Localized FRBs.):
 $P_{i}\left(\mathrm{DM}^{\prime}_{i}\right)$ $=\int_{30\,\mathrm{pc\,cm}^{-3}}^{100\,\mathrm{pc\,cm}^{-3}}P_{\rm halo}\left(\mathrm{DM}_{\rm halo}\right)\,\mathrm{dDM}_{\rm halo}$
@@
 $\times P_{\rm IGM}\left(\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}-\mathrm{DM}_{\rm host}\right)\,\mathrm{dDM}_{\rm host}.$ (7)
-We also consider a scenario in which the empirical DM model parameters are fixed, though our main results treat them as free.
-Specifically, we adopt values calibrated from cosmological hydrodynamical simulations, notably the IllustrisTNG simulation, which provides redshift-dependent distributions for both $P_{\rm IGM}(\mathrm{DM}_{\rm IGM})$ Zhang:2020xoc and $P_{\rm host}(\mathrm{DM}_{\rm host})$ Zhang:2020mgq.
-The parameters $A(z)$, $C_{0}(z)$, and $\sigma_{\rm IGM}(z)$ in Eq.
-(4), as well as $\mu_{\rm host}(z)$ and $\sigma_{\rm host}(z)$ in Eq.
-(II.1), are interpolated from best-fit simulation results across redshift.
-Note that $\mu_{\rm host}$ and $\sigma_{\rm host}$ are further fixed based on the host galaxy types (see the seventh column of Table LABEL:tab:localized_FRBs).
-Finally, we select a sample of $104$ localized FRBs from an initial set of $119$, based on the following criteria: (1) $\mathrm{DM}_{\mathrm{MW,ISM}}$ is required to be less than $40\%$ of $\mathrm{DM}_{\rm obs}$ connor2025gas; (2) $\mathrm{DM}_{\rm obs}-\rm DM_{MW,ISM}$ is taken to exceed 80 $\rm{pc}\,\rm{cm}^{-3}\,$, ensuring a positive $\mathrm{DM}_{\rm host}$; (3) FRB 20190520B and FRB 20220831A are excluded due to excess local DM contributions from their host environments.
-The final sample is listed in Appendix B (see Tables LABEL:tab:localized_FRBs).
-II.2 GW standard siren data
-We use 47 GW standard sirens with SNRs greater than 11 and an inverse false alarm rate exceeding 4 yr from GWTC-3222https://gwosc.org/GWTC-3, covering a redshift range of approximately $z<0.8$.
+Finally, we select a sample of $104$ localized FRBs (see the Supplemental Material SM (see also Refs.
+Chatterjee et al.
+(2017); Tendulkar et al.
+(2017); Bhandari et al.
+(2022); Bannister et al.
+(2019); Prochaska et al.
+(2019); Bhardwaj et al.
+(2024); Bhandari et al.
+(2020); Ibik et al.
+(2024); Michilli et al.
+(2023); Ravi et al.
+(2019); Chittidi et al.
+(2021); Heintz et al.
+(2020); Law et al.
+(2020); Shin et al.
+(2025); Ravi et al.
+(2022); Shannon et al.
+(2025); Caleb et al.
+(2023); Cassanelli et al.
+(2024); Gordon et al.
+(2023); Sharma et al.
+(2024); Law et al.
+(2024); Li et al.
+(2026); Ryder et al.
+(2023); Rajwade et al.
+(2024); Anna-Thomas et al.
+(2025); Hanmer et al.
+(2025); Connor et al.
+(2025); Eftekhari et al.
+(2025); Amiri et al.
+(2025); Tian et al.
+(2024) therein).
+The dataset is selected from an initial set of $119$, based on the following criteria: (1) $\mathrm{DM}_{\mathrm{MW,ISM}}$ is required to be less than $40\%$ of $\mathrm{DM}_{\rm obs}$ Connor et al.
+(2025); (2) $\mathrm{DM}_{\rm obs}-\rm DM_{MW,ISM}$ is taken to exceed 80 $\rm{pc}\,\rm{cm}^{-3}\,$, ensuring a positive $\mathrm{DM}_{\rm host}$; (3) FRB 20190520B and FRB 20220831A are excluded due to excess local DM contributions from their host environments.
+GW standard sirens.
+We use 47 GW standard sirens with SNRs greater than 11 and an inverse false alarm rate exceeding 4 yr from GWTC-3111https://gwosc.org/GWTC-3, covering a redshift range of approximately $z<0.8$.
 The sample consists of 42 BBHs, three neutron star-black holes (NSBHs), and two BNSs.
-We do not consider GW events detected in the LVK O4a run, mainly because during this run, Virgo is not online, and all GW events are detected by two LIGOs, resulting in relatively poorer sky localizations and luminosity distance measurements LIGOScientific:2025jau.
-We obtain the redshift of GW170817 from the electromagnetic (EM) counterpart LIGOScientific:2017adf; LIGOScientific:2017vwq; LIGOScientific:2017zic.
-For GW events without EM counterparts, we infer their redshifts through the dark siren method Schutz:1986gp; Mastrogiovanni:2023emh; Gray:2023wgj and incorporate the GLADE+ catalog Dalya:2021ewn.
+We do not consider GW events detected in the LVK O4a run, mainly because during this run, Virgo is not online, and all GW events are detected by two LIGOs, resulting in relatively poorer sky localizations and luminosity distance measurements Abac et al.
+(2025).
+We obtain the redshift of GW170817 from the electromagnetic (EM) counterpart Abbott et al.
+(2017a, b, c).
+For GW events without EM counterparts, we infer their redshifts through the dark siren method Schutz (1986); Mastrogiovanni et al.
+(2023); Gray et al.
+(2023) and incorporate the GLADE+ catalog Dálya et al.
+(2022).
 In the dark siren method, assuming the GW sources all originate from galaxies, we can cross-match sky localizations of GW events with galaxies collected in galaxy catalogs, and identify a serious potential host galaxy.
 These potential host galaxies provide redshift priors for GW events, which help to infer cosmological parameters in a hierarchical Bayesian framework.
-Furthermore, GW population models can help break degeneracies of masses and redshifts in GW detections, thereby providing additional redshift information Taylor:2011fs; Ezquiaga:2022zkx; Mastrogiovanni:2023emh; Gray:2023wgj.
-In this paper, we use the public pipeline ICAROGW Mastrogiovanni:2023zbw to compute the GW likelihood function, denoted as $\mathcal{L}_{\rm GW}$.
-The explicit formulas can be found in Refs Mastrogiovanni:2023emh; Gray:2023wgj.
+Furthermore, GW population models can help break degeneracies of masses and redshifts in GW detections, thereby providing additional redshift information Taylor et al.
+(2012); Ezquiaga and Holz (2022); Mastrogiovanni et al.
+(2023); Gray et al.
+(2023).
+In this paper, we use the public pipeline ICAROGW Mastrogiovanni et al.
+(2024) to compute the GW likelihood function, denoted as $\mathcal{L}_{\rm GW}$.
+The explicit formulas can be found in Refs Mastrogiovanni et al.
+(2023); Gray et al.
+(2023).
 We adopt the same population model assumptions as Ref.
-LIGOScientific:2021aug, but restrict our analysis to the Power Law + Peak BBH mass model, which is best supported by the GWTC-3 data.
+Abbott et al.
+(2023b), but restrict our analysis to the Power Law + Peak BBH mass model, which is best supported by the GWTC-3 data.
 Considering additional models would increase computational cost and lies beyond the scope of this work.
-For BNSs, we assume that neutron star masses are uniformly distributed in the range of 1–3 $M_{\odot}$.
+For BNSs, we assume that neutron star masses are uniformly distributed in the range of 131-3 $M_{\odot}$.
 In the case of NSBH systems, the black hole component is drawn from the same distribution as the primary black holes in BBHs, while the neutron star component follows the mass distribution assumed for BNSs.
@@
 Consistent with Ref.
-LIGOScientific:2021aug, we disregard potential redshift evolution in the BBH mass distribution as well as variations in the GW detection rate arising from the spin distribution.
+Abbott et al.
+(2023b), we disregard potential redshift evolution in the BBH mass distribution as well as variations in the GW detection rate arising from the spin distribution.
 These effects are currently subdominant compared to the statistical uncertainties present in the existing GW observations.
-II.3 Joint analysis of FRB and GW data
-Figure 1: Normalized posterior distributions of $\Omega_{\rm b}$~from 104 FRBs (orange) and 104 FRBs + 47 GWs (indigo).
-The median and $1\sigma$ credible interval from FRB + GW is $\Omega_{\rm b}=0.0488\pm 0.0064$, marginalized over $f_{\rm d}$ and empirical DM model parameters.
-The errorbars in different colors and linestyles are shown under different assumptions: FRB + $H_{0}$ (Planck18 prior) and FRB + $H_{0}$ (SH0ES prior) adopt Gaussian priors on $H_{0}$ from Planck18 2020Planck and SH0ES Riess:2021jrx, respectively; FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW adopt priors on empirical DM model parameters from the IllustrisTNG simulation Zhang:2020mgq; Zhang:2020xoc and on $f_{\rm d}$ from the redshift-dependent form of Macquart20 Macquart:2020lln, respectively.
-The blue, gray, and red shaded bands show the results from CMB + BBN Cooke:2017cwo, Macquart20, and Yang22 Yang:2022ftm, respectively.
-Note that Yang22 adopts the TNG prior, a flat prior on $H_{0}$ $\in$ $\mathcal{U}(67,73)$ km s-1 Mpc-1, and $f_{\rm d}=0.82$.
-All results are at the $1\sigma$ level.
-We jointly constrain cosmological and FRB-related DM parameters, and adopt flat priors for the following seven parameters: $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ Yang:2022ftm, $\Omega_{\rm b}\in\mathcal{U}(0.01,0.10)$, $H_{0}\in\mathcal{U}(20,140)~\rm km\ s^{-1}\ Mpc^{-1}$, $f_{\rm d}\in\mathcal{U}(0.84,0.96)$, [^3] This prior on $f_{\rm d}$ is informed by observations of the baryon fraction in stars ($f_{*}$) and cold gas ($f_{\rm cold}$), independent of FRBs.
-Assuming a conservative flat prior for $f_{*}$ extended by a factor of five, the diffuse baryon fraction is inferred via $f_{\rm d}=1-f_{*}-f_{\rm cold}$ connor2025gas.
-$F\in\mathcal{U}(0.011,1.0)$, $\mu_{\rm host}\in\mathcal{U}(3.0,5.3)$, and $\sigma_{\rm host}\in\mathcal{U}(0.2,2.0)$ Macquart:2020lln.
-For the GW population model parameters, we consider two scenarios.
-In the first, following Ref.
-LIGOScientific:2021aug, we fix all population-model parameters to their fiducial values.
-In the second, we allow the BBH mass-distribution parameters and the redshift-distribution parameters of GW sources to vary freely, adopting the same prior ranges as the spectral siren analysis in Ref.
-LIGOScientific:2021aug, to assess the impact of population-model uncertainties on our results.
+Joint analysis of FRB and GW data.
+We jointly constrain cosmological and all DM parameters with flat priors for the following seven parameters: $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ Yang et al.
+(2022), $\Omega_{\rm b}\in\mathcal{U}(0.01,0.10)$, $H_{0}\in\mathcal{U}(20,140)~\rm km\ s^{-1}\ Mpc^{-1}$, $f_{\rm d}\in\mathcal{U}(0.84,0.96)$, [^2] This prior on $f_{\rm d}$ is informed by observations of the baryon fraction in stars ($f_{*}$) and cold gas ($f_{\rm cold}$), independent of FRBs.
+Assuming a conservative flat prior for $f_{*}$ extended by a factor of five, the diffuse baryon fraction is inferred via $f_{\rm d}=1-f_{*}-f_{\rm cold}$ Connor et al.
+(2025).
+$F\in\mathcal{U}(0.011,1.0)$, $\mu_{\rm host}\in\mathcal{U}(3.0,5.3)$, and $\sigma_{\rm host}\in\mathcal{U}(0.2,2.0)$ Macquart et al.
+(2020).
+To test the robustness of our analysis, we examine the effect of prior assumptions on our results.
+For the $\Omega_{\rm m}$ prior, we relax it to a broad range $\mathcal{U}(0,1)$ to assess its impact on the constraints.
+For the GW population-model parameters, following Ref.
+Abbott et al.
+(2023b), we first fix all parameters to their fiducial values and then, vary the BBH mass- and redshift-distribution parameters of GW sources, adopting the same prior ranges as in the spectral-siren analysis of Ref.
+Abbott et al.
+(2023b), to assess the impact of population-model uncertainties on our results.
 This treatment is similar to Ref.
-Gray:2023wgj; however, unlike Ref.
-LIGOScientific:2025jau, we do not free the mass-distribution parameters of neutron-star sources.
+Gray et al.
+(2023); however, unlike Ref.
+Abac et al.
+(2025), we do not free the mass-distribution parameters of neutron-star sources.
 We utilize the bilby [^3] https://github.com/bilby-dev/bilby package to perform Markov Chain Monte Carlo (MCMC) analysis and choose the emcee [^4] https://emcee.readthedocs.io/en/stable sampler.
@@
 Table 1: $1\sigma$ constraints on cosmological parameters $\Omega_{\rm b}$ and $H_{0}$.
-The results are derived from $104$ localized FRBs alone or in combination with $47$ GW events, under different assumptions.
-For all cases, the $1\sigma$ constraints on DM parameters are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$.
+The results are derived from 104104 localized FRBs alone or in combination with $47$ GW events, under different assumptions.
 Unless otherwise specified, the FRB + GW case assumes a flat prior of $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ and fixed GW population parameters.
@@
 FRB $0.0520^{+0.0120}_{-0.0280}$ $77.0^{+20.0}_{-40.0}$
-    FRB + GW     $0.0488\pm 0.0064$     $71.6^{+4.4}_{-8.6}$
-    FRB + GW ($\Omega_{\rm m}\in\mathcal{U}(0,1)$)     $0.0454^{+0.0066}_{-0.0082}$     $72.7^{+4.7}_{-9.0}$
-    FRB (TNG) + GW     $0.0491\pm 0.0054$     $71.2^{+4.4}_{-8.0}$
-    FRB (TNG + $f_{\rm d}(z)$) + GW     $0.0510^{+0.0055}_{-0.0050}$     $71.8^{+4.5}_{-8.6}$
 FRB + $H_{0}$ (Planck18 prior) $0.0513\pm 0.0046$ $67.4\pm 0.5$
 FRB + $H_{0}$ (SH0ES prior) $0.0473\pm 0.0043$ $73.0\pm 1.0$
+FRB + $\Omega_{\rm b}h^{2}$ (BBN prior) $66.2^{+5.3}_{-7.3}$
+FRB + GW $0.0488\pm 0.0064$ $71.6^{+4.4}_{-8.6}$
 FRB + GW (GW parameters free) $0.0448^{+0.0094}_{-0.0076}$ $79.9^{+5.0}_{-17.0}$
-III Results and discussions
-In this section, we first report the constraints on the baryon density as our main results in Sec.
-III.1, and then extend the analysis to assess the impacts of FRB and GW population modeling assumptions in Secs.
-III.2 and III.3, respectively.
-III.1 Main results
-First, we report the main results for the baryon-density estimate in the first block of Table 1.
+FRB + GW ($\Omega_{\rm m}\in\mathcal{U}(0,1)$) $0.0454^{+0.0066}_{-0.0082}$ $72.7^{+4.7}_{-9.0}$
+Results.
+Figure 1: Posteriors for $\Omega_{\rm b}$~from 104 FRBs (orange) and 104 FRBs + 47 GWs (indigo).
+The joint constraint is $\Omega_{\rm b}=0.0488\pm 0.0064$, obtained by marginalized over all DM parameters.
+Errorbars in different colors and linestyles are shown under different assumptions: FRB + $H_{0}$ (Planck18 prior) and FRB + $H_{0}$ (SH0ES prior) adopt Gaussian priors on $H_{0}$ from Planck18 Aghanim et al.
+(2020) and SH0ES Riess et al.
+(2022), respectively; Shaded bands in different colors show the results from CMB + BBN Cooke et al.
+(2018), Macquart20, and Yang22 Yang et al.
+(2022).
+Note that Yang22 adopts the prior on DM model parameters from the TNG simulation Zhang et al.
+(2020, 2021), a flat prior on $H_{0}$ $\in$ $\mathcal{U}(67,73)$ km s-1 Mpc-1, and $f_{\rm d}=0.82$.
+All results are at the $1\sigma$ level.
+Figure 2: Posteriors ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm m}$, $H_{0}$, and the FRB DM parameter $f_{\rm d}$, $F$, $\sigma_{\rm host}$, $\mu_{\rm host}$, and the typical GW population parameter $\mu_{\rm g}$ from FRB + GW data.
+Two cases are shown: GW parameters fixed within the PowerLaw + Peak model (indigo) and the parameters allowed to vary freely and marginalized over (gray).
+The dotted lines indicate the median values of each posterior, except for $\mu_{\rm g}$ in the GW-parameter-fixed case, where the dotted line marks the prior value $\mu_{\rm g}=32.27~M_{\odot}$ adopted following Ref.
+Abbott et al.
+(2023b).
+We report the main results for the baryon-density estimate in the first block of Table 1.
 The posteriors for $\Omega_{\rm b}$~ and the corresponding credible intervals are shown in Fig.
 1.
+Figure 2 shows the joint posterior distributions of the cosmological and DM parameters.
 Using FRB data alone yields a relatively weak constraint of $\Omega_{\mathrm{b}}=0.052_{-0.028}^{+0.012}$ at the $1\sigma$ level.
 Incorporating GW standard sirens significantly narrows this to $\Omega_{\mathrm{b}}=0.0488\pm 0.0064$, by breaking the strong degeneracy between $\Omega_{\mathrm{b}}$ and $H_{0}$ via the independent $H_{0}$ measurement provided by GWs.
-Remarkably, this late-Universe estimate exhibits concordance with early-Universe values from BBN and CMB at the $\sim 13\%$ level, linking the gap between the early and late Universe.
-Note that a similarly precise constraint is simultaneously obtained for the Hubble constant, $H_{0}=71.6^{+4.4}_{-8.6}~\mathrm{kms^{-1}~Mpc^{-1}}$, primarily driven by the GW dataset alone, which gives $H_{0}=72.6^{+5.2}_{-8.9}~\mathrm{kms^{-1}~Mpc^{-1}}$.
-More data is required to arbitrate the $H_{0}$ tension.
-A detailed discussion on $H_{0}$ and DM parameters is provided in Appendix A.
+The $\Omega_{\mathrm{b}}$-$H_{0}$ panel in Fig.
+2 shows that this correlation is markedly stronger than those between $\Omega_{\mathrm{b}}$ and the other parameters.
+This late-Universe estimate exhibits remarkable concordance with early-Universe values from BBN and CMB at the $\sim 13\%$ level.
+For the FRB DM parameters, we obtain $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$, corresponding to a median rest-frame host DM of $108\pm 20~\mathrm{pc~cm^{-3}}$ and a mean host DM of $168^{+57}_{-40}~\mathrm{pc~cm^{-3}}$, consistent with previous estimates Yang et al.
+(2025); Connor et al.
+(2025).
 We compare our $\Omega_{\rm b}$~results with those from previous works.
@@
 1.
-This suggests that the $H_{0}$ tension would lead to two biased $\Omega_{\rm b}$~estimates.
+Obviously, the $H_{0}$ tension would lead to two biased $\Omega_{\rm b}$~estimates.
 Although the induced $\Omega_{\rm b}$ inconsistency” in our sample is within $1\sigma$ uncertainty ($\sim 0.95\sigma$), it may become non-negligible with larger samples in updated DM analyses.
-Nevertheless, the joint FRB + GW result assumes no prior on $H_{0}$ and instead provides a late-Universe measurement of baryons, independent of both CMB and local distance ladder.
-This FRB + GW synergy may establish a novel paradigm for the cosmic baryon census, analogous to CMB + BBN, and serve as a powerful low-redshift cosmological probe in the era of next-generation FRB and GW surveys Zhang:2023gye.
+Nevertheless, the joint FRB + GW result assumes no prior on $H_{0}$ and instead provides a late-Universe measurement of baryons, independent of both CMB and the local distance ladder.
+This FRB + GW synergy may establish a novel paradigm for the cosmic baryon census, analogous to CMB + BBN, and serve as a powerful low-redshift cosmological probe in the era of next-generation FRB and GW surveys Zhang et al.
+(2023).
+We assess the impact of GW population modeling on our baryon census.
+In our main analysis, no FRB population model is imposed, whereas the PowerLaw + Peak GW population model is adopted for the GW population, as is commonly used in GW-based cosmological inference Abbott et al.
+(2023b).
+To investigate the robustness of our $\Omega_{\rm b}$~estimate against this assumption, we release the GW population model parameters and marginalize over them.
+We find that joint FRB + GW analysis with GW model parameters free yields a higher value of $H_{0}=79.9^{+5.0}_{-17.0}$ km s-1  Mpc-1and a correspondingly lower value of $\Omega_{\rm b}=0.0448^{+0.0094}_{-0.0076}$.
+The $\Omega_{\rm b}$ result ($\sim 20\%$ precision) remains broadly consistent with the CMB + BBN determination at the $1\sigma$ level.
+To illustrate how the GW population prior affects $\Omega_{\rm b}$, we focus on the parameter $\mu_{\rm g}$, which denotes the mean of the Gaussian component in the primary mass distribution within the PowerLaw + Peak model.
+This parameter is closely correlated with the inferred $H_{0}$ from GW observations Abbott et al.
+(2023b), making it suitable for accounting for the impact of population modeling.
+As shown in Fig.
+2, $\mu_{\rm g}$ exhibits a negative degeneracy with $H_{0}$ and a corresponding positive degeneracy with $\Omega_{\rm b}$.
+Consequently, compared with the prior value $\mu_{\rm g}=32.27~M_{\odot}$ adopted in the GW-parameter-fixed (indigo) case, the lower inferred value of $\mu_{\rm g}=31.5~M_{\odot}$ in the GW-parameter-free (gray) case leads to a higher $H_{0}$ and thus a lower $\Omega_{\rm b}$.
+Thus, to obtain a robust estimate of $\Omega_{\rm b}$, caution is required when adopting GW-population assumptions, due to their strong impact on the inferred $H_{0}$.
+Nevertheless, our test without imposing any GW population model reinforces that our fiducial result obtained with the PowerLaw + Peak model is reliable and not significantly biased.
 We also assess the dependence on the prior for $\Omega_{\rm m}$.
-Relaxing the prior from $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ to a broad flat prior $\Omega_{\rm m}\in\mathcal{U}(0,1)$ yields $\Omega_{\rm b}=0.0454^{+0.0066}_{-0.0082}$ and $H_{0}=72.7^{+4.7}_{-9.0}$ km s-1  Mpc-1, as shown in the third row of Table 1.
+Relaxing the prior from $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ to a broad flat prior $\Omega_{\rm m}\in\mathcal{U}(0,1)$ yields $\Omega_{\rm b}=0.0454^{+0.0066}_{-0.0082}$ and $H_{0}=72.7^{+4.7}_{-9.0}$ km s-1  Mpc-1.
 Compared with the fiducial case ($\Omega_{\rm b}=0.0488\pm 0.0064$), the slightly lower value of $\Omega_{\rm b}$ is driven by the biased constraint on $\Omega_{\rm m}$ under the broad prior, $\Omega_{\rm m}=0.225^{+0.060}_{-0.180}$, which propagates through the positive correlation between $\Omega_{\rm b}$~and $\Omega_{\rm m}$.
 Therefore, we caveat that our present analysis moderately depends on the $\Omega_{\rm m}$ prior, which can be improved with larger future FRB and GW samples to better constrain $\Omega_{\rm m}$.
-III.2 Impact of FRB model parameters
-Figure 2: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm b}$ and $\Omega_{\rm b}h_{70}$ versus $H_{0}$, $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$, using FRB (orange) and FRB + GW (indigo) data.
-The correlation coefficient $r$ is labeled in each panel to quantify parameter degeneracies.
-Note that $H_{0}$ is in units of km s-1  Mpc-1.
-Next, we assess the impact of FRB population modeling on the baryon census, as listed in the second block of Table 1.
-In our fiducial analysis in Sec.
-III.1, no specific FRB population model is imposed and all DM parameters are left free to avoid potential bias.
-To explore how such modeling affects the $\Omega_{\rm b}$ inference, we apply several commonly used FRB population models and evaluate their potential to improve the precision of $\Omega_{\rm b}$.
-FRB cosmology is subject to uncontrolled systematics mainly from modeling $\mathrm{DM}_{\mathrm{IGM}}$ and $\mathrm{DM}_{\mathrm{host}}$ distributions, as well as the weakly constrained $f_{\rm d}$, which affects the mean value of $\mathrm{DM}_{\mathrm{IGM}}$.
-To reduce these systematics, cosmological simulations are commonly used to calibrate these DM parameters ($F$, $\mu_{\rm host}$, $\sigma_{\rm host}$, and $f_{\rm d}$) at different redshifts.
-Accordingly, we consider two cases with simulation-based priors: (i) fixing the empirical DM model parameters to the best-fit results from the IllustrisTNG simulation (hereafter TNG) Zhang:2020mgq; Zhang:2020xoc; and (ii) further fixing $f_{\rm d}$ to a redshift-dependent form following Macquart20, as implemented in the public FRB code [^7] https://github.com/FRBs/FRB.
-The constraints from FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW are $\Omega_{\mathrm{b}}=0.0491\pm 0.0054$ and $\Omega_{\mathrm{b}}=0.0507\pm 0.0052$, respectively, with the corresponding errorbars shown in Fig.
-1.
-Both results are consistent with the FRB + GW constraint within the $1\sigma$ uncertainty, suggesting that the simulation-based priors can improve current constraints without introducing significant bias.
-However, such agreement at the current level does not ensure the absence of systematic errors when applied to larger future FRB and GW samples, in which systematic errors may become dominant.
-Therefore, a detailed assessment of how the DM parameters affect the inferred baryon density is of great importance.
-To assess the impact of systematic uncertainties on $\Omega_{\mathrm{b}}$ and verify the robustness of our constraint, we analyze the parameter correlations using the FRB + GW dataset.
-The lower row of Fig.
-2 shows the marginalized posteriors of $\Omega_{\rm b}$~and other relevant parameters, with correlation coefficients $r$ labeled in each panel to quantify parameter degeneracies.
-On the one hand, the degeneracy between $\Omega_{\rm b}$ and $H_{0}$ is the strongest among all parameter pairs (r=0.73r=-0.73), implying that the inferred value of $\Omega_{\rm b}$ is highly sensitive to how $H_{0}$ is determined.
-While FRB + $H_{0}$ analyses rely on external priors, which are controversial due to the $H_{0}$ tension, FRB + GW analyses rely on GW data to provide $H_{0}$, which is uniquely derived from self-calibrating absolute distance measurements.
-Therefore, the $\Omega_{\rm b}$ constraint from FRB + GW is robust against the $H_{0}$-related bias and free from assumptions affected by the $H_{0}$ tension.
-On the other hand, the correlations between $\Omega_{\rm b}$~and DM parameters are moderate.
-These DM parameters are not properly constrained, both in this work (see Appendix A for details) and earlier studies James:2022dcx, and thus act as nuisance parameters that bias the inferred $\Omega_{\rm b}$.
-In the joint FRB + GW analysis, the correlation coefficients between $\Omega_{\rm b}$ and $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$ are 0.28-0.28, $0.40$, 0.40-0.40, and $0.17$, respectively.
-For comparison, the upper row of Fig.
-2 shows the constraint on $\Omega_{\rm b}h_{70}$, where $h_{70}=H_{0}/(70~\mathrm{km~s^{-1}~Mpc^{-1}})$.
-This parameter absorbs the strong degeneracy, enabling it to be effectively constrained using FRB data alone Macquart:2020lln; connor2025gas.
-In this case, the corresponding correlations are notably stronger: 0.41-0.41, $0.57$, 0.62-0.62, and $0.27$, respectively.
-This comparison suggests that directly measuring $\Omega_{\rm b}$ reduces degeneracies with the nuisance parameters and is more robust against the associated systematic uncertainties than $\Omega_{\rm b}h_{70}$.
-We conclude that the impact of FRB population modeling on the inferred $\Omega_{\rm b}$ is negligible within our analysis.
-The results also show that, within our FRB + GW framework, incorporating GW data into the analysis of FRB dispersion measures enables a direct and effective measurement of $\Omega_{\rm b}$ and, more importantly, avoids biases from the $H_{0}$ tension and mitigates those associated with nuisance model parameters.
-This demonstrates that the synergy between FRBs and GWs provides a robust estimate of the cosmic baryon density.
-III.3 Impact of GW model parameters
-Figure 3: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm m}$, $H_{0}$, and the typical GW population parameter $\mu_{g}$ using FRB + GW data.
-Two cases are shown: GW parameters fixed within the PowerLaw + Peak model (indigo) and the parameters allowed to vary freely and marginalised over (gray).
-The dotted lines indicate the median values of each posterior, except for $\mu_{g}$ in the GW-parameter-fixed case, where the dotted line marks the prior value $\mu_{g}=32.27$ adopted following Ref.
-LIGOScientific:2021aug.
-Note that $H_{0}$ is in units of km s-1  Mpc-1.
-Last, we assess the impact of GW population modeling on our baryon census, as listed in the third block of Table 1.
-In our fiducial analysis in Sec.
-III.1, no FRB population model is imposed, whereas the PowerLaw + Peak GW population model is adopted for the GW population, as is commonly used in GW-based cosmological inference LIGOScientific:2021aug.
-To investigate the robustness of our $\Omega_{\rm b}$~estimate against this assumption, we release the GW population model parameters and marginalize over them.
-We find that joint FRB + GW analysis with GW model parameters free yields a higher value of $H_{0}=79.9^{+5.0}_{-17.0}$ km s-1  Mpc-1and a correspondingly lower value of $\Omega_{\rm b}=0.0448^{+0.0094}_{-0.0076}$.
-The $\Omega_{\rm b}$ result ($\sim 20\%$ precision) remains consistent with the CMB + BBN determination at the $1\sigma$ level.
-To illustrate how the GW population prior affects $\Omega_{\rm b}$, we focus on the parameter $\mu_{g}$, which denotes the mean of the Gaussian component in the primary mass distribution within the PowerLaw + Peak model.
-This parameter is moderately correlated with the inferred $H_{0}$ from GW observations LIGOScientific:2021aug, making it suitable for accounting for the impact of population modeling.
-As shown in Fig.
-3, $\mu_{g}$ exhibits a negative degeneracy with $H_{0}$ and a corresponding positive degeneracy with $\Omega_{\rm b}$.
-Consequently, compared with the prior value $\mu_{g}=32.27$ adopted in the GW–parameter-fixed (indigo) case, the lower inferred value of $\mu_{g}=31.5$ in the GW–parameter-free (gray) case leads to a higher $H_{0}$ and thus a lower $\Omega_{\rm b}$.
-Thus, within our FRB + GW framework, the inferred $\Omega_{\rm b}$ is more sensitive to GW-related parameters than to FRB-related ones (see Sec.
-III.2).
-It is because $H_{0}$, the essential parameter to infer $\Omega_{\rm b}$ in our framework, is more strongly degenerate with the GW population parameters than with the FRB-related parameters.
-We conclude that, to obtain an unbiased and robust estimate of $\Omega_{\rm b}$, caution is required when adopting GW-population assumptions, due to their strong impact on the inferred $H_{0}$.
-Nevertheless, our test without imposing any GW population model reinforces that our fiducial result obtained with the PowerLaw + Peak model is reliable and not significantly biased.
-IV Conclusion
-In this paper, we combine two emerging cosmological probes, i.e., $104$ localized FRBs and $47$ GW standard sirens, within a unified framework to measure the cosmic baryon density in a $H_{0}$-free approach.
-Most previous baryon censuses assumed a fixed value of $H_{0}$ or adopted an external prior, which could bias the inference in the presence of the Hubble tension.
-If FRB DMs are instead used to trace the baryons, additional parameter degeneracies arise in the DM modeling, particularly between $H_{0}$ and $\Omega_{\rm b}$.
+Figure 3: Same as Fig.
+1 but for posteriors for $H_{0}$.
+The joint FRB + GW constraint is $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1.
+The orange and black errorbars correspond to constraints from FRB + $\Omega_{\rm b}h^{2}$ (BBN prior Cooke et al.
+(2018)) and GW data alone, respectively.
+The green, yellow, red, and gray shaded bands denote the results from Planck18 Aghanim et al.
+(2020), SH0ES Riess et al.
+(2022), Refs.
+Wu et al.
+(2022) (Wu22), and James et al.
+(2022) (James22), respectively.
+Note that Wu22 adopts the TNG prior, and both Wu22 and James22 adopt $\Omega_{\rm b}h^{2}$ priors from BBN Cooke et al.
+(2018) and CMB Aghanim et al.
+(2020), respectively.
+All results are at $1\sigma$ uncertainty.
+Last, the corresponding constraints on $H_{0}$ are also listed in Table 1, with the posteriors and credible intervals shown in Fig.
+3.
+A direct measurement of $H_{0}$ using FRBs alone is difficult due to strong parameter degeneracies.
+Concretely, our FRB-only measurement gives $H_{0}=77.0^{+20.0}_{-40.0}$ km s-1  Mpc-1, while including GW data narrows the constraint to $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1.
+This improvement is primarily driven by the GW dataset itself, which yields $H_{0}=72.6^{+5.2}_{-8.9}$ km s-1  Mpc-1.
+Previous FRB studies commonly inferred $H_{0}$ by fixing the baryon density to early-Universe values.
+For example, Refs.
+Wu et al.
+(2022) and James et al.
+(2022) inferred $H_{0}$ from different FRB samples by adopting external priors on $\Omega_{\rm b}h^{2}$, namely $\Omega_{\rm b}h^{2}\sim\mathcal{N}(0.02235,0.00049)$ from BBN and $\Omega_{\rm b}h^{2}=0.02242$ from Planck18, respectively, as illustrated by the red and gray bands in Fig.
+3.
+Using the same BBN prior, we obtain $H_{0}=66.2^{+5.3}_{-7.3}$ km s-1  Mpc-1 from FRB + $\Omega_{\rm b}h^{2}$, with a precision comparable to that of FRB + GW.
+Unlike these prior-based methods, the FRB + GW constraint provides a late-Universe determination of $H_{0}$.
+More data is required to offer further insight into arbitrating the $H_{0}$ tension.
+Conclusion.
+In this Letter, we combine two emerging cosmological probes, i.e., $104$ localized FRBs and $47$ GW standard sirens, within a unified framework to measure the cosmic baryon density in a $H_{0}$-free approach.
+Previous baryon censuses often fixed $H_{0}$ or used external priors, which could bias the inference under the Hubble tension.
+If FRB DMs are employed to trace the baryons, serious parameter degeneracies arise in the DM modeling, particularly between $H_{0}$ and $\Omega_{\rm b}$.
 GWs independently infer $H_{0}$ from self-calibrated luminosity distances obtained through waveform analysis as sirens, thus breaking the $\Omega_{\rm b}$-$H_{0}$ degeneracy.
@@
 Together, these results exhibit $\Omega_{\rm b}$-$H_{0}$ concordance between the early and late Universe.
-Moreover, we explore two prior-based scenarios: (i) $H_{0}$-prior-based cases using FRB data alone, where adopting Gaussian priors from Planck18 and SH0ES yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively.
-These results indicate that the $H_{0}$ tension would bias baryon density estimates, although the induced $\Omega_{\rm b}$~inconsistency” remains within $1\sigma$ uncertainty in our sample.
-(ii) Simulation-based prior cases using FRB + GW, where fixing empirical DM model parameters from the IllustrisTNG simulation and further fixing $f_{\rm d}$ to a redshift-dependent form yield $\Omega_{\rm b}=0.0491\pm 0.0054$ and $0.0507\pm 0.0052$, respectively.
-These two results agree with that of FRB + GW at the $1\sigma$ level, indicating that simulation-based priors can improve current constraints without introducing significant bias.
-However, in both (i) and (ii), systematic effects from external assumptions (related to $H_{0}$, $f_{\rm d}$, and empirical DM model parameters) may become non-negligible when applied to larger future FRB and GW samples.
-Hence, we assess the associated systematic uncertainties through a detailed correlation analysis.
-On the one hand, the dominant uncertainty is the degeneracy between $\Omega_{\rm b}$ and $H_{0}$, implying the importance of carefully addressing the $H_{0}$-related bias.
-Unlike prior-based analyses that rely on assumed values of $H_{0}$, the joint FRB + GW analysis avoids such bias by uniquely inferring $H_{0}$ from the absolute distances provided by GWs, independent of both CMB and local distance ladder.
-On the other hand, the correlations between $\Omega_{\rm b}$ and DM parameters—treated as nuisance parameters—are moderate, and notably weaker than those of $\Omega_{\rm b}h_{70}$, making the direct inference of $\Omega_{\rm b}$ more robust against the uncertainties induced by the nuisance parameters.
-Therefore, combining FRB and GW data not only avoids the bias induced by the $H_{0}$ tension, but also mitigates model-dependent uncertainties.
-We caveat that our baryon census still depends on the adopted $\Omega_{\rm m}$ prior and the assumed GW population model.
-Nevertheless, the $\Omega_{\rm b}$ values inferred when removing these assumptions show only modest shifts and remain consistent with CMB + BBN values at the $1\sigma$ level.
+We note that our baryon census still depends on the adopted $\Omega_{\rm m}$ prior and the assumed GW population model.
+Nevertheless, the $\Omega_{\rm b}$ values inferred when removing these assumptions show only modest shifts and remain consistent with CMB + BBN determinations at the $1\sigma$ level.
 Using larger samples of localized FRBs and GW standard sirens, future analyses would jointly constrain $\Omega_{\rm m}$, GW population parameters, and $\Omega_{\rm b}$ more robustly, thereby reducing this dependence.
 Overall, the novel combination of FRBs and GWs enables a direct, unique, and robust measurement of both the baryon content and cosmological distances at low redshifts.
-With an increasing number of GW detections expected in the upcoming LVK run, and potentially thousands of localized FRBs expected from CHIME/Outriggers masui2023pinpointing and DSA-110 Law:2023ibd, their synergy will soon emerge as a powerful low-redshift cosmological probe.
-V Acknowledgments
+With an increasing number of GW detections expected in the upcoming LVK run, and potentially thousands of localized FRBs expected from CHIME/Outriggers Masui (2023) and DSA-110 Law et al.
+(2024), their synergy will soon emerge as a powerful low-redshift cosmological probe.
+Acknowledgments.
 We thank Tian-Nuo Li, Meng-Lin Zhao, Yang Liu, and Qin Wu for fruitful discussions.
@@
 N2405008).
-Appendix A Constraints on $H_{0}$ and DM parameters
-Figure 4: Same as Fig.
-1 but for posterior distribution of $H_{0}$.
-The median and $1\sigma$ credible interval from FRB + GW is $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1.
-The orange and black errorbars correspond to constraints from FRB + $\Omega_{\rm b}h^{2}$ (BBN prior Cooke:2017cwo) and GW data alone, respectively.
-The green, yellow, red, and gray shaded bands denote the results from Planck18 2020Planck, SH0ES Riess:2021jrx), Wu22 Wu:2021jyk, and James22 James:2022dcx, respectively.
-Note that Wu22 adopts the TNG prior, and both Wu22 and James22 adopt $\Omega_{\rm b}h^{2}$ priors from BBN Cooke:2017cwo and CMB 2020Planck, respectively.
-All results are at $1\sigma$ uncertainty.
-In this Appendix, we report and discuss the constraints on $H_{0}$ and DM parameters ($f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$).
-The constraints on $H_{0}$ are listed in Table 1, and posteriors and corresponding credible intervals for $H_{0}$ are shown in Fig.
-4.
-A direct measurement of $H_{0}$ using FRBs alone is difficult due to strong parameter degeneracies.
-Specifically, our measurement using FRBs alone is $H_{0}=77.0^{+20.0}_{-40.0}$ km s-1  Mpc-1, and the inclusion of GW data narrows the constraint to $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1.
-This improvement is mainly driven by the GW data itself, which yields $H_{0}=72.6^{+5.2}_{-8.9}$ km s-1  Mpc-1.
-Previous studies predominantly inferred $H_{0}$ by fixing the baryon density parameter at the early-Universe value.
-For example, Refs.
-Wu:2021jyk (hereafter Wu22) and James:2022dcx (hereafter James22) used different FRB samples to derive $H_{0}=69.31^{+6.21}_{-6.63}~\mathrm{km~s^{-1}~Mpc^{-1}}$ and $H_{0}=73^{+12}_{-8}~\mathrm{km~s^{-1}~Mpc^{-1}}$, by assuming external priors on $\Omega_{\rm b}h^{2}\sim\mathcal{N}(0.02235,0.00049)$ from BBN and $\Omega_{\rm b}h^{2}=0.02242$ from Planck18, respectively, which are shown as red and gray bands in Fig.
-4.
-Recently, Refs.
-Wang:2025ugc and Gao:2025fcr also used $92$ and $108$ localized FRBs to obtain $H_{0}=69.04_{-2.07}^{+2.30}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$ and $H_{0}=69.40_{-1.97}^{+2.14}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$, respectively, both assuming BBN and TNG priors.
-Using the same BBN prior, we obtain $H_{0}=66.2^{+5.3}_{-7.3}$ km s-1  Mpc-1from FRB + $\Omega_{\rm b}h^{2}$, which slightly favors the Planck18 result with a precision comparable to that of FRB + GW.
-Unlike prior-based methods, the FRB + GW constraint provides a late-Universe result of $H_{0}$.
-More data is required to offer further insight into arbitrating the $H_{0}$ tension.
-Notably, time-delay cosmography using strongly lensed FRBs by massive galaxies can offer an independent measurement of $H_{0}$, since the time delays between lensed FRB images can be determined with exceptional precision Li:2017mek; Zhang:2024rra.
-For DM parameters, the constraints are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$.
-The inclusion of GW data has a negligible impact on this constraint.
-For the host contribution, the results of $\mu_{\rm host}$ and $\sigma_{\rm host}$ correspond to a median rest-frame host DM of $108\pm 20~\mathrm{pc~cm^{-3}}$ (corresponding mean DM is $168^{+57}_{-40}~\mathrm{pc~cm^{-3}}$), which is consistent with both $\mathrm{DM}_{\rm host}=103_{-48}^{+68}~\mathrm{pc~cm^{-3}}$ Yang:2024vqq and $\mathrm{DM}_{\rm host}=130_{-23}^{+25}~\mathrm{pc~cm^{-3}}$ connor2025gas.
-Appendix B FRB data
-In this Appendix, we list the properties of the $104$ localized FRBs used in this work in Table LABEL:tab:localized_FRBs.
-Table 2: Properties of $104$ localized FRBs.
-The listed parameters include redshift, right ascension (R.A.), declination (Decl.), observed dispersion measure (DM), the Milky Way ISM contribution derived from the NE2001 model, and types of their host galaxies.
-The host types are classified into three categories: type 1 refers to repeating FRBs in dwarf galaxies, type 2 to repeating FRBs in spiral galaxies, and type 3 to one-off FRBs.
-Name Redshift $\mathrm{DM}_{\mathrm{obs}}$ $\mathrm{DM}_{\mathrm{MW,ISM}}$ (NE2001) R.A. (J2000) Decl. (J2000) Host Type Reference
-(pc cm-3) (pc cm-3) () ()
-FRB 20121102A 0.19273 557 188.4 82.9946 $33.1479$ 1 Chatterjee:2017dqg; Tendulkar2016
-FRB 20180301A 0.3304 552 151.7 93.2268 $4.6711$ 1 Bhandari2022
-FRB 20180924B 0.3214 361.42 40.5 326.1053 40.9-40.9 3 Bannister2019
-FRB 20181112A 0.4755 589.27 41.7 327.3485 52.9709-52.9709 3 Prochaska2019
-FRB 20181223C 0.03024 111.61 19.9 180.9207 $27.5476$ 3 Bhardwaj2024
-FRB 20190102C 0.2912 364.5 57.4 322.4157 79.4757-79.4757 3 Bhandari2020
-FRB 20190110C 0.12244 221.6 37.1 249.3185 $41.4434$ 2 Ibik2024
-FRB 20190303A 0.064 223.2 29.8 207.9958 $48.1211$ 2 Michilli2023
-FRB 20190418A 0.07132 182.78 70.2 65.8123 $16.0738$ 3 Bhardwaj2024
-FRB 20190523A 0.66 760.8 37.2 207.065 $72.4697$ 3 Ravi2019
-FRB 20190608B 0.11778 338.7 37.3 334.0199 7.8983-7.8983 3 Chittidi2021
-FRB 20190611B 0.3778 321.4 57.8 320.7456 79.3976-79.3976 3 Heintz2020
-FRB 20190614D 0.6 959.2 87.8 65.0755 $73.7067$ 3 Law2020
-FRB 20190711A 0.522 593.1 56.5 329.4193 80.358-80.358 1 Heintz2020
-FRB 20190714A 0.2365 504.13 38.5 183.9797 13.021-13.021 3 Heintz2020
-FRB 20191001A 0.234 506.92 44.2 323.3513 54.7478-54.7478 3 Heintz2020
-FRB 20191106C 0.10775 332.2 25 199.5801 $42.9997$ 2 Ibik2024
-FRB 20191228A 0.2432 297.5 32.9 344.4304 28.5941-28.5941 3 Bhandari2022
-FRB 20200223B 0.06024 201.8 45.6 8.2695 $28.8313$ 2 Ibik2024
-FRB 20200430A 0.1608 380.1 27.2 229.7064 $12.3763$ 3 Heintz2020
-FRB 20200723B 0.0085 244 33 190.1583 5.135-5.135 3 2024arXiv241007307S
-FRB 20200906A 0.3688 577.8 35.8 53.4962 14.0832-14.0832 3 Bhandari2022
-FRB 20201124A 0.098 413.52 139.9 77.0146 $26.0607$ 2 Ravi2022
-FRB 20210117A 0.2145 729.1 34.4 339.9792 16.1515-16.1515 3 Bhandari2023
-FRB 20210320C 0.2797 384.8 39.3 204.4608 16.1227-16.1227 3 Shannon2024
-FRB 20210410D 0.1415 571.2 56.2 326.0863 79.3182-79.3182 3 Caleb2023
-FRB 20210603A 0.1772 500.147 39.5 10.2741 $21.2263$ 3 Cassanelli2023
-FRB 20210807D 0.1293 251.3 121.2 299.2214 0.7624-0.7624 3 Gordon2023
-FRB 20211127I 0.0469 234.83 42.5 199.8082 18.8378-18.8378 3 Gordon2023
-FRB 20211203C 0.3439 636.2 63.7 204.5625 31.3801-31.3801 3 Gordon2023
-FRB 20211212A 0.0715 206 27.1 157.3509 $1.676889$ 3 Gordon2023
-FRB 20220105A 0.2785 583 22 208.8039 $22.4665$ 3 Gordon2023
-FRB 20220204A 0.4012 612.584 50.7 274.2263 $69.7225$ 3 Sharma2024
-FRB 20220207C 0.04304 262.38 76.1 310.1995 $72.8823$ 3 Law2024
-FRB 20220208A 0.351 437 101.6 319.3483 $71.54$ 3 Sharma2024
-FRB 20220307B 0.2507 499.15 128.2 350.8745 $72.1924$ 3 Law2024
-FRB 20220310F 0.477958 462.24 46.3 134.7204 $73.4908$ 3 Law2024
-FRB 20220330D 0.3714 468.1 38.6 165.7256 $71.7535$ 3 Sharma2024
-FRB 20220418A 0.622 623.25 36.7 219.1056 $70.0959$ 3 Law2024
-FRB 20220501C 0.381 449.5 30.6 352.3792 32.4907-32.4907 3 Shannon2024
-FRB 20220506D 0.30039 396.97 84.6 318.0448 $72.8273$ 3 Law2024
-FRB 20220509G 0.0894 269.53 55.6 282.67 $70.2438$ 3 Law2024
-FRB 20220521B 1.354 1342.9 138.8 351.036 $71.138$ 3 Shannon2024
-FRB 20220529A 0.1839 246 40 19.1042 $20.6325$ 1 Li:2025ckl
-FRB 20220610A 1.016 1458.15 31 351.0732 33.5137-33.5137 3 Ryder2023
-FRB 20220717A 0.36295 637.34 118.3 293.3042 19.2877-19.2877 3 Rajwade2024
-FRB 20220725A 0.1926 290.4 30.7 353.3152 35.9902-35.9902 3 Shannon2024
-FRB 20220726A 0.361 686.55 89.5 73.94567 $69.9291$ 3 Sharma2024
-FRB 20220825A 0.241397 651.24 78.5 311.9815 $72.585$ 3 Law2024
-FRB 20220914A 0.1139 631.28 54.7 282.0568 $73.3369$ 3 Law2024
-FRB 20220918A 0.491 656.8 153.1 17.5921 $70.8113$ 3 Shannon2024
-FRB 20220920A 0.158239 314.99 39.9 240.2571 $70.9188$ 3 Law2024
-FRB 20221012A 0.284669 441.08 54.3 280.7987 $70.5242$ 3 Law2024
-FRB 20221027A 0.229 452.5 47.2 129.6104 $71.7315$ 3 Sharma2024
-FRB 20221029A 0.975 1391.75 43.8 141.9634 $72.4523$ 3 Sharma2024
-FRB 20221101B 0.2395 491.554 131.2 342.2162 $70.6812$ 3 Sharma2024
-FRB 20221106A 0.2044 343.8 34.8 56.7048 25.5698-25.5698 3 Shannon2024
-FRB 20221113A 0.2505 411.027 91.7 71.411 $70.3074$ 3 Sharma2024
-FRB 20221116A 0.2764 643.448 132.3 21.2102 $72.6539$ 3 Sharma2024
-FRB 20221219A 0.553 706.708 44.4 257.6298 $71.6268$ 3 Sharma2024
-FRB 20230124A 0.0939 590.574 38.6 231.9163 $70.9681$ 3 Sharma2024
-FRB 20230203A 0.1464 420.1 67.299 151.66159 $35.6941$ 3 Amiri:2025sbi
-FRB 20230216A 0.531 828 38.5 155.9717 $1.4678$ 3 Sharma2024
-FRB 20230222A 0.1223 706.1 33.31 106.96036 $11.22452$ 3 Amiri:2025sbi
-FRB 20230222B 0.11 187.8 56.9977 238.73912 $30.8987$ 3 Amiri:2025sbi
-FRB 20230307A 0.271 608.9 37.6 177.7813 $71.6956$ 3 Sharma2024
-FRB 20230311A 0.1918 364.3 67.2352 91.10966 $55.94595$ 3 Amiri:2025sbi
-FRB 20230501A 0.301 532.5 125.6 340.0272 $70.9222$ 3 Sharma2024
-FRB 20230506C 0.3896 772 68 12.0998 $42.0061$ 2 AnnaThomas2025
-FRB 20230526A 0.157 361.4 31.9 22.2326 52.7173-52.7173 3 Shannon2024
-FRB 20230626A 0.327 451.2 39.2 240.7125 $71.7142$ 3 Sharma2024
-FRB 20230628A 0.1265 344.952 39 166.7867 $72.2818$ 3 Sharma2024
-FRB 20230703A 0.1184 291.3 38.1566 184.62445 $48.72993$ 3 Amiri:2025sbi
-FRB 20230708A 0.105 411.51 60.3 303.1155 55.3563-55.3563 3 Shannon2024
-FRB 20230712A 0.4525 587.567 39.2 167.3585 $72.5578$ 3 Sharma2024
-FRB 20230730A 0.2115 312.5 114.6465 54.66456 $33.1593$ 3 Amiri:2025sbi
-FRB 20230808F 0.3472 653.2 31 53.3041 51.9353-51.9353 3 2025MNRAS.538.1800H
-FRB 20230814A 0.553 696.4 104.8 335.9748 $73.0259$ 3 connor2025gas
-FRB 20230902A 0.3619 440.1 34.1 52.1398 47.3335-47.3335 3 Shannon2024
-FRB 20230930A 0.0925 456 70 10.5 $41.4$ 3 AnnaThomas2025
-FRB 20231011A 0.0783 186.3 58.5524 18.2411 $41.7491$ 3 Amiri:2025sbi
-FRB 20231017A 0.245 344.2 31.1515 346.75429 $36.65268$ 3 Amiri:2025sbi
-FRB 20231025B 0.3238 368.7 65.5787 270.78807 $63.98908$ 3 Amiri:2025sbi
-FRB 20231120A 0.07 438.9 43.8 143.6169 $71.7574$ 3 Sharma2024
-FRB 20231123A 0.0729 302.1 37.1085 82.62325 $4.47554$ 3 Amiri:2025sbi
-FRB 20231123B 0.2625 396.7 40.2 242.5382 $70.7851$ 3 Sharma2024
-FRB 20231128A 0.1079 331.6 64.7368 199.5782 $42.99271$ 1 Amiri:2025sbi
-FRB 20231204A 0.0644 221 34.9372 207.99903 $48.116$ 2 Amiri:2025sbi
-FRB 20231206A 0.0659 457.7 37.5114 112.44284 $56.25627$ 3 Amiri:2025sbi
-FRB 20231220A 0.3355 491.2 49.9 123.9087 $73.6599$ 3 connor2025gas
-FRB 20231226A 0.1569 329.9 38.1 155.3638 $6.1103$ 3 Shannon2024
-FRB 20231229A 0.019 198.5 65.1335 26.46783 $35.11292$ 3 Amiri:2025sbi
-FRB 20231230A 0.0298 131.4 32.2305 72.79761 $2.39398$ 3 Amiri:2025sbi
-FRB 20240114A 0.13 527.65 49.7 321.9161 $4.3292$ 1 Tian:2024ygd
-FRB 20240119A 0.376 483.1 38 224.4672 $71.6118$ 3 connor2025gas
-FRB 20240123A 0.968 1462 90.2 68.2625 $71.9453$ 3 connor2025gas
-FRB 20240201A 0.042729 374.5 38.6 149.9056 $14.088$ 3 Shannon2024
-FRB 20240209A 0.1384 176.518 63.1 289.8504 $86.0609$ 2 2025ApJ...979L..22E
-FRB 20240210A 0.023686 283.73 28.7 8.7796 28.2708-28.2708 3 Shannon2024
-FRB 20240213A 0.1185 357.4 40 166.1683 $74.0754$ 3 connor2025gas
-FRB 20240215A 0.21 549.5 47.9 268.4413 $70.2324$ 3 connor2025gas
-FRB 20240229A 0.287 491.15 38 169.9835 $70.6762$ 3 connor2025gas
-FRB 20240304A 0.2423 652.6 73.9 136.3304 16.2282-16.2282 3 Shannon2024
-FRB 20240310A 0.127 601.8 30.1 17.6219 44.4394-44.4394 3 Shannon2024
-Table 2: Continued